### v1

# Stochastic Calculus for Pathwise Observables of Markov-Jump Processes: Unification of Diffusion and Jump Dynamics

Abstract
Path-wise observables—functionals of stochastic trajectories—are at the heart of time-average statistical mechanics and are central to thermodynamic inequalities such as uncertainty relations, speed limits, and correlation-bounds. They provide a means of thermodynamic inference in the typical situation, when not all dissipative degrees of freedom in a system are experimentally accessible. So far, theories focusing on path-wise observables have been developing in two major directions, diffusion processes and Markov-jump dynamics, in a virtually disjoint manner. Moreover, even the respective results for diffusion and jump dynamics were derived with a patchwork of different approaches that are predominantly indirect. Stochastic calculus was recently shown to provide a direct approach to path-wise observables of diffusion processes, while a corresponding framework for jump dynamics remained elusive. In our work we develop, in an exact parallelism with continuous-space diffusion, a complete stochastic calculus for path-wise observables of Markov-jump processes. We formulate a “Langevin equation” for jump processes, define general path-wise observables, and establish their covariation structure, whereby we fully account for transients and time-inhomogeneous dynamics. We prove the known kinds of thermodynamic inequalities in their most general form and discus saturation conditions. We determine the response of path-wise observables to general (incl. thermal) perturbations and carry out the continuum limit to achieve the complete unification of diffusion and jump dynamics. Our results open new avenues in the direction of discrete-state analogs of generative diffusion models and the learning of stochastic thermodynamics from fluctuating trajectories.
preprint: APS/123-QED
I Introduction
A large variety of systems in and out of equilibrium with nominally discrete state spaces, or with continuous state spaces but well-defined metastable states, can be accurately and thermodynamically consistently described by Markov-jump dynamics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Intense research over the past few decades led to quite deep understanding of the (stochastic) thermodynamics [3, 5, 6, 7, 6, 20, 9, 21, 22, 23, 24, 25] and kinetics [26, 27, 28, 29, 30, 31, 32, 24, 33, 34, 35] of Markov dynamics under so-called local detailed balance [11, 6, 36, 37], including fluctuation theorems [38, 39, 40, 41, 42, 5, 43], speed limits [44, 45, 46, 47, 48, 49, 50, 51], thermodynamic uncertainty relations (TUR) [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 9], anomalous relaxation phenomena [62, 32, 63, 64], nonequilibrium response relations [65, 66, 67, 68, 69, 18, 4, 70, 71], and recently descriptions within the framework of information geometry [72, 21, 30, 31, 30, 64].
From a practical perspective, state-of-the-art experiments, in particular single-particle tracking [73, 74] and single-molecule spectroscopy [75, 76, 77, 78, 79] probe stochastic time series. Single-molecule techniques invariably track low-, predominantly one-dimensional projections of high-dimensional dynamics (e.g., FRET efficiencies reporting on the instantaneous distance between fluorescent dyes [80, 81, 82, 83, 84], molecular extensions [77, 85, 78, 79, 86, 87, 84, 88] or on the instantaneous distance between plasmonic particles [89, 90] attached to the molecule). These projected time series are typically analyzed by averaging along individual realizations, thus giving rise to path-wise observables (i.e., functionals of stochastic paths) and leading naturally to the field of time-average statistical mechanics [91, 92, 93, 94, 73, 95].
Currently, the quantity of central interest is typically the thermodynamic cost—the total entropy production—of nonequilibrium processes that may be considered as the nonequilibrium “counterpart” of free energy. However, one almost never has experimental access to all dissipative degrees of freedom and therefore probes only coarse observables (i.e., projections) of Markov-jump dynamics. Diverse approaches have meanwhile been developed for thermodynamic inference from coarse observations [96, 9, 97, 98, 99, 100, 101, 102, 103, 19, 104]. Thermodynamic inequalities, that is, diverse speed limits [44, 45, 46, 47, 48, 49, 50, 51] and uncertainty relations [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 9, 105, 106, 107, 108, 109], in particular, attracted much attention, as they provide a general means to infer a lower bound on dissipation of the entire system from a measured observable. These thermodynamical inequalities in various forms (and valid under different conditions) provide important insight into how dissipation bounds fluctuations [52, 53, 54, 55, 56, 57, 58, 59, 60, 61], transport [45, 46, 110, 51], and correlations [111, 112, 113, 114] and, in turn, allow for efficient thermodynamic inference based on (generally coarse) observables. In addition to these three major classes of thermodynamic inequalities, alternative inference methods have been developed as well, based on the kinetic uncertainty relation [54], statistics of waiting times [97, 115], first-passage times [116, 55, 26, 53], via the so-called variance-sum rule [117, 118], from the statistics of coarsely observed trajectories [119, 100, 104], and others.
However, despite the general consensus about their importance and applicability, there is no unified approach to deriving and proving the above results. For example, there are strategies involving the Cramér-Rao inequality [120, 121, 54, 57, 122], a variational approach [113], the scaled cumulant generating function [59], large-deviation theory [123, 124], and the master fluctuation theorem [56], to name a few. Because the above methods are generally indirect, not much is known about the sharpness and conditions required to saturate the various inequalities.
In the case of diffusion processes, a direct method based on stochastic calculus was recently introduced [125, 61, 51, 126, 114, 71], which allows for a direct and unified approach. While the mathematical intricacies in correctly handling nowhere differentiable functions seem to still pose a barrier to the proliferation of the above ideas into the physics literature, the underlying concept of Itô calculus is well established.
In contrast, Markov-jump processes enjoy a much greater popularity in physics, and in particular in stochastic thermodynamics [6, 9, 8, 23, 127]. Here, master-equation approaches [128, 3, 29] (and corresponding “Quantum Hamiltonian” analogy [129]), Feynman-Kac “tilting” [130, 95], as well as approaches relying on path measures [131, 20, 6, 22] are widely adopted in the community. Nevertheless, as far as proving thermodynamic inequalities is concerned, these methods are not direct and do not provide insight about the sharpness and saturation conditions. A stochastic-calculus approach in the spirit of [125, 61, 51, 126, 114, 71] is therefore desired to provide a unification of results for diffusion and Markov-jump dynamics, and specifically of different approaches to Markov-jump dynamics.
The first steps in this direction, instigated in the Appendix of [61], have indeed already been made in the context of TURs [60] and linear response [70]. However, the full battery of results, which is required for a true unification and for the approach to be able compete with existing methods, remains to be developed.
Here, we carry out the complete program. We develop, in exact parallelism with the continuous diffusion counterpart, the stochastic equations of motions — a Langevin equation for Markov-jump processes with a central noise-time-Correlation Lemma — and define general pathwise observables (i.e., densities and currents). The results are throughout presented in direct analogy with the continuous-space framework. We allow for time-inhomogeneous dynamics (e.g., time-dependent driving). Equivalently to diffusions [96], Stratonovich increments emerge that govern current-type observables and allow for a direct proof of the complete covariations (incl. in transients) of general density- and current observables that takes the form of (generalized) Green-Kubo relations. We use the developed stochastic calculus to prove directly the known thermodynamic inequalities (TUR, correlation TUR, transport bounds, bounds on correlations etc.) in their most general form, assess their sharpness, and establish the respective saturation conditions. Moreover, fully within the continuous-space stochastic-calculus approach [71] we derive the response of an observable to a general (incl. thermal) perturbations for Markov-jump dynamics. Finally, we carry out the continuum limit to achieve the unification of diffusion and jump dynamics. We conclude with an outlook on open questions and future directions.
The manuscript is structured as follows. In Sec. II we first provide a summary of our main results. Next, in Sec. III we develop the stochastic calculus for jump processes, formalize general pathwise observables and establish their complete covariation structure. In Sec. IV we use the stochastic calculus to prove the (quite complete) list of thermodynamic inequalities in their most general form and apply them to thermodynamic inference on a selection of biophysical model systems. Following this, we show in Sec. V how the stochastic calculus can be applied to derive results for the response to general perturbations for discrete-state systems and how this relates to the fluctuation-dissipation theorem in equilibrium. In Sec. VI we take the continuum limit to show the equivalence of stochastic calculi for continuous- and discrete-state systems and in Sec. VII we extend the results to time-inhomogeneous dynamics. We conclude in Sec. VIII with an outlook. The mathematical proofs of the noise-time correlation Lemma and increment correlations, and other technical details are given in the Appendix A.
II Summary of Main Results
We aim to bridge the gap between continuous-space dynamics and Markov jump processes (MJP) using stochastic calculus. Throughout we will develop the results for MJP in analogy with those of diffusions, all results will therefore be shown in parallel to make the analogy as clear and natural as possible. The central results of the stochastic calculus for continuous-space diffusion and Markov-jump processes is highlighted in Fig. 1.
Figure 1: Comparison of key quantities in continuous and discrete space, namely the equations of motion, noise statistics, fundamental path observables, thermodynamic bounds, fluctuation-scale function, and response to perturbations (incl. temperature changes).
In the context of continuous-space dynamics we will focus on overdamped Langevin dynamics in $d$ dimensions governed by [132, 6, 61, 133]
$\mathrm{d}\bm{x}_{\tau}=\bm{F}(\bm{x}_{\tau})\mathrm{d}\tau+\bm{\sigma}\mathrm{d}\bm{W}_{\tau},$ (1)
for the $d$-dimensional position $\bm{x}_{\tau}$ in a drift field $\bm{F}(\bm{x}_{\tau})$ subject to thermal noise $\bm{\sigma}\mathrm{d}\bm{W}_{\tau}$, where $\mathrm{d}\bm{W}_{\tau}$ are increments of the $d$-dimensional Wiener process (i.e., $d$-dimensional “white noise”). In Sec. VII we will also allow for time-inhomogeneous (temporally driven) dynamics, while for clarity we will develop the results within the time-homogeneous setting.
We will prove in Sec. III that trajectories $x_{\tau}$ of MJP follow the matrix stochastic differential equation (i.e., a “discrete-space Langevin equation”)
$\mathrm{d}\bm{n}(\tau)=\bm{R}(x_{\tau},\tau)\mathrm{d}\tau+\mathrm{d}\bm{\varepsilon}(\tau),$ (2)
where the displacements have elements $(\mathrm{d}\bm{n}(\tau))_{xy}=\mathrm{d}n_{xy}(\tau)$ describing jumps from state $x$ to $y$ in $\tau,\tau+\mathrm{d}\tau]$. The matrix $(\bm{R}(x_{\tau},\tau))_{xy}\mathrm{d}\tau=\mathds{1}_{x_{\tau}=x}r_{xy}\mathrm{d}\tau$ encodes the expected (average) drift in $\tau,\tau+\mathrm{d}\tau]$ with time-independent transition rate $r_{xy}$ (following from the corresponding master equation with generator $(\bm{L})_{xy}=-\delta_{xy}\sum_{y\neq x}r_{xy}+(1-\delta_{xy})r_{yx}$), and the noise is encoded in the time-inhomogeneous shifted Poisson process $(\mathrm{d}\bm{\varepsilon}(\tau))_{xy}=\mathrm{d}\varepsilon_{xy}(\tau)$ [10, 134, 61, 60]. This noise has zero mean $\langle\mathrm{d}\varepsilon_{xy}(\tau)\rangle=0$ and the “white noise” property $\langle\mathrm{d}\varepsilon_{xy}(\tau)\mathrm{d}\varepsilon_{ij}(\tau^{\prime})\rangle_{x_{\tau}=k}=\delta_{ix}\delta_{jy}\delta_{ik}\delta(\tau-\tau^{\prime})r_{ij}\mathrm{d}\tau\mathrm{d}\tau^{\prime}$, where the expectation is conditioned on $x_{\tau}=k$.
We prove the central time-noise correlation lemma with $\mathrm{d}\tau_{i}(\tau)=\mathds{1}_{x_{\tau}=i}\mathrm{d}\tau$,
$\frac{\langle\mathrm{d}\varepsilon_{kl}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle_{x_{\tau}=x}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}$ (3)
$=\delta_{xk}\mathds{1}_{\tau<\tau^{\prime}}\left[P(i,\tau^{\prime}|l,\tau)-P(i,\tau^{\prime}|k,\tau)\right]r_{kl}+\mathcal{O}(\mathrm{d}\tau),$
which gives access to nontrivial correlations between differentials in Eq. (2)
$\langle\mathrm{d}n_{xy}(\tau^{\prime})\mathrm{d}\tau_{i}(\tau)\rangle^{x_{\tau^{\prime}}=x}_{x_{\tau}=i}$ $=r_{xy}P(x,\tau^{\prime}|i,\tau)p_{i}(\tau)\mathrm{d}\tau\mathrm{d}\tau^{\prime},$ (4)
$\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle_{x_{\tau}=x}^{x_{\tau^{\prime}}=i}$ $=r_{xy}P(i,\tau^{\prime},|y,\tau)p_{x}(\tau)\mathrm{d}\tau\mathrm{d}\tau^{\prime},$
$\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}n_{ij}(\tau^{\prime})\rangle_{x_{\tau}=x}^{x_{\tau^{\prime}}=i}$ $=r_{ij}r_{xy}P(i,\tau^{\prime},|y,\tau)p_{x}(\tau)\mathrm{d}\tau\mathrm{d}\tau^{\prime},$
where $P(x,\tau^{\prime}|i,\tau)$ is the transition probability (i.e., the “propagator”) from state $x$ at $\tau$ to state $i$ at $\tau\leq\tau^{\prime}$ and $p_{i}(\tau)$ is the probability to find the system in state $i$ at time $\tau$. Note that the averages are “pinned”, e.g., $\langle\cdot\rangle^{x_{\tau^{\prime}}=x}_{x_{\tau}=i}$ is the average conditioned on $x_{\tau}=i$} and $x_{\tau^{\prime}}=x$} at times $\tau\leq\tau^{\prime}$. The above correlations were shown in [60] using a different approach.
These correlations are required for the covariation structure of observables, such as currents and densities. Specifically, time-integrated currents are defined as [61]
$J_{t}^{\mathrm{c}}$ $=\int_{\tau=0}^{\tau=t}\bm{U}(\bm{x}_{\tau},\tau)^{T}\circ\mathrm{d}\bm{x}_{\tau},$
$J_{t}$ $=\int_{\tau=0}^{\tau=t}\mathrm{Tr}[\bm{\kappa}(\tau)^{T}\mathrm{d}\bm{n}(\tau)],$ (5)
for continuous-space and MJP, respectively, where the former is a Stratonovich integral over some vector-valued $\bm{U}(\bm{x}_{\tau},\tau)$. For MJP the transitions are weighted by a time-dependent antisymmetric transition weight $\bm{\kappa}(\tau)=-\bm{\kappa}(\tau)$. Hence, we identify Stratonovich integrals for MJP as integrals over traces of a two-point function and the jump increments. The above currents naturally decompose into a stochastic and “usual” integral [61], $J^{(\mathrm{c})}_{t}=J_{t}^{\mathrm{(c),I}}+J_{t}^{\mathrm{(c),II}}$, where the stochastic part is
$J_{t}^{\mathrm{c,I}}$ $=\int_{\tau=0}^{\tau=t}\bm{U}(\bm{x}_{\tau},\tau)^{T}\bm{\sigma}\mathrm{d}\bm{W}_{\tau}$
$J_{t}^{\mathrm{I}}$ $=\int_{\tau=0}^{\tau=t}\mathrm{Tr}[\bm{\kappa}(\tau)^{T}\mathrm{d}\bm{\varepsilon}(\tau)],$ (6)
respectively, while the ”usual” part reads $J_{t}^{\mathrm{(c),II}}=\int_{0}^{t}\mathcal{U}_{\tau}^{(c)}\mathrm{d}\tau$ where $\mathcal{U}_{\tau}^{c}\equiv\bm{U}(\bm{x}_{\tau},\tau)^{T}\bm{F}(\bm{x}_{\tau})+\nabla\cdot\left[\bm{D}\bm{U}(\bm{x}_{\tau},\tau)\right]$ and $\mathcal{U}_{\tau}\equiv\mathrm{Tr}[\bm{\kappa}(\tau)^{T}\bm{R}(x_{\tau})]$. Thus, we identify in the second line of Eq. (6) the Ito-type integral for observables of MJP by analogy to diffusions.
Similarly, time-integrated densities are defined as [61]
$\rho_{t}^{\mathrm{c}}$ $=\int_{0}^{t}\mathrm{d}\tau V(\bm{x}_{\tau},\tau),$
$\rho_{t}$ $=\int_{\tau=0}^{\tau=t}\bm{V}(\tau)^{T}\mathrm{d}\bm{\tau}(\tau),$ (7)
for continuous and discrete dynamics, respectively, where $V(\bm{x}_{\tau})$ and $\bm{V}(\tau)=(V_{1}(\tau),V_{2}(\tau),\dots)^{T}$ are scalar- and vector-valued functions, respectively.
(Co)variances of stationary densities and currents can be expressed in terms of integral operators [133, 96]
$\hat{\mathcal{I}}^{t,\mathrm{c}}_{M,N}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{\tau}^{t}\mathrm{d}\tau^{\prime}\int\mathrm{d}\bm{z}\int\mathrm{d}\bm{z}^{\prime}M(\bm{z})^{T}[\cdot]N(\bm{z}^{\prime}),$
$\hat{{I}}^{t}_{U^{\alpha},U^{\beta}}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{\tau}^{t}\mathrm{d}\tau^{\prime}\sum_{x,y}U_{xy}^{\alpha}\sum_{i,j}U^{\beta}_{ij}[\cdot],$ (8)
for continuous and discrete dynamics respectively. The operator in the first line of Eq (8) is already known [96, 125]. For $X_{t}^{\mathrm{(c)}}\in\{J_{t}^{\mathrm{(c)},k},\rho_{t}^{\mathrm{(c)},k}\}$ and $Y_{t}^{\mathrm{(c)}}\in\{J_{t}^{\mathrm{(c)},l},\rho_{t}^{\mathrm{(c)},l}\}$, $k,l\in\{1,2\}$, we find
$\mathrm{cov}(X_{t}^{\mathrm{c}},Y_{t}^{\mathrm{c}})$ $=\hat{\mathcal{I}}^{t,\mathrm{c}}_{X,Y}[{\Xi}_{m}^{\bm{z}\bm{z}^{\prime}}]-\langle X_{t}^{\mathrm{c}}\rangle\langle Y_{t}^{\mathrm{c}}\rangle,$
$\mathrm{cov}(X_{t},Y_{t})$ $=\hat{{I}}^{t}_{U^{X},U^{Y}}[{\Xi}^{m}_{ijxy}]-\langle X_{t}\rangle\langle Y_{t}\rangle,$ (9)
where $X_{t}^{\mathrm{(c)}},Y^{\mathrm{(c)}}_{t}$ are the appropriate pathwise observables and $U^{X},U^{Y}$ “window functions” that define them ($\bm{\kappa}$ in the case of MJP). Moreover, $m\in\{1,2,3\}$ indicates density-density ($m=1$), current-density ($m=2$), and current-current ($m=3$) covariances. For the density covariance ($m=1$), $X_{t}^{\mathrm{(c)}}=\rho_{t}^{\mathrm{(c)},k}$ and $Y_{t}^{\mathrm{(c)}}=\rho_{t}^{\mathrm{(c)},l}$, we get
${\Xi}_{1}^{\bm{z}\bm{z}^{\prime}}$ $=\langle\mathds{1}\rangle_{\bm{x}_{\tau}=\bm{z}}^{\bm{x}_{\tau^{\prime}}=\bm{z}^{\prime}}+\langle\mathds{1}\rangle_{\bm{x}_{\tau}=\bm{z}^{\prime}}^{\bm{x}_{\tau^{\prime}}=\bm{z}},$ (10)
${\Xi}^{1}_{ijxy}$ $=\frac{\langle\mathrm{d}\tau_{x}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle_{x_{\tau}=x}^{x_{\tau^{\prime}}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}+\frac{\langle\mathrm{d}\tau_{i}(\tau)\mathrm{d}\tau_{x}(\tau^{\prime})\rangle^{x_{\tau^{\prime}}=x}_{x_{\tau}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}},$
for current-density covariances ($m=2$) we recover
${\Xi}_{2}^{\bm{z}\bm{z}^{\prime}}$ $=\frac{\langle\circ\mathrm{d}\bm{x}_{\tau}\rangle_{\bm{x}_{\tau}=\bm{z}}^{\bm{x}_{\tau^{\prime}}=\bm{z}^{\prime}}}{\mathrm{d}\tau}+\frac{\langle\circ\mathrm{d}\bm{x}_{\tau^{\prime}}\rangle_{\bm{x}_{\tau}=\bm{z}^{\prime}}^{\bm{x}_{\tau^{\prime}}=\bm{z}}}{\mathrm{d}\tau^{\prime}},$ (11)
${\Xi}_{ijxy}^{2}$ $=\frac{\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle_{x_{\tau}=x}^{x_{\tau^{\prime}}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}+\frac{\langle\mathrm{d}n_{xy}(\tau^{\prime})\mathrm{d}\tau_{i}(\tau)\rangle^{x_{\tau^{\prime}}=x}_{x_{\tau}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}},$
and for the current-current covariance ($m=3$) we find
${\Xi}_{3}^{\bm{z}\bm{z}^{\prime}}$ $=\frac{\langle\circ\mathrm{d}\bm{x}_{\tau}\circ\mathrm{d}\bm{x}_{\tau^{\prime}}^{T}\rangle_{\bm{x}_{\tau}=\bm{z}}^{\bm{x}_{\tau^{\prime}}=\bm{z}^{\prime}}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}+\frac{\langle\circ\mathrm{d}\bm{x}_{\tau^{\prime}}\circ\mathrm{d}\bm{x}_{\tau}^{T}\rangle_{\bm{x}_{\tau}=\bm{z}^{\prime}}^{\bm{x}_{\tau^{\prime}}=\bm{z}}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}$ (12)
${\Xi}_{ijxy}^{3}$ $=\frac{\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}n_{ij}(\tau^{\prime})\rangle_{x_{\tau}=x}^{x_{\tau^{\prime}}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}+\frac{\langle\mathrm{d}n_{xy}(\tau^{\prime})\mathrm{d}n_{ij}(\tau)\rangle^{x_{\tau^{\prime}}=x}_{x_{\tau}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}.$
These above correlations can all be evaluated explicitly. For MJP we find, using $\int_{0}^{t}\mathrm{d}\tau\int_{0}^{\tau}\mathrm{d}\tau^{\prime}\to\int_{0}^{t}\mathrm{d}t^{\prime}(t-t^{\prime})$ [96], that
$\mathrm{cov}({\rho}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P_{\mathrm{s}}(x,\tau;i,\tau^{\prime})+P_{\mathrm{s}}(x,\tau^{\prime};i,\tau)-p_{x}^{\mathrm{s}}p_{i}^{\mathrm{s}}\right]$
$\mathrm{cov}({J}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\hat{j}_{xy}P(x,t^{\prime}|i)p_{i}^{\mathrm{s}}+\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}p_{i}^{\mathrm{s}}\right],$
$\mathrm{cov}(J_{t}^{k},{J}_{t}^{l})=$ $t\sum_{x,y}\kappa_{xy}^{k}\kappa_{xy}^{l}r_{xy}p_{x}^{\mathrm{s}}+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\hat{j}_{xy}\hat{j}_{ji}^{\ddagger}P(x,t^{\prime}|j)p_{j}^{\mathrm{s}}+\hat{j}_{ij}\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}J^{\mathrm{s}}_{ij}\right],$ (13)
where $V^{\alpha/\beta}\delta$ is a shorthand notation for $V^{\alpha/\beta}_{i}\delta_{ij}$, $J^{\mathrm{s}}_{xy}$ is the local steady-state current between $x$ and $y$, and $\hat{j}_{xy}^{\ddagger}$ is the so-called dual-reverse current operator (reversed steady-state current) [133, 96, 125] which for MJP reads $\hat{j}_{xy}=r_{xy}\to\hat{j}_{xy}^{\ddagger}=r_{yx}p_{y}^{\mathrm{s}}/p_{x}^{\mathrm{s}}$ [6]. The results for continuous-space diffusion are given in Refs. [96, 125]. For stationary initial conditions and with time-independent state function and transition weights, Eqs. (13) have the form of generalized Green-Kubo-like relations, in the sense that they relate (co)variances of observables to time integrals of (generalized) correlation functions of density- and current-observables (see Refs. [133, 96] for a discussion of diffusion processes). Notably, Eqs. (13) have a striking similarity to the continuous space expressions [96, 133]. Later we show how to to generalize Eqs. (13) to include time-dependent transition weights and state functions, as well as transient dynamics.
Thermodynamic inequalities
Since we now have a stochastic calculus for MJP at our disposal (incl. fluctuation and correlation relations), we can employ the same strategy as in [61] and derive thermodynamic inequalities from the Cauchy-Schwarz inequality by appropriately choosing an auxiliary functional. The first family of inequalities we prove in Sec. IV are generalized (incl. transient) correlation TURs [61, 57]
$\frac{\left(t\partial_{t}\langle J^{\mathrm{(c)}}_{t}\rangle-\langle\tilde{J}_{t}^{\mathrm{(c)}}\rangle-c(t)\left\{\left[t\partial_{t}-1\right]\langle\rho_{t}^{\mathrm{(c)}}\rangle-\langle\tilde{\rho}_{t}^{\mathrm{(c)}}\rangle\right\}\right)^{2}}{\mathrm{var}\left(J_{t}^{\mathrm{(c)}}-c(t)\rho_{t}^{\mathrm{(c)}}\right)}$
$\leq\frac{\Delta S_{\mathrm{tot}}^{\mathrm{(c)}}(t)}{2},$ (14)
where $c(t):\mathds{R}\to\mathds{R}$ is a deterministic weight function for the density and $\Delta S_{\mathrm{tot}}^{\mathrm{(c)}}$ is the discrete (continuous) space entropy production. In analogy with [61] we introduced the discrete space modified time-integrated density and current $\tilde{\rho}_{t}=\int_{\tau=0}^{\tau=t}\tau\partial_{\tau}\bm{V}(\tau)^{T}\mathrm{d}\bm{\tau}(\tau)$ and $\tilde{J}_{t}=\int_{\tau=0}^{\tau=t}\tau\mathrm{Tr}\left[(\partial_{\tau}\bm{\kappa}(\tau))^{T}\mathrm{d}\bm{n}(\tau)\right]$ that account for the dependence of $\bm{V}(\tau)$ and $\bm{\kappa}(\tau)$ on $\tau$.
Setting $c(t)=0$ we recover the transient TUR. The CTUR (in contrast to the TUR) allows to turn the Cauchy-Schwarz inequality into an equality for all times given a suitable choice of $\bm{V}(\tau)$ and $\bm{\kappa}(\tau)$. This choice, however, requires explicit and detailed information about the system and is therefore not applicable in most realistic applications. We can nevertheless optimize Eq. (14) with respect to $c(t)$ for given $\bm{V}(\tau)$ and $\bm{\kappa}(\tau)$ to find
$c^{*}(t)$ $=\frac{a(t)\mathrm{cov}({\rho}_{t},{J}_{t})-b(t)\mathrm{var}({J}_{t})}{a(t)\mathrm{var}({\rho}_{t})-b(t)\mathrm{cov}({\rho}_{t},{J}_{t})}\,$, (15)
where $a(t)=t\partial_{t}\langle J^{\mathrm{(c)}}_{t}\rangle-\langle\tilde{J}_{t}^{\mathrm{(c)}}\rangle$ and $b(t)=\left[t\partial_{t}-1\right]\langle\rho_{t}^{\mathrm{(c)}}\rangle-\langle\tilde{\rho}_{t}^{\mathrm{(c)}}\rangle$. Note that this optimal CTUR contains only operationally accessible quantities and affords an improved thermodynamic inference compared to using currents or densities alone. Interestingly, in contrast to the steady-state result [57], choosing an arbitrary $c(t)$ may even deteriorate the inference compared to the bare current or density TURs.
The second thermodynamic inequality we prove is the thermodynamic bound on the transport of any scalar observable $z(t)\equiv z(x_{\tau},\tau)$ [51]
$\frac{\langle z(t)-z(0)-\int_{0}^{t}\mathrm{d}\tau\partial_{\tau}z(\tau)\rangle^{2}}{t\mathcal{D}^{\mathrm{(c)}}_{t}}\leq\frac{\Delta S_{\mathrm{tot}}^{\mathrm{(c)}}}{2}\,$, (16)
where $\mathcal{D}^{\mathrm{(c)}}_{t}$ is the so-called fluctuation-scale function accounting for how much the observation stretches microscopic coordinates. In continuous space it was shown [113, 51] that
$\mathcal{D}_{t}^{\mathrm{c}}=\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\left\langle\left[\nabla_{\bm{x}}z_{\tau}\right]^{T}\bm{\gamma}^{-1}(\tau)\nabla_{\bm{x}}z_{\tau}\right\rangle.$ (17)
For MJP we in turn find
$\mathcal{D}_{t}\equiv\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\sum_{x,y}(z_{y}(\tau)-z_{x}(\tau))^{2}r_{xy}p_{x}(\tau).$ (18)
The fluctuation-scale function for both, continuous and discrete space dynamics, is operationally accessible by considering increments $\mathrm{d}z(\tau)=z(x_{\tau+\mathrm{d}\tau},\tau+\mathrm{d}\tau)-z(x_{\tau},\tau)$ along a trajectory and recognizing (as in continuous space [113, 51]) that
$\mathcal{D}_{t}^{\mathrm{(c)}}=\frac{1}{t}\int_{\tau=0}^{\tau=t}\mathrm{var}(\mathrm{d}z(\tau))\,.$ (19)
Contrary to the CTUR, the transport bound cannot be saturated in general. We provide a counterexample to illustrate how an attempt to saturate Eq. (15) in general fails.
The last family of thermodynamic bounds we prove are thermodynamic bounds on observable correlations (or, in short, “correlation bounds”) as a generalization of the bounds in Ref. [113]. It involves correlations of time-independent observables $V_{\tau}\equiv V(x_{\tau})=\sum_{k}\delta_{k,x_{\tau}}V_{k}$ and $z_{\tau}\equiv z(x_{\tau})=\sum_{k}\delta_{k,x_{\tau}}z_{k}$ and their time-average along a trajectory, e.g., $\overline{V}_{t}=t^{-1}\int_{0}^{t}V_{\tau}\mathrm{d}\tau$, resulting in a lower bound of the integrated entropy production rate $\dot{S}_{\mathrm{tot}}(\tau)$ weighted by a “pseudo variance” $\mathrm{pvar}_{\mathrm{ps}}^{F}(z_{\tau})=\sum_{x,y}p_{xy}^{\mathrm{ps}}(\tau)(z_{x}+z_{y}-2F(\tau))$ [114]
$\frac{\left[2\mathrm{cov}({V}_{t},\overline{{z}}_{t})-\frac{2}{t}\int_{0}^{t}\mathrm{d}\tau\mathrm{cov}(V_{\tau},z_{\tau})\right]^{2}}{t\mathrm{var}(\overline{V}_{t})}-\mathcal{D}_{t}+\frac{2}{t}\left[\mathrm{var}(z_{t})-\mathrm{var}(z_{0})\right]\leq\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\dot{S}_{\mathrm{tot}}(\tau)\mathrm{pvar}_{\mathrm{ps}}^{F}(z_{\tau}),$ (20)
where $F(\tau)$ may be any time-(in)dependent function. The tilted two-point probability $p_{xy}^{\mathrm{ps}}(\tau)$ requires knowledge about transition rates, hence the pseudo variance is not operationally accessible. However, it can be controlled by choosing bounded observables. In NESS, Eq. (20) reads
$\frac{4\left[\mathrm{cov}(V_{t},\overline{{z}}_{t})-\mathrm{cov}_{\mathrm{s}}({V},{z})\right]^{2}}{t\mathrm{var}(\overline{V}_{t})}-\mathcal{D}_{t}$
$\leq\frac{\Delta S_{\mathrm{tot}}(t)}{2t}\mathrm{pvar}_{\mathrm{ps}}^{F}({z})\,.$ (21)
In Sec. IV we provide an example to showcase how the correlation bound can be applied to systems where both, the CTUR and transport bound, fail to infer a non-zero entropy production.
Time Inhomogeneous Dynamics
The generalization of the above results to systems exposed to time-dependent driving with constant protocol speed $v$ (altering the generator as $\bm{L}\to\bm{L}(v\tau)$) is in most cases straightforward. Indeed, the covariation results (9), the CTUR (14), and the transport bound (16) generalize immediately. While the results for the covariation and the transport bound remain unaltered (up to the obvious change in the expectation operation $\langle\cdot\rangle$), the CTUR becomes altered with the differential operator $\hat{\Lambda}\equiv t\partial_{t}-v\partial_{v}$ and reads
$\frac{\left(\hat{\Lambda}\langle J_{t}\rangle-c(t)\left\{\left[\hat{\Lambda}-1\right]\langle\rho_{t}\rangle\right\}\right)^{2}}{\mathrm{var}(J_{t}-c(t)\rho_{t})}\leq\frac{\Delta S_{\mathrm{tot}}}{2},$ (22)
Conversely, the correlation bound does not easily generalize to include time-dependent driving.
Response to External Perturbations
The response of a single-time observable $O(t)$ of a continuous-space diffusion to a perturbation in the noise amplitude, which effectively is a temperature perturbation [135], $\bm{\sigma}\to\bm{\sigma}+\delta\tilde{\bm{\sigma}}$ was recently derived using stochastic calculus and reads [136]
$\partial_{\delta}O_{t}$ $\equiv\lim_{\delta\to 0}\frac{1}{\delta}\left(\langle O(t)\rangle_{\delta}-\langle O(t)\rangle\right)$ (23)
$=-\left\langle O(t)\int_{s=0}^{s=t}\bm{\sigma}(x_{s},s)^{-1}\bm{b}_{\mathrm{eff}}^{0}(x_{s},s)\mathrm{d}\bm{W}_{s}\right\rangle$ (24)
where $\langle\cdot\rangle$ is the average w.r.t. the unperturbed process in Eq. (1) [now with multiplicative noise $\bm{\sigma}(x,\tau)$] and $\langle\cdot\rangle_{\delta}$ is the average w.r.t. the perturbed process
$\mathrm{d}\bm{x}^{\delta}_{\tau}=F(\bm{x}^{\delta}_{\tau})\mathrm{d}\tau+\left[\bm{\sigma}(\bm{x}_{\tau}^{\delta},\tau)+\delta\tilde{\bm{\sigma}}(\bm{x}_{\tau}^{\delta},\tau)\right]\mathrm{d}\bm{W}_{\tau}.$ (25)
The effective drift $\bm{b}_{\mathrm{eff}}^{0}(\bm{x},s)]_{i}=\frac{1}{2}\left[\bm{\Sigma}(\bm{x},s)\bm{s}^{0}(\bm{x},s)\right]_{i}+\frac{1}{2}\sum_{j}\partial_{x_{j}}\Sigma_{ij}(\bm{x},s)$ can be written in terms of the score function $\bm{s}^{0}(\bm{x},t)=\nabla_{\bm{x}}\log p(\bm{x},t)$ of the instantaneous probability distribution of $\bm{x}_{\tau}$ in Eq. (1) and $\bm{\Sigma}=\bm{\sigma}\tilde{\bm{\sigma}}^{T}+\tilde{\bm{\sigma}}\bm{\sigma}^{T}$ [136].
We derive the equivalent results for MJP perturbing a general control parameter $\chi\to\chi+\delta$, where we find that Eq. (23) takes the form
$\partial_{\delta}O_{t}=$ $\left\langle O(t)\int_{s=0}^{s=t}\sum_{x,y\neq x}\frac{\tilde{r}_{xy}}{r_{xy}}\mathrm{d}\varepsilon_{xy}(s)\right\rangle,$ (26)
where the unperturbed system evolves according to Eq. (2) whereas the perturbed system obeys
$\mathrm{d}\bm{n}^{\delta}(\tau)=\bm{R}^{\delta}(x^{\delta}_{\tau})\mathrm{d}\tau+\mathrm{d}\bm{\varepsilon}^{\delta}(\tau),$ (27)
with rates $r^{\delta}_{xy}$ perturbed as
$r_{xy}^{\delta}$ $=r_{xy}+\delta\tilde{r}_{xy}+\mathcal{O}(\delta^{2}).$ (28)
For equilibrium systems where
$r_{xy}=D\mathrm{e}^{\left(\lambda E_{x}-(1-\lambda)E_{y}\right)/T},$ (29)
with mixing parameter $\lambda\in[0,1]$ and free energies $E_{x}$ of state $x$, we find that Eq. (26) for temperature perturbations $T\to T+\delta$ equals the equilibrium correlation function of $O$ and free energy $E$
$\partial_{\delta}O_{t}=$ $\frac{1}{T^{2}}\left(\langle OE\rangle_{\mathrm{eq}}-\langle O(t)E(0)\rangle_{\mathrm{eq}}\right)\equiv\frac{1}{T^{2}}C_{OE}(t).$ (30)
The result in Eq. (26) relates the response to an observable $O$ to a general small perturbation of the dynamics to correlation functions of the unperturbed dynamics (which, notably, for both MJP and diffusions are not necessarily operationally accessible). In the case of equilibrium dynamics, Eq. (26) reduces to the Fluctuation Dissipation Theorem (FDT) in Eq. (30). Therefore, Eq. (26) can in some sense be seen as a (mathematical) extension of the FDT to irreversible and transient systems. In Appendix A.8 we generalize the response in Eq. (26) to perturbations of pathwise observables of MJP, e.g.
$O(t)=\int_{s=0}^{s=t}\mathrm{Tr}[\bm{b}(s)^{T}\mathrm{d}\bm{\varepsilon}(s)],$
for some time-dependent hollow matrix $\bm{b}(s)$.
The results in Eqs. (24) and  (26) may be useful in the context of computer simulations of high-dimensional systems, where knowing the drift, diffusivity, and $\mathrm{d}\bm{W}_{s}$ (and their discrete-state counterparts) allows to evaluate perturbations without the need to solve the high-dimensional Fokker-Planck (or master) equation of the perturbed system. In this context, Eq. (26) is expected to be useful, e.g. when the diagonalization of the transition matrix becomes computationally too demanding.
In the following section we provide detailed definitions and derivations of the above results and present applications. Technical steps and Lemmas are deferred to the Appendix.
III Theory
III.1 Stochastic Equations of Motion
To facilitate the comparison between discrete and continuous state spaces we briefly summarize the basics of overdamped (for simplicity additive-noise) diffusions.
Let $(\bm{x}_{\tau})_{0\leq\tau\leq t}$ be a trajectory of $\bm{x}_{\tau}\in\mathds{R}^{d}$ evolving according to overdamped Langevin dynamics [61]
$\mathrm{d}\bm{x}_{\tau}=\bm{F}(\bm{x}_{\tau})\mathrm{d}\tau+\bm{\sigma}\mathrm{d}\bm{W}_{\tau},$ (31)
where $\mathrm{d}\bm{W}_{\tau}$ is the $d$-dimensional Wiener process with $\langle\mathrm{d}\bm{W}_{\tau}\rangle=0$ and $\langle(\mathrm{d}\bm{W}_{\tau})_{i}(\mathrm{d}\bm{W}_{\tau^{\prime}})_{j}\rangle=\delta_{ij}\delta(\tau-\tau^{\prime})\mathrm{d}\tau\mathrm{d}\tau^{\prime}$, $\bm{F}$ is a drift field, and $\bm{\sigma}$ is the noise amplitude yielding the diffusion matrix $\bm{D}=\bm{\sigma}\bm{\sigma}^{T}/2$. While some bounds for continuous trajectories presented in the next section also apply to underdamped trajectories, we here limit the discussion to overdamped systems. This restriction allows for consistent comparisons with jump dynamics, since these descriptions are valid for negligible and marginalized momenta [132, 1], respectively. The evolution of the probability density $P(\bm{x},\tau)$ of $\bm{x}_{\tau}$ is governed by the Fokker-Planck equation [137] $\partial_{\tau}P(\bm{x},\tau)=-{\nabla}\cdot\bm{j}(\bm{x},\tau)$, written as a continuity equation with probability current $\bm{j}(\bm{x},\tau)=\left[\bm{F}(\bm{x})-\bm{D}\nabla\right]P(\bm{x},\tau)$.
Figure 2: Schematic comparison of discrete and continuous trajectories. In (a), a continuous trajectory is shown. (b) shows a discrete trajectory in a three-state MJP. The transitions $2\to 3$ (blue), $3\to 1$ (orange), $1\to 2$ (green), and $3\to 2$ (red) are marked with the respective colors. The displacement and jump increments $\mathrm{d}x_{\tau}$ and $\mathrm{d}n_{xy}(\tau)$ are presented in (c) and (d) for the continuous and discrete trajectories. The colors in (d) correspond to the different transitions in (b).
Consider now a discrete state space $\mathcal{N}\subseteq\mathds{N}$ which, without loss of generality, we can enumerate as $\mathcal{N}=\{1,2,\dots,N\}$. Focusing for the time being on time-homogeneous systems (i.e., time-independent transition rates $r_{ij}$ between states $i$ and $j$) the state probability vector $p_{i}(\tau)$ evolves according to the master equation [14, 8]
$\frac{\mathrm{d}}{\mathrm{d}\tau}p_{i}(\tau)=\sum_{j\neq i}\left[r_{ji}p_{j}(\tau)-r_{ij}p_{i}(\tau)\right],$ (32)
where the summation is only over $j$. In terms in the incoming and outgoing fluxes $J_{xy}(\tau)=r_{xy}p_{x}(\tau)$ we have as $\frac{\mathrm{d}}{\mathrm{d}\tau}p_{i}(\tau){\color[rgb]{0,0,1}=}\sum_{j\neq i}[J_{ij}(\tau)-J_{ji}(\tau)]$, and in vector notation $\frac{\mathrm{d}}{\mathrm{d}\tau}\bm{p}(\tau)=\bm{L}\bm{p}(\tau)$ with generator $\bm{L}$ having entries $L_{ij}=\delta_{ij}r_{ii}+(1-\delta_{ij})r_{ji}$ and exit rate $r_{ii}=-\sum_{j\neq i}r_{ij}$. The transition probability (or “propagator”) is therefore given by $P(i,\tau^{\prime}-\tau|j)=P(i,\tau^{\prime}|j,\tau)=\left(\exp{(\tau^{\prime}-\tau)\bm{L}}\right)_{ij}$.
A path is fully determined by the sequence of states $(x_{i})$ and transition times $t_{i}$, $(x_{\tau})_{0\leq\tau\leq t}=(x_{0},t_{0}=0;x_{1},t_{1};\dots)$. The probability of such a trajectory is then [6, 36, 95, 138, 139]
$\mathds{P}\left[(x_{\tau})_{0\leq\tau\leq t}\right]=p_{x_{0}}(0)\mathrm{e}^{\sum_{k}\tau_{k}(t)r_{kk}}\prod_{i,j}(r_{ij})^{n_{ij}(t)},$ (33)
where $p_{x_{0}}(0)$ is the probability of the initial state $x_{0}$ and $n_{ij}(t)$ and $\tau_{k}(t)$ are the number of transitions $i\to j$ and the total time spent in state $k$ along the trajectory of length $t$, respectively, and can be formulated as [61]
$n_{ij}(t)$ $=\int_{\tau=0}^{\tau=t}\mathrm{d}n_{ij}(\tau),$
$\tau_{i}(t)$ $=\int_{\tau=0}^{\tau=t}\mathrm{d}\tau_{i}(\tau).$ (34)
Suppose $x_{\tau}=i\neq j$, then $\mathrm{d}n_{ij}(\tau)\sim\mathrm{Poi}(r_{ij}\mathrm{d}\tau)$ is Poisson distributed and describes the number of transitions $i$ to $j$ that occur in the time interval $\tau,\tau+\mathrm{d}\tau]$. In particular [10, 134, 61, 60],
$\mathrm{d}n_{ij}(\tau)\xrightarrow{\mathrm{d}\tau\to 0}\begin{cases}1,&\text{$x_{\tau}=i$ and transition $i\to j$}\\ &\text{occurs in $[\tau,\tau+\mathrm{d}\tau]$,}\\ 0,&\text{else}.\end{cases}$ (35)
Note that $n_{ij}(\tau)$ is not continuous (and therefore not differentiable). Similarly, the differential describing the time spent in a state is [10, 134, 61, 60]
$\mathrm{d}\tau_{i}(\tau)=\begin{cases}\mathrm{d}\tau,&x_{\tau}=i,\\ 0,&\text{else}.\end{cases}$ (36)
For any state $i$ along the trajectory, the probability to transition to another state $j$ in the $\mathrm{d}\tau$ is given by $r_{ij}\mathrm{d}\tau$. Therefore, it is natural to define a noise [61, 60, 70, 10]
$\mathrm{d}\varepsilon_{ij}(\tau)\equiv\mathrm{d}n_{ij}(\tau)-r_{ij}\mathrm{d}\tau_{i}(\tau),$ (37)
by subtracting the expected number of jumps, $r_{ij}\mathrm{d}\tau_{i}(\tau)$, from the actual jumps $\mathrm{d}n_{ij}(\tau)$. This noise has also been studied in the context of martingales [10].
Rearranging Eq. (37) yields the consistent stochastic equation of motion for $x\neq y$ for MJP
$\underbrace{\mathrm{d}n_{xy}(\tau)}_{\text{``displacement"}}=\underbrace{r_{xy}\mathds{1}_{x_{\tau}=x}}_{\text{``drift"}}\mathrm{d}\tau+\underbrace{\mathrm{d}\varepsilon_{xy}(\tau)}_{\text{``noise"}}.$ (38)
Defining the matrix $(\bm{R}(x_{\tau}))_{ij}\equiv(1-\delta_{ij})r_{ij}\mathds{1}_{x_{\tau}=i}$, we can write Eq. (38) as a stochastic matrix equation
$\mathrm{d}\bm{n}(\tau)=\bm{R}(x_{\tau})\mathrm{d}\tau+\mathrm{d}\bm{\varepsilon}(\tau).$ (39)
The differential matrices have entries $(\mathrm{d}\bm{n}(\tau))_{ij}=\mathrm{d}n_{ij}(\tau)$ and $(\mathrm{d}\bm{\varepsilon}(\tau))_{ij}=\mathrm{d}\varepsilon_{ij}(\tau)$, where
$\!\!\!\langle\mathrm{d}\varepsilon_{kl}(\tau)\mathrm{d}\varepsilon_{ij}(\tau^{\prime})\rangle_{x_{\tau}=x}$ $=\delta_{kx}\delta_{ik}\delta_{jl}\delta(\tau-\tau^{\prime})r_{kl}\mathrm{d}\tau\mathrm{d}\tau^{\prime},$ (40)
in complete analogy with the Wiener process $\mathrm{d}W_{\tau}\propto\sqrt{\mathrm{d}\tau}$. Equation (39) is a matrix evolution equation, while the continuous evolution, Eq. (31), is vectorial. However, it has the same structure as the Langevin equation, as it consist of a “displacement”, “deterministic drift”, and “noise”, see Eq. (38). We highlight the comparison in Tab. 1. It should be stressed that the diagonal terms in Eq. (39) are to be neglected, as they do not have any physically relevant meaning.
Note that $\mathrm{d}\bm{\varepsilon}(\tau)$ is not Gaussian noise, but rather a shifted (due to the $\bm{R}(x_{\tau})\mathrm{d}\tau$ term) time-inhomogeneous Poisson process. Explicitly, because for $x_{\tau}=x$ and $y\neq x$ the differential $\mathrm{d}n_{xy}(\tau)\sim\mathrm{Poi}(r_{xy}\mathrm{d}\tau)$ is Poisson, the values $\mathrm{d}\varepsilon_{xy}(\tau)$ can take for any $x_{\tau}$ are $\mathrm{d}\varepsilon_{xy}(\tau)\in\{0,-r_{xy}\mathrm{d}\tau,1-r_{xy}\mathrm{d}\tau,2-r_{xy}\mathrm{d}\tau,3-r_{xy}\mathrm{d}\tau,\dots\}$, where $\mathrm{d}\varepsilon_{xy}(\tau)=0$ happens if $x_{\tau}\neq x$. In case $x_{\tau}=x$, one can identify $P(\mathrm{d}\varepsilon_{xy}(\tau)=k-r_{xy}\mathrm{d}\tau|x_{\tau}=x)=\mathrm{poi}_{r_{xy}\mathrm{d}\tau}(k)$ for any $k\in\mathds{N}_{\geq 0}$. In the limit of small $\mathrm{d}\tau$ to linear order in $\mathrm{d}\tau$ the noise therefore takes values $\mathrm{d}\varepsilon_{xy}(\tau)\in\{0,-r_{xy}\mathrm{d}\tau,1-r_{xy}\mathrm{d}\tau\}$ with probabilities $1-p_{x}(\tau),\left(1-r_{xy}\mathrm{d}\tau\right)p_{x}(\tau),r_{xy}\mathrm{d}\tau p_{x}(\tau)$, respectively. The probability for any $k\geq 2$ are $\mathcal{O}(\mathrm{d}\tau^{k})$ and can therefore be neglected.
Additionally, the scaling of the noise $\mathrm{d}\bm{\varepsilon}(\tau)\sim\sqrt{\mathrm{d}\tau}$ [see Eq. (40)] implies that $|\mathrm{d}\bm{\varepsilon}(\tau)/\mathrm{d}\tau|\sim 1/\sqrt{\mathrm{d}\tau}$ and, therefore, $|\mathrm{d}\bm{n}(\tau)/\mathrm{d}\tau|\sim 1/\sqrt{\mathrm{d}\tau}$. The limit $\mathrm{d}\tau\to 0$ does not exist. This is similar to the well-known (and often ignored) issue with writing the Langevin equation in the “physicist’s” way using $\dot{\bm{x}}_{\tau}$.
Further similarities in the equations of motion become obvious, e.g., in Fig. 2. Any trajectory is fully characterized by $\mathrm{d}\bm{n}(\tau)$ or $\mathrm{d}\bm{x}_{\tau}$ (see Fig. 2). Moreover, as shown in Sec. A.3, the master equation follows from Eq. (39) in the same manner as the Fokker-Planck equation follows from Eq. (31).
We further require the statistical properties of the above stochastic differentials. Conditioning the expectation on being in a specific state, e.g. $x_{\tau}=x$, we recover for $\tau\leq\tau^{\prime}$
$\langle\mathrm{d}n_{ij}(\tau)\rangle_{x_{\tau}=x}$ $=\delta_{ix}r_{ij}\mathrm{d}\tau,$
$\langle\mathrm{d}\tau_{i}(\tau)\rangle_{x_{\tau}=x}$ $=\delta_{ix}\mathrm{d}\tau,$
$\langle\mathrm{d}\varepsilon_{ij}(\tau)\rangle_{x_{\tau}=x}$ $=0,$
$\langle\mathrm{d}\tau_{i}(\tau)\mathrm{d}\tau_{j}(\tau^{\prime})\rangle_{x_{\tau}=x}$ $=\delta_{ix}P(j,\tau^{\prime}|i,\tau)\mathrm{d}\tau^{\prime}\mathrm{d}\tau.$ (41)
We are also required to evaluate cross-correlations between the noise and the time spent in a state, which follows from the noise-time correlation lemma
$\frac{\langle\mathrm{d}\varepsilon_{kl}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle_{x_{\tau}=x}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}$ (42)
$=\delta_{xk}\mathds{1}_{\tau<\tau^{\prime}}\left[P(i,\tau^{\prime}|l,\tau)-P(i,\tau^{\prime}|k,\tau)\right]r_{kl}+\mathcal{O}(\mathrm{d}\tau).$
The above lemma has been shown before [60]; in App. A.1 we provide a new alternative proof. Note that, due to the Markovianty of the system, we do not need to restrict on $\tau\leq\tau^{\prime}$ in evaluating the expectation, as $\tau>\tau^{\prime}$ vanishes.
The above correlators allow, for example, to evaluate the covariance between the number of jumps ate different space-time points along the trajectory, i.e., for $\tau\leq\tau^{\prime}$, [60]
$\langle$ $\mathrm{d}n_{kl}(\tau)\mathrm{d}n_{ij}(\tau^{\prime})\rangle_{x_{\tau}=x}=\mathrm{d}\tau\mathrm{d}\tau^{\prime}\delta_{kx}r_{kl}$ (43)
$\times\left[\delta(\tau-\tau^{\prime})\delta_{ik}\delta_{jl}+\mathds{1}_{\tau=\tau^{\prime}}\delta_{ik}r_{ij}+\mathds{1}_{\tau<\tau^{\prime}}r_{ij}P(i,\tau^{\prime}|l,\tau)\right].$
The middle term can be neglected when integrating over $\tau$ and $\tau^{\prime}$ as it has measure zero. The first term accounts for jumps occurring at $\tau=\tau^{\prime}$, while the last term accounts for the jumps $k\to l$ at time $\tau$ that result in $i\to j$ at some later time $\tau^{\prime}>\tau$.
We may rewrite the path measure Eq. (33) with the stochastic differentials discussed above as
$\mathds{P}[(x_{\tau})_{0\leq\tau\leq t}]=p_{x_{0}}(0)\exp\left[\int_{s=0}^{s=t}\sum_{x}r_{xx}\mathrm{d}\tau_{x}(s)+\int_{s=0}^{s=t}\sum_{x,y\neq x}\log r_{xy}\mathrm{d}n_{xy}(s)\right],$ (44)
where the second sum is taken over both $x$ and $y\neq x$. The advantage of writing the path measure as in Eq. (44) instead of Eq. (33) is that it allows for possible time dependencies of the rates.
Meaning Continuous Discrete
Displacement $\mathrm{d}\bm{x}_{\tau}$ $\mathrm{d}\bm{n}(\tau)$
Deterministic Drift $\bm{F}(\bm{x}_{\tau})\mathrm{d}\tau$ $\bm{R}(x_{\tau})\mathrm{d}\tau$
Noise $\bm{\sigma}\mathrm{d}\bm{W}_{\tau}$ $\mathrm{d}\bm{\varepsilon}(\tau)$
Table 1: Overview of quantities entering the stochastic equations of motion in continuous and discrete space.
III.2 Pathwise Observables: Currents and Densities
We now introduce pathwise observables, i.e., time-integrated densities and currents that were recently analyzed extensively in the context of diffusions [61, 133, 96, 125]. As before, we start with the continuous space setting, where a time-integrated current is defined as [61]
$J_{t}^{\mathrm{c}}=\int_{\tau=0}^{\tau=t}\bm{U}(\bm{x}_{\tau},\tau)^{T}\circ\mathrm{d}\bm{x}_{\tau},$ (45)
as a Stratonovich integral over some (square integrable) vector-valued window function $\bm{U}(\bm{x}_{\tau},\tau)$, and a time-integrated density is defined as [61]
$\rho_{t}^{\mathrm{c}}=\int_{0}^{t}\mathrm{d}\tau V(\bm{x}_{\tau},\tau),$ (46)
where $V(\bm{x}_{\tau},\tau)$ is some (square integrable) scalar-valued function.
In the context of MJP the time-integrated currents are weigh the transitions occurring along a trajectory with a fully anti-symmetric, possibly time-dependent weight $\bm{\kappa}(\tau)$, i.e., $\kappa_{ij}(\tau)=-\kappa_{ji}(\tau)$. The current defined this way is manifestly anti-symmetric such that the current changes sign upon time reversal. Note that $\bm{\kappa}(\tau)$ describes experimentally accessible transitions. We therefore define a (scalar) time-integrated current as [61]
$J_{t}$ $\equiv\int_{\tau=0}^{\tau=t}\mathrm{Tr}[\bm{\kappa}(\tau)^{T}\mathrm{d}\bm{n}(\tau)].$ (47)
A comparison of Eqs. (45) and (47) highlights the fact that integrating over time and tracing over $\mathrm{d}\bm{n}(\tau)$ is the MJP correspondence to the Stratonovich integral (45). The stochastic integral Eq. (47) is time antisymmetric in analogy to the current defined for diffusion processes in the form of a Stratonovich integral Eq. (45). The current in Eq. (47) naturally decomposes (as for diffusions [61]) into a “usual time integral” $J_{t}^{\mathrm{II}}$ and the stochastic integral, $J_{t}=J_{t}^{\mathrm{I}}+J_{t}^{\mathrm{II}}$ where
$J_{t}^{\mathrm{I}}$ $=\int_{\tau=0}^{\tau=t}\mathrm{Tr}[\bm{\kappa}(\tau)^{T}\mathrm{d}\bm{\varepsilon}(\tau)],$ (48a)
$J_{t}^{\mathrm{II}}$ $=\int_{0}^{t}\mathrm{Tr}[\bm{\kappa}(\tau)^{T}\bm{R}(x_{\tau})]\mathrm{d}\tau,$ (48b)
i.e., integrals with traces over $\mathrm{d}\bm{\varepsilon}(\tau)$ are not necessarily time antisymmetric and therefore may be identified (per analogy) as Itô integrals due to the correspondence to the stochastic part of Eq. (45) [61]. This analogy (i) allows us to easier distinguish between the integrals over noise or jump increments and (ii) is fully consistent with observables of continuous-space diffusion. By means of the latter we will later demonstrate how continuous space Itô integrals emerge from the corresponding integrals of MJP with traces over $\mathrm{d}\bm{\varepsilon}(\tau)$ in the continuum limit. This will later provide a unified description of MJP and overdamped diffusions within the framework of stochastic calculus.
Similarly, a density assigns each state $i$ a weight $V_{i}(\tau)$ and then integrates these along the trajectory. These weights can be combined into the state function vector $\bm{V}(\tau)=(V_{1}(\tau),V_{2}(\tau),\dots)^{T}$. With $\bm{V}(\tau)$ we define the density [61]
$\rho_{t}=\int_{\tau=0}^{\tau=t}\bm{V}(\tau)^{T}\mathrm{d}\bm{\tau}(\tau),$ (49)
where $\mathrm{d}\bm{\tau}(\tau)=(\mathrm{d}\tau_{1}(\tau),\mathrm{d}\tau_{2}(\tau),\dots)^{T}$.
An exemplary trajectory with corresponding time-integrated density and current is shown in Fig. 3 to visualize Eqs. (47) and (49).
Figure 3: Example trajectory and examples of an arising current and density. The part of the trajectory in (a) which contributes to the density $\rho_{\tau}$ with state function, $\bm{V}^{T}=(0,1,0)$, is marked in blue. Similarly, the transitions contributing to the current $J_{\tau}$ with transition weights, $\kappa_{ij}=\delta_{i2}\delta_{j3}-\delta_{i3}\delta_{j2}$, are marked by orange arrows with numbers representing the weight of the transition. The resulting density and current can be seen in (b) and (c), respectively.
III.3 From Stratonovich Increments to Observable Correlations
In the case of diffusion processes, correlations of pathwise observables were shown to provide deep information about the passage of “pinned” trajectories through the respective window in terms of generalized Green-Kubo like relations [96]. They are furthermore relevant in the context of thermodynamic inequalities, e.g., TURs [61, 126] and thermodynamic correlation bounds (CB) [113, 114]. Following Refs. [96, 125, 133] we now show how the steady-state correlations of stochastic differentials relate to covariances of pathwise observables of MJP. In particular, we emphasize how these compare to the corresponding continuous-space results.
In order to express (co)variances of continuous and discrete currents, Eq. (45) and Eq. (47), and densities, Eq. (46) and Eq. (49), in a compact unifying form [96, 125, 133] we define the integration operators
$\hat{\mathcal{I}}^{t,\mathrm{c}}_{M,N}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{\tau}^{t}\mathrm{d}\tau^{\prime}\int\mathrm{d}\bm{z}\int\mathrm{d}\bm{z}^{\prime}M(\bm{z})^{T}[\cdot]N(\bm{z}^{\prime}),$ (50a)
$\hat{{I}}^{t}_{U^{\alpha},U^{\beta}}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{\tau}^{t}\mathrm{d}\tau^{\prime}\sum_{x,y}U_{xy}^{\alpha}\sum_{i,j}U^{\beta}_{ij}[\cdot],$ (50b)
where $\cdot$] in the first line denotes functions which depend on $\bm{z}$, $\bm{z}^{\prime}$, $\tau$, and $\tau^{\prime}$. Similarly, the functions entering $\hat{{I}}^{t}_{U^{\alpha},U^{\beta}}[\cdot]$ depend on $x$, $y$, $i$, $j$, $\tau$, and $\tau^{\prime}$. The functions $M(\bm{z})$ and $N(\bm{z})$ may be scalar- or vector-valued functions, while $U^{\alpha}$ and $U^{\beta}$ are two-point quantities, e.g., $\bm{\kappa}$. Equation (50b) corresponds to Eq. (50a) where each continuous spatial integral has been changed to a discrete double sum. While one may expect the integrals to be substituted by a single sum, the double sum is indeed required to include non-local quantities, e.g., currents.
We will now consider two observables $X_{t}^{\mathrm{(c)}}\in\{J_{t}^{\mathrm{(c),k}},\rho_{t}^{\mathrm{(c),k}}\}$ and $Y_{t}^{\mathrm{(c)}}\in\{J_{t}^{\mathrm{(c),l}},\rho_{t}^{\mathrm{(c),l}}\}$, where the superscript is used to distinguish discrete from the continuous setting and $k,l\in\{1,2\}$ is the index for currents $J_{t}^{\mathrm{(c),1}},J_{t}^{\mathrm{(c),2}}$ and densities $\rho_{t}^{\mathrm{(c),1}},\rho_{t}^{\mathrm{(c),2}}$. These currents and densities may have different current-/density-defining “window” functions, which we take into consideration here.
As shown in Refs. [96, 125, 133], the continuous (co)variances can be written as compactly in terms of Eq. (50a) together with fundamental increment correlations. We show here that the same structure appears for the discrete case. In fact, the covariances can be written as
$\mathrm{cov}(X_{t}^{\mathrm{c}},Y_{t}^{\mathrm{c}})$ $=\hat{\mathcal{I}}^{t,\mathrm{c}}_{X,Y}[{\Xi}_{m}^{\bm{z}\bm{z}^{\prime}}]-\langle X_{t}^{\mathrm{c}}\rangle\langle Y_{t}^{\mathrm{c}}\rangle,$
$\mathrm{cov}(X_{t},Y_{t})$ $=\hat{{I}}^{t}_{U^{X},U^{Y}}[{\Xi}^{m}_{ijxy}]-\langle X_{t}\rangle\langle Y_{t}\rangle,$ (51)
where $m=1$ for $X_{t}^{\mathrm{(c)}}=\rho_{t}^{\mathrm{(c)},k}$ and $Y_{t}^{\mathrm{(c)}}=\rho_{t}^{\mathrm{(c)},l}$, w.l.o.g. $m=2$ for $X_{t}^{\mathrm{(c)}}=J_{t}^{\mathrm{(c)},k}$ and $Y_{t}^{\mathrm{(c)}}=\rho_{t}^{\mathrm{(c)},l}$, and $m=3$ for $X_{t}^{\mathrm{(c)}}=J_{t}^{\mathrm{(c)},k}$ and $Y_{t}^{\mathrm{(c)}}=J_{t}^{\mathrm{(c)},l}$. For density-density covariance ($m=1$), the increment correlations read [96]
${\Xi}_{1}^{\bm{z}\bm{z}^{\prime}}$ $=\langle\mathds{1}\rangle_{\bm{x}_{\tau}=\bm{z}}^{\bm{x}_{\tau^{\prime}}=\bm{z}^{\prime}}+\langle\mathds{1}\rangle_{\bm{x}_{\tau}=\bm{z}^{\prime}}^{\bm{x}_{\tau^{\prime}}=\bm{z}},$ (52)
${\Xi}^{1}_{ijxy}$ $=\frac{\langle\mathrm{d}\tau_{x}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle_{x_{\tau}=x}^{x_{\tau^{\prime}}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}+\frac{\langle\mathrm{d}\tau_{x}(\tau^{\prime})\mathrm{d}\tau_{i}(\tau)\rangle^{x_{\tau^{\prime}}=x}_{x_{\tau}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}.$
Here, $\langle\cdot\rangle_{\bm{x}_{\tau}=\bm{z}^{\prime}}^{\bm{x}_{\tau^{\prime}}=\bm{z}}$ is expectation conditioned on $\bm{x}_{\tau}=\bm{z}$ and $\bm{x}_{\tau^{\prime}}=\bm{z}^{\prime}$ with $\tau\leq\tau^{\prime}$ and analogously for the discrete case. Evaluating these correlations yields the stationary covariances [96, 125]
$\mathrm{cov}({\rho}_{t}^{\mathrm{c},k},{\rho}_{t}^{\mathrm{c},l})=$ $\hat{\mathcal{I}}^{t,\mathrm{c}}_{V^{k},V^{l}}\left[P_{\mathrm{s}}(\bm{z},\tau;\bm{z}^{\prime},\tau^{\prime})+P_{\mathrm{s}}(\bm{z}^{\prime},\tau;\bm{z},\tau^{\prime})-p_{\mathrm{s}}(\bm{z})p_{\mathrm{s}}(\bm{z}^{\prime})\right],$
$\mathrm{cov}({\rho}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P_{\mathrm{s}}(x,\tau;i,\tau^{\prime})+P_{\mathrm{s}}(x,\tau^{\prime};i,\tau)-p_{x}^{\mathrm{s}}p_{i}^{\mathrm{s}}\right].$ (53)
The two-point function $V^{k}\delta$ stands for the (effectively one-point) quantity $V^{k}_{i}\delta_{ij}$  [^1] Note that the $\delta$ is introduced for convenience to allow for compact notation using the integration operator $\hat{{I}}^{t}_{U^{\alpha},U^{\beta}}[\cdot]$. and $P_{\mathrm{s}}(\cdot;\cdot)$ denotes the joint probability for stationary systems, i.e., $P_{\mathrm{s}}(x,\tau;i,\tau^{\prime})=P(i,\tau^{\prime}|x,\tau)p_{x}^{\mathrm{s}}$.
Similarly, for current-density covariances ($m=2$), we need the following increment correlations [96]
${\Xi}_{2}^{\bm{z}\bm{z}^{\prime}}$ $=\frac{\langle\circ\mathrm{d}\bm{x}_{\tau}\rangle_{\bm{x}_{\tau}=\bm{z}}^{\bm{x}_{\tau^{\prime}}=\bm{z}^{\prime}}}{\mathrm{d}\tau}+\frac{\langle\circ\mathrm{d}\bm{x}_{\tau^{\prime}}\rangle_{\bm{x}_{\tau}=\bm{z}^{\prime}}^{\bm{x}_{\tau^{\prime}}=\bm{z}}}{\mathrm{d}\tau^{\prime}},$ (54)
${\Xi}_{ijxy}^{2}$ $=\frac{\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle_{x_{\tau}=x}^{x_{\tau^{\prime}}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}+\frac{\langle\mathrm{d}n_{xy}(\tau^{\prime})\mathrm{d}\tau_{i}(\tau)\rangle^{x_{\tau^{\prime}}=x}_{x_{\tau}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}},$
which include correlations of a Stratonovich increment $\circ\mathrm{d}\bm{x}_{\tau}$ and $\mathrm{d}n_{xy}(\tau)$ with a future event. The evaluation of such correlations is possible by means of the generalized time-reversal symmetry, i.e., dual-reversal symmetry [141, 133, 96, 125]. In both cases, the dual system (denoted by ) has the same steady state probability whereas the irreversible drift $\bm{j}_{\mathrm{s}}(\bm{x})$ has the opposite sign $\bm{j}^{\ddagger}_{\mathrm{s}}(\bm{x})=-\bm{j}_{\mathrm{s}}(\bm{x})$. It is useful to define current and dual current operators; in the continuous-space setting, these are [133, 96, 125] $\hat{\bm{j}}_{\bm{x}}=p_{s}(\bm{x})^{-1}\bm{j}_{\mathrm{s}}(\bm{x})-p_{\mathrm{s}}(\bm{x})\bm{D}\nabla_{\bm{x}}p_{\mathrm{s}}(\bm{x})^{-1}$ and $\hat{\bm{j}}_{\bm{x}}^{\ddagger}=p_{s}(\bm{x})^{-1}\bm{j}_{\mathrm{s}}(\bm{x})+p_{\mathrm{s}}(\bm{x})\bm{D}\nabla_{\bm{x}}p_{\mathrm{s}}(\bm{x})^{-1}$. The corresponding operators for discrete-state systems are much simpler and read $\hat{j}_{xy}=r_{xy}$ and $\hat{j}_{xy}^{\ddagger}=r_{yx}p_{y}^{\mathrm{s}}/p_{x}^{\mathrm{s}}$, respectively. The latter is the dual current operator reversing the steady-state currents [6] . When these current operators act on $p_{x}^{\mathrm{s}}$, they result in the probability flux and dual probability flux on edge $x\to y$, i.e., $\hat{j}_{xy}p^{\mathrm{s}}_{x}=r_{xy}p_{x}^{\mathrm{s}}=J_{xy}^{\mathrm{s}}$ and $\hat{j}^{\ddagger}_{xy}p^{\mathrm{s}}_{x}=r_{yx}p_{y}^{\mathrm{s}}=J_{yx}^{\mathrm{s}}$. By means of the flux operators, it can be shown that
$\langle\mathrm{d}n_{xy}(\tau^{\prime})\mathrm{d}\tau_{i}(\tau)\rangle^{x_{\tau^{\prime}}=x}_{x_{\tau}=i}$ $=\hat{j}_{xy}P(x,\tau^{\prime}-\tau|i)p^{\mathrm{s}}_{i}\mathrm{d}\tau\mathrm{d}\tau^{\prime}$
$=r_{xy}P(x,\tau^{\prime}-\tau|i)p^{\mathrm{s}}_{i}\mathrm{d}\tau\mathrm{d}\tau^{\prime},$ (55)
which is the current on edge $x\to y$ at time $\tau^{\prime}>\tau$ in addition to $x_{\tau}=i$. With an approach analogous to Ref. [96], we use the dual current operator to express the correlation of an initial transition with future events. To be precise, we find that (see App. A.2.1)
$\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle_{x_{\tau}=x}^{x_{\tau^{\prime}}=i}$ $=\hat{j}^{\ddagger}_{yx}P(i,\tau^{\prime}-\tau|y)p^{\mathrm{s}}_{y}\mathrm{d}\tau\mathrm{d}\tau^{\prime}$
$=r_{xy}P(i,\tau^{\prime}-\tau|y)p^{\mathrm{s}}_{x}\mathrm{d}\tau\mathrm{d}\tau^{\prime},$ (56)
as well as (see App. A.2.2)
$\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}n_{ij}(\tau^{\prime})\rangle_{x_{\tau}=x}^{x_{\tau^{\prime}}=i}$ $=\hat{j}_{ij}\hat{j}^{\ddagger}_{yx}P(i,\tau^{\prime}-\tau|y)p^{\mathrm{s}}_{y}\mathrm{d}\tau\mathrm{d}\tau^{\prime}$
$=r_{ij}r_{xy}P(i,\tau^{\prime}-\tau|y)p^{\mathrm{s}}_{x}\mathrm{d}\tau\mathrm{d}\tau^{\prime}.$ (57)
Equations (56) and (57) enable the calculation of (co)variances of $n_{ij}$ and thus provide valuable insight about, e.g., how noisy are transitions in a given system (for an explicit example see Figs. 5a and 5b on the statistics of transitions and dwell-time, respectively, in a model of secondary active transport (SAT) in Sec. IV). Moreover, we can rewrite the current-density covariances compactly and in a unified form as [96, 125]
$\mathrm{cov}({J}_{t}^{\mathrm{c},k},{\rho}_{t}^{\mathrm{c},l})$ $=\hat{\mathcal{I}}^{t,\mathrm{c}}_{\bm{U}^{k},V^{l}}\left[\hat{\bm{j}}_{\bm{z}^{\prime}}P(\bm{z}^{\prime},t^{\prime}|\bm{z})p_{\mathrm{s}}(\bm{z})+\hat{\bm{j}}^{\ddagger}_{\bm{z}}P(\bm{z}^{\prime},t^{\prime}|\bm{z})p_{\mathrm{s}}(\bm{z})-\bm{j}_{\mathrm{s}}(\bm{z})p_{\mathrm{s}}(\bm{z}^{\prime})\right],$
$\mathrm{cov}({J}_{t}^{k},{\rho}_{t}^{l})$ $=\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\hat{j}_{xy}P(x,t^{\prime}|i)p_{i}^{\mathrm{s}}+\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}p_{i}^{\mathrm{s}}\right].$ (58)
The only time dependence in the integrands in Eq. (58) is the time-translation invariant propagators $P(i,t|y)$, that is, the integrands depend only on the time difference $t$. Hence, the time integrals in Eqs. (50a) and (50b) may be simplified to $\int_{0}^{t}\mathrm{d}\tau\int_{0}^{\tau}\mathrm{d}\tau^{\prime}\to\int_{0}^{t}\mathrm{d}t^{\prime}(t-t^{\prime})$ [96].
It remains to consider current-current covariances, which require the correlations of Stratonovich increments [in case of MJP, Stratonovich differentials $\mathrm{d}n_{xy}(\tau)$ are defined by analogy; see Eq. (47)] at different times [96]
${\Xi}_{3}^{\bm{z}\bm{z}^{\prime}}$ $=\frac{\langle\circ\mathrm{d}\bm{x}_{\tau}\circ\mathrm{d}\bm{x}_{\tau^{\prime}}^{T}\rangle_{\bm{x}_{\tau}=\bm{z}}^{\bm{x}_{\tau^{\prime}}=\bm{z}^{\prime}}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}+\frac{\langle\circ\mathrm{d}\bm{x}_{\tau^{\prime}}\circ\mathrm{d}\bm{x}_{\tau}^{T}\rangle_{\bm{x}_{\tau}=\bm{z}^{\prime}}^{\bm{x}_{\tau^{\prime}}=\bm{z}}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}},$ (59)
${\Xi}_{ijxy}^{3}$ $=\frac{\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}n_{ij}(\tau^{\prime})\rangle_{x_{\tau}=x}^{x_{\tau^{\prime}}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}+\frac{\langle\mathrm{d}n_{xy}(\tau^{\prime})\mathrm{d}n_{ij}(\tau)\rangle^{x_{\tau^{\prime}}=x}_{x_{\tau}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}.$
Using again the generalized time-reversal symmetry, we can evaluate these correlations to obtain
$\mathrm{cov}({J}_{t}^{\mathrm{c},k},J_{t}^{\mathrm{c},l})=$ $2t\int\mathrm{d}\bm{z}(\bm{U}^{k}(\bm{z}))^{T}\bm{D}\bm{U}^{l}(\bm{z})p_{\mathrm{s}}(\bm{z})$
$+\hat{\mathcal{I}}^{t,\mathrm{c}}_{\bm{U}^{k},\bm{U}^{l}}\left[\hat{\bm{j}}_{\bm{z}}\left(\hat{\bm{j}}^{\ddagger}_{\bm{z}^{\prime}}\right)^{T}P(\bm{z},t^{\prime}|\bm{z}^{\prime})p_{\mathrm{s}}(\bm{z}^{\prime})+\hat{\bm{j}}^{\ddagger}_{\bm{z}}\hat{\bm{j}}_{\bm{z}^{\prime}}^{T}P(\bm{z}^{\prime},t^{\prime}|\bm{z})p_{\mathrm{s}}(\bm{z})-\bm{j}_{\mathrm{s}}(\bm{z})\bm{j}_{\mathrm{s}}(\bm{z}^{\prime})^{T}\right]$
$\mathrm{cov}(J_{t}^{k},{J}_{t}^{l})=$ $t\sum_{x,y}\kappa_{xy}^{k}\kappa_{xy}^{l}r_{xy}p_{x}^{\mathrm{s}}+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\hat{j}_{xy}\hat{j}_{ji}^{\ddagger}P(x,t^{\prime}|j)p_{j}^{\mathrm{s}}+\hat{j}_{ij}\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}J^{\mathrm{s}}_{ij}\right],$ (60)
where the analogy between continuous and discrete state spaces is evident. Note that one can derive the results for covariances in different, less direct, ways, e.g., using Feynman-Kac and (spectral) perturbation theory (see [95, 142] and references therein).
Equations (53), (58), and (60) have the form of generalized Green-Kubo relations, connecting (co)variances of pathwise observables to generalized (auto)correlation functions, which in the continuous-space setting have been discussed in Ref. [96]. Here we highlight the similarities and differences between discrete space and continuous space.
The mathematical structure of the results is (up to the analogy) the same, which should not come as a surprise, at least (but not only) in the sense that the MJP process can (under appropriate conditions on the drift $\bm{F}$ in Eq. (1), i.e., that the reversible contribution stems from a generalized potential with deep minima separated by high barriers [1, 2, 3, 11, 16]) be seen as the long-time coarse-grained limit of the continuous system.
The continuous space correlations in Eqs. (52), (54), and (59) are valid for all $t\geq 0$ and similarly for systems with nominally discrete degrees of freedom (such as spin systems). Conversely, if the MJP is meant as an approximate, long-time effective description of a system with continuous degrees of freedom, Eqs. (52), (54), and (59) described covariances of pathwise observables at sufficiently long times.
A non-zero $\mathrm{cov}(J_{t}^{\mathrm{c}},\rho_{t}^{\mathrm{c}})$ necessarily indicates that detailed balance is broken [96]. The same observation can be made in the discrete case, as the equilibrium dual operator $\hat{j}_{xy}^{{\ddagger},\mathrm{eq}}=\hat{j}_{xy}^{\mathrm{eq}}$ together with $\bm{\kappa}=\bm{\kappa}(\tau)=-\bm{\kappa}^{T}$ yields $\mathrm{cov}_{\mathrm{eq}}(J_{t},\rho_{t})=0$.
To provide preliminary intuition, density (co)variances (scaled by $t^{-1}$) are shown in Fig. 4 for the secondary active transport (SAT) model, which will be discussed in more detail in Sec. IV. In addition, in Fig. 5c, we show the mean current $J_{t}$ arising in the stationary SAT model with the magnitude of the typical fluctuations using the standard deviation, $\sqrt{\mathrm{var}(J_{t})}$, and how these can be bounded from below using the stationary thermodynamic uncertainty relation (TUR).
Figure 4: Two densities, ${\rho}_{t}^{\alpha}$ and ${\rho}_{t}^{\beta}$, are used to visualize density (co)variances in the SAT model, see Sec. IV.1.1. These correspond to the state functions $V_{i}^{\alpha}=\delta_{i1}+\delta_{i2}$ and $V_{i}^{\beta}=\delta_{i5}+\delta_{i6}$ measuring if molecules of type AA and $B$ are in the channel, respectively. In (a), the (scaled) variances $\mathrm{var}({\rho}^{\alpha/\beta}_{t})/t$ are shown together with the (scaled) covariance between them $\mathrm{cov}({\rho}^{\alpha}_{t},{\rho}^{\beta}_{t})/t$ from numerical simulations (colored) and analytical solutions Eq. (53) (dashed lines). The analytical variance of ${\rho}_{t}^{\alpha}$ (solid lines) and covariance (black dashed lines), both modulated by $1/t$, are shown in (b) for various values of $e^{\mathrm{out}}_{B}$. The inset shows $\mathrm{var}({\rho}_{t}^{\beta})/t$. The numerics in (a) are evaluated using $N=10^{4}$ trajectories sampled using the celebrated Gillespie algorithm [143, 144]. The initial condition is $p_{i}=(\delta_{i1}+\delta_{i3}+\delta_{i5})/3$ and the values of the parameters can be found in Tab. 3. In (b), $e^{\mathrm{out}}_{B}=40$ is used.
Figure 5: Statistics of transitions and waiting times in the stationary SAT model. The average number of transition $\langle n_{ij}\rangle$ are shown in $(a)$ for a selection of edges. In (b), the average time spent $\langle\tau_{i}\rangle$ in each state is presented. The steady-state currents resulting from the transitions in (a) are presented in (c). The red errorbars are the respective standard deviations $\sqrt{\mathrm{var}(J_{t})}$. In (c), the lower bound $|\langle J_{t}\rangle|\sqrt{2/\Delta S_{\mathrm{tot}}}\leq\sqrt{\mathrm{var}(J_{t})}$ resulting from the steady-state TUR is added as purple errorbars. We use $t=10$ in every panel.
III.4 Transient Covariances
The generalizations of Eqs. (53), (58), and (60) to transient dynamics has been performed in Ref. [133] for continuous space dynamics. We give a brief overview of the generalization for MJP in this section. The first difference arises in the integration operator Eq. (50b)
$\hat{{I}}^{t}_{U^{\alpha},U^{\beta}}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{0}^{t}\mathrm{d}\tau^{\prime}\sum_{x,y}U_{xy}^{\alpha}(\tau)\sum_{i,j}U^{\beta}_{ij}(\tau^{\prime})[\cdot],$ (61)
where $U^{\alpha}$ and $U^{\beta}$ are explicitly time-dependent. Note that also the integration limits differ, as the explicit time-dependence of the $U$s and probabilities requires the full range of both temporal integrals. It is at this point easier to consider the covariances not in terms of probability flux operators, but rather express them in terms of transition rates. The discrete covariances in Eqs. (53), (58), and (60) generalize to
$\mathrm{cov}({\rho}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P(x,\tau;i,\tau^{\prime})-p_{x}(\tau)p_{i}(\tau^{\prime})\right],$
$\mathrm{cov}({J}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\mathds{1}_{\tau>\tau^{\prime}}r_{xy}P(x,\tau|i,\tau^{\prime})p_{i}(\tau^{\prime})+\mathds{1}_{\tau<\tau^{\prime}}r_{yx}P(i,\tau^{\prime}|y,\tau)p_{y}(\tau)-r_{xy}p_{x}(\tau)p_{i}(\tau^{\prime})\right],$
$\mathrm{cov}(J_{t}^{k},{J}_{t}^{l})=$ $\int_{0}^{t}\mathrm{d}\tau\sum_{x,y}\kappa_{xy}^{k}(\tau)\kappa_{xy}^{l}(\tau)r_{xy}p_{x}(\tau)$
$+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\mathds{1}_{\tau>\tau^{\prime}}r_{xy}r_{ij}P(x,\tau|j,\tau^{\prime})p_{i}(\tau^{\prime})+\mathds{1}_{\tau<\tau^{\prime}}r_{ij}r_{xy}P(i,\tau^{\prime}|y,\tau)p_{y}(\tau)-r_{xy}p_{x}(\tau)r_{ij}p_{i}(\tau^{\prime})\right],$ (62)
For a generalization to time-dependent rates see Sec. VII, and for time-inhomogeneous diffusion see Ref. [133].
IV Application of Stochastic Calculus to Thermodynamic Bounds
IV.1 Model systems
While discussing various applications of the developed stochastic calculus for MJP we will throughout elude to biophysically relevant model systems shown in Fig. 6. Thereby we aim mostly at illustrating how to apply the theoretical results rather than to discuss the respective biological implications.
Figure 6: Overview of models used to visualize the results: (a) Markov model for secondary active transport (SAT); the states $(1,3,5)$ correspond to the protein funnel being open toward the cell exterior while states $(2,4,6)$ describe the funnel being open toward the cell interior. The transition rate for the conversion of the opening from the interior to the exterior (and vice versa) is $\gamma$, which in the presence of ligands AA or $B$ is scaled by a dimensionless factor $x$. AA and $B$ leave the funnel with rate $l_{A/B}$ and enter it with rate ${\rm e}^{\mathrm{out}/\mathrm{in}}_{A/B}$ outside/inside the cell. Of particular interest in the (extended) SAT model is the impact of ${\rm e}^{\mathrm{out}}_{B}$, i.e., the concentration of the helper molecule. (b) Schematic of how the SAT model in (a) may be extended to a “train” of SAT models to capture the number of AA molecules, $n$, transported through the cell membrane. (c) Schematic of the Markov model for calmodulin folding. We focus on the inference of dissipation by observing a current with weights $\kappa_{ij}^{(23)}=\delta_{i2}\delta_{j3}-\delta_{j2}\delta_{i3}$ and/or density with weight $V_{i}=(\delta_{i1}+\delta_{i2})/2$. (d) Four-state ring model with equilibrium transition rates parameterized by free energies $E_{i}$ for state $i$, local diffusion coefficient $D$, temperature $T$, and mixing parameter $\lambda$ [see Eq. (63)], and upon constant driving, respectively. We are interested in the response to temperature perturbations and investigate the effect on a single-time observable $A_{t}=\delta_{x_{t}1}+\delta_{x_{t}2}$ and a current with transition weights $\kappa_{ij}=\delta_{i4}\delta_{j1}-\delta_{i1}\delta_{j4}+\delta_{i3}\delta_{j2}-\delta_{i2}\delta_{j3}$. (e) 4-state ring model with constant driving having rates 1 in the clockwise and $\epsilon$ in counterclockwise direction; the stationary distribution is independent of $\epsilon$. The model is used for a comparison of the quality of thermodynamic bounds in situations where only the mesostates A and B, corresponding to states $(1,2)$ and $(3,4)$, respectively, are observed without the ability to distinguish the respective transition pathways between them.
IV.1.1 Secondary Active Transport
We first consider a Markov model of secondary active transport (SAT) of molecule $A$ through a protein funnel facilitated by molecule $B$ [145, 146, 147]. The model, shown in Fig. 6a, consists of six states; three [i.e., $(1,3,5)$] corresponding to the funnel being open toward the exterior of the cell and other three [i.e., $(2,4,6)$] open toward the interior of cell. In the absence of $A$ or $B$ bound, the funnel changes conformation with rate $\gamma$. Conversely, the crossing via $A$ or $B$-filled states occurs with $\gamma x$ for a dimensionless speedup factor $x$. AA and $B$ leave or enter the funnel inside and outside the cell with rates $l_{A/B}$ and ${\rm e}^{\mathrm{out/in}}_{A/B}$, respectively.
We further consider a periodic extension of coupled SAT models to describe the transport of molecules $A$ through the membrane (Fig. 6b). We leave the rates in the system unchanged, whereas the state space is extended to $\mathcal{N}\times\mathds{Z}$, with the additional degree of freedom accounting for the total number of $A$ molecules outside the cell.
IV.1.2 Calmodulin Folding Dynamics
Being a $\mathrm{Ca}^{2+}$ receptor, calmodulin plays an important role in physiological signaling processes [148]. The Markov model for the calmodulin folding pathway [88] is commonly used as a paradigmatic example in, e.g., the study of first passage times [149] and stochastic thermodynamics [59, 150]. The respective states represent different conformations of the protein and the transitions are changes of conformation [88]. Here we adopt the transition rates from Ref. [149] summarized in Tab. 2 while the topology of the folding network is shown in Fig. 6c. The calmodulin example will be used to gauge the quality of thermodynamic inference via the derived thermodynamic inequalities.
IV.1.3 Four-State Ring Toy Model
Here we consider, as a minimal example, a four-state model with states $\{1,2,3,4\}$ on a ring to visualize results. We will choose two different parameterizations of the model, visualized in Figs. 6d and 6e.
In the equilibrium (detailed-balance) setting, we assign to each state a free energy $E_{i}$, $i\in\{1,2,3,4\}$, such that the stationary probability is $p_{i}^{\mathrm{eq}}\propto\mathrm{e}^{-E_{i}/T}$ (note that we explicitly include the temperature here). For any $\lambda\in[0,1]$, the following transition rates obey detailed balance [20]
$r_{xy}=D\mathrm{e}^{(\lambda E_{x}-(1-\lambda)E_{y})/T}.$ (63)
We use the free energies listed in Tab. 4.
In the constant driving setting, we set the clockwise rates to unity whereas the counterclockwise rates are controlled by the parameter $\epsilon$. This way the steady-state distribution $p_{i}^{\mathrm{s}}=p=1/4$ is independent of $\epsilon$.
IV.2 Thermodynamic inequalities and inference
To highlight the power of the developed stochastic calculus for observables of MJP we now apply it to directly prove a broad selection of thermodynamic inequalities. Throughout we draw comparisons to continuous-space results and highlight some fundamental differences between the two. We immediately gauge the quality of the bounds by means of model examples and discuss the respective saturation conditions. The section thus entails a brief discussion of the crucial steps in the proofs of the thermodynamic bounds for diffusions.
IV.2.1 Entropy and Pseudo-Entropy Production
The central object of interest for thermodynamic inference is the mean total entropy production, which consist of the entropy change in the system as well as the surrounding medium. For continuous systems, the total entropy production (EP) is [6, 5, 139]
$\Delta S_{\mathrm{tot}}^{\mathrm{c}}(t)=\int_{0}^{t}\mathrm{d}\tau\int\mathrm{d}\bm{x}\frac{\bm{j}(\bm{x},\tau)^{T}\bm{D}(x)\bm{j}(\bm{x},\tau)}{P(\bm{x},\tau)}\,.$ (64)
In the direct proofs in Refs. [61, 51], the stochastic integral
$A_{t}^{\mathrm{c}}\equiv\int_{\tau=0}^{\tau=t}\frac{\bm{j}(\bm{x}_{\tau},\tau)}{P(\bm{x}_{\tau},\tau)}\left[2\bm{D}(\bm{x}_{\tau})\right]^{-1}\bm{\sigma}(\bm{x}_{\tau})\mathrm{d}\bm{W}_{\tau},$ (65)
is a central object, chosen to immediately yield $\langle(A_{t}^{\mathrm{c}})^{2}\rangle=\Delta S_{\mathrm{tot}}^{\mathrm{c}}(t)/2$ which allows for crucial simplifications in the proofs.
Conversely, in discrete-state systems the total entropy production reads [7, 5, 139]
$\Delta S_{\mathrm{tot}}(t)=\frac{1}{2}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}(r_{ij}p_{i}(\tau)-r_{ji}p_{j}(\tau)\log\frac{r_{ij}p_{i}(\tau)}{r_{ji}p_{j}(\tau)}.$ (66)
As a key quantity in the various proofs of thermodynamic bounds we introduce the auxiliary integral
$A_{t}$ $=\int_{\tau=0}^{\tau=t}\mathrm{Tr}[z_{\tau}^{T}\mathrm{d}\bm{\varepsilon}(\tau)],$ (67)
where
$(z_{\tau})_{ij}=\frac{r_{ij}p_{i}(\tau)-r_{ji}p_{j}(\tau)}{r_{ij}p_{i}(\tau)+r_{ji}p_{j}(\tau)}.$ (68)
Using Eq. (41) the mean is $\langle A_{t}\rangle=0$ and the variance can be identified as a “pseudo entropy production” [122, 45]
$\langle A_{t}^{2}\rangle=\frac{1}{2}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\frac{\left[r_{ij}p_{i}(\tau)-r_{ji}p_{j}(\tau)\right]^{2}}{r_{ij}p_{i}(\tau)+r_{ji}p_{j}(\tau)}.$ (69)
Applying the logarithmic inequality $(a-b)^{2}/(a+b)\leq(a+b)\log(a/b)/2$ $\forall a,b>0$, the total entropy production Eq. (66) can be bounded from below using Eq. (69), i.e., $\langle A_{t}^{2}\rangle\leq\Delta S_{\mathrm{tot}}(t)/2$. The logarithmic inequality is saturated only for $a=b$ where the right hand side vanishes, and therefore (for discrete state spaces) $\langle A_{t}^{2}\rangle=\Delta S_{\mathrm{tot}}(t)/2$ only at equilibrium where $\Delta S_{\mathrm{tot}}(t)=0$.
Note that the inability to saturate $\langle A_{t}^{2}\rangle\leq\Delta S_{\mathrm{tot}}(t)/2$ outside of equilibrium is the first fundamental difference between discrete and continuous systems. The former bounds the pseudo-entropy production (which in turn bounds the entropy production) and the latter directly the total entropy production. As a consequence, when discussing the saturation of bounds in discrete space we refer to lower bounds on the pseudo-entropy production, not the total entropy production. However, we will show in Sec. VI, if the continuum limit exist the discrete pseudo-entropy production in Eqs. (69) and (66) both converge to continuous entropy production in Eq. (64), which may be saturated far from equilibrium [61].
IV.2.2 General TUR and Correlation TUR
In the direct proof of the TUR in Refs. [61, 60], the time-integrated current is split as $J_{t}^{\mathrm{c}}=J_{t}^{\mathrm{c,I}}+J_{t}^{\mathrm{c,II}}$, with
$J_{t}^{\mathrm{c,I}}$ $=\int_{\tau=0}^{\tau=t}\bm{U}(\bm{x}_{\tau},\tau)^{T}\bm{\sigma}\mathrm{d}\bm{W}_{\tau},$
$J_{t}^{\mathrm{c,II}}$ $=\int_{0}^{t}{U}(\bm{x}_{\tau},\tau)\mathrm{d}\tau,$ (70)
where ${U}(\bm{x}_{\tau},\tau)=\bm{U}(\bm{x}_{\tau},\tau)^{T}\bm{F}(\bm{x}_{\tau})+\nabla\cdot\left[\bm{D}\bm{U}(\bm{x}_{\tau},\tau)\right]$.
The proof of the transient TUR for MJP, which can be found, e.g., in Ref. [60], is analogous to the continuous case [61]. We will reiterate the proof here, focusing on the use of stochastic calculus, as well as including densities to obtain the correlation TUR (CTUR).
We start from Eq. (39) and split the current in Eq. (47) as
$J_{t}$ $=J_{t}^{\mathrm{I}}+J_{t}^{\mathrm{II}},$ (71)
where
$J_{t}^{\mathrm{I}}$ $=\int_{\tau=0}^{\tau=t}\mathrm{Tr}[\bm{\kappa}(\tau)^{T}\mathrm{d}\bm{\varepsilon}(\tau)],$
$J_{t}^{\mathrm{II}}$ $=\int_{0}^{t}\mathrm{Tr}[\bm{\kappa}(\tau)^{T}\bm{R}(x_{\tau})]\mathrm{d}\tau,$ (72)
which are a stochastic and a “usual” integral respectively. Moreover, we have $\langle J_{t}^{\mathrm{I}}\rangle=0$ and
$\langle J_{t}^{\mathrm{II}}\rangle=\langle J_{t}\rangle=\int_{0}^{t}\mathrm{d}\tau\sum_{x,y}\kappa_{xy}(\tau)r_{xy}p_{x}(\tau).$ (73)
To prove the TUR we use the Cauchy-Schwarz inequality
$\langle A_{t}(J_{t}-\langle J_{t}\rangle)\rangle^{2}\equiv\langle A_{t}\Delta J_{t}\rangle^{2}\leq\mathrm{var}(J_{t})\langle A_{t}^{2}\rangle.$ (74)
One can easily show that $\langle A_{t}J_{t}^{\mathrm{I}}\rangle=\langle J_{t}\rangle$. Evaluating
$\langle$ $A_{t}J_{t}^{\mathrm{II}}\rangle$
$=$ $\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}\mathds{1}_{\tau<\tau^{\prime}}$ (75)
$\times\sum_{x,y}\left[P(i,\tau^{\prime}|y,\tau)-P(i,\tau^{\prime}|x,\tau)\right]p_{x}(\tau)r_{xy}Z_{xy}(\tau),$
requires Eq. (42) and two integrations by part (for details see App. A.4), which yield
$\langle A_{t}J_{t}^{\mathrm{II}}\rangle=(t\partial_{t}-1)\langle J_{t}\rangle-\langle\tilde{J}_{t}\rangle,$ (76)
where we introduced the modified current
$\tilde{J}_{t}$ $\equiv\int_{\tau=0}^{\tau=t}\tau\mathrm{Tr}[\left(\partial_{\tau}\bm{\kappa}(\tau)^{T}\right)\mathrm{d}\bm{n}(\tau)],$ (77)
that accounts for the change of $\bm{\kappa}(\tau)$ in time $\tau$. We thus proved the transient TUR
$\frac{\left[t\partial_{t}\langle J_{t}\rangle-\langle\tilde{J}_{t}\rangle\right]^{2}}{\mathrm{var}(J_{t})}\leq\langle A_{t}^{2}\rangle\leq\frac{\Delta S_{\mathrm{tot}}(t)}{2}.$ (78)
As in the continuous case [61], the modified current is necessary for the validity of Eq. (78).
With slight modifications Eq. (78) also holds for time-dependent driving as proven in Ref. [59] using scaled cumulant generating functions. In Ref. [60], the same TUR was proven using a stochastic calculus approach. Note that Eq. (78) is a special case of the results in Refs. [59, 60] for systems with constant rates and $\bm{\kappa}(v\tau)$ with $v=1$ as the protocol velocity. In other words, the modified current in Eq. (77) corresponds to $v\partial_{v}J_{t}|_{v=1}$ where the generator is constant in time.
As shown in Ref. [61], one needs to include a density to saturate Eq. (78). By means of a calculation that is analogous to the evaluation of density-current correlations in Sec. III.4 we here further show that
$\langle A_{t}\rho_{t}\rangle$ $=(t\partial_{t}-1)\langle\rho_{t}\rangle-\langle\tilde{\rho}_{t}\rangle,$
$\tilde{\rho}_{t}$ $\equiv\int_{\tau=0}^{\tau=t}\tau\partial_{\tau}\bm{V}(\tau)^{T}\mathrm{d}\bm{\tau}(\tau).$ (79)
By shifting the current $\Delta J_{t}\to\Delta J_{t}-c(t)\Delta\rho_{t}$ with some $c:\mathds{R}\to\mathds{R}$ in Eq. (74) already yields the transient correlation TUR (CTUR)
$\frac{\left[t\partial_{t}\langle J_{t}\rangle-\langle\tilde{J}_{t}\rangle-c(t)\left((t\partial_{t}-1)\langle\rho_{t}\rangle-\langle\tilde{\rho}_{t}\rangle\right)\right]^{2}}{\mathrm{var}(J_{t}-c(t)\rho_{t})}\leq\frac{\Delta S_{\mathrm{tot}}(t)}{2}.$ (80)
This route to proving thermodynamic bounds may thus legitimately called be referred to as direct and allows to discuss saturation in a straightforward manner, as we show next.
The l.h.s. of Eq. (80) corresponds to the estimated total entropy production $\Sigma_{\mathrm{est}}(t)$. To effectively visualize how much of the total entropy produce is recovered in the estimate we use the quality factor
$\mathcal{Q}=\frac{\Sigma_{\mathrm{est}}(t)}{\Delta S_{\mathrm{tot}}(t)}.$ (81)
For the TUR, CTUR, and TB, the quality factor is bounded by $0\leq\mathcal{Q}\leq 1$. We will later show that the correlation bound the quality factor only is trivially bounded from above $\mathcal{Q}^{\mathrm{CB}}\leq 1$.
IV.2.3 Optimization of TUR and CTUR
There are two ways to optimize Eq. (80), either by saturating Eq. (74) or by choosing the optimal $c(t)$ for fixed $J_{t}$ and $\rho_{t}$. The former requires $\Delta J_{t}-c(t)\Delta\rho_{t}\propto A_{t}$. Since $c(t)\neq 0$, we see that the saturation is achieved when $V_{i}(\tau)=\hat{\bm{e}}_{i}^{T}\bm{\kappa}(\tau)\bm{L}\hat{\bm{e}}_{i}/c(t)=\sum_{j}\kappa_{ij}(\tau)r_{ij}/c(t)$ and $\bm{\kappa}=c^{\prime}\bm{Z}$ for any $c^{\prime}\in\mathds{R}$. Here, $\hat{\bm{e}}_{i}$ is the $i$th unit vector. However, these choices do not saturate Eq. (80) but only the estimate for the pseudo-entropy production [see first inequality in Eq. (78)]. In other words, the “saturated” CTUR quality factor reads
$\mathcal{Q}^{\mathrm{cTUR}}_{\mathrm{sat}}=\frac{2\langle A_{t}^{2}\rangle}{\Delta S_{\mathrm{tot}}(t)}\leq 1.$ (82)
We again highlight, that $\mathcal{Q}^{\mathrm{CTUR}}_{\mathrm{sat}}=1$ in equilibrium or (as we show below) in the continuum limit (if said limit exists). Moreover, the optimization requires the rates and transient probability distribution to be known, which makes the saturated bound superfluous. Namely, knowing the rates and the transient probability distribution allows the EP to be determined exactly.
The saturation approach shown here is analogous to the approach for continuous space dynamics [61] as it involves choosing $V(\bm{x}_{\tau},\tau)$ such that $c(t)\rho_{t}^{\mathrm{c}}=J_{t}^{\mathrm{c,II}}$. Then, the vector-valued function $\bm{U}(\bm{x}_{\tau},\tau)$ needs to be chosen proportional to the integrand in $A_{t}^{\mathrm{c}}$.
Hence, it seems to be more practical to optimize the bound w.r.t. $c(t)$ for a given $J_{t}$ and $\rho_{t}$. The l.h.s. of Eq. (80), which we denote $h(t)$, has two extrema in $c(t)$
$c^{\dagger}(t)$ $=\frac{a(t)}{b(t)}\,$,
$c^{*}(t)$ $=\frac{a(t)\mathrm{cov}({\rho}_{t},{J}_{t})-b(t)\mathrm{var}({J}_{t})}{a(t)\mathrm{var}({\rho}_{t})-b(t)\mathrm{cov}({\rho}_{t},{J}_{t})}\,$, (83)
where $a(t)\equiv t\partial_{t}\langle{J}_{t}\rangle-\langle{\tilde{J}}_{t}\rangle$ and $b(t)\equiv(t\partial_{t}-1)\langle{\rho}_{t}\rangle-\langle{\tilde{\rho}}_{t}\rangle$. All quantities entering Eq. (83) are accessible from $J_{t}$ and $\rho_{t}$, making the optimization practically feasible. Note that $c^{\dagger}(t)$ and $c^{*}(t)$ correspond to the minimum and maximum of $h(t)$, respectively, as
$\frac{\partial^{2}}{\partial c^{2}}h(c,t)|_{c(t)=c^{\dagger}(t)}$ $=\frac{4b(t)^{2}}{\mathrm{var}\left({J}_{t}-\frac{a(t)}{b(t)}\rho_{t}\right)}\geq 0\,$, (84)
$\frac{\partial^{2}}{\partial c^{2}}h(c,t)|_{c(t)=c^{*}(t)}$ $=\frac{-4b(t)^{2}\mathrm{var}\left({J}_{t}-\frac{a(t)}{b(t)}\rho_{t}\right)}{\left(\mathrm{var}\left({J}_{t}-c^{*}(t)\rho_{t}\right)\right)^{2}}\leq 0\,.$
Moreover, one can immediately see that using $c^{\dagger}(t)$ one recovers the second law of thermodynamics, as $t\partial_{t}\langle J_{t}\rangle-\langle\tilde{J}_{t}\rangle-c^{\dagger}(t)\left[(t\partial_{t}-1)\langle\rho_{t}\rangle-\langle\tilde{\rho}_{t}\rangle\right]=0$. The optimal quality factor therefore reads
$\mathcal{Q}^{\mathrm{CTUR}}_{\mathrm{opt}}$ (85)
$=\frac{2\left[a(t)^{2}\mathrm{var}(\rho_{t})-2a(t)b(t)\mathrm{cov}(J_{t},\rho_{t})+b(t)^{2}\mathrm{var}(J_{t})\right]}{\left[\mathrm{var}(J_{t})\mathrm{var}(\rho_{t})-\mathrm{cov}(J_{t},\rho_{t})^{2}\right]\Delta S_{\mathrm{tot}}(t)}.$
The saturated quality factor $\mathcal{Q}_{\mathrm{sat}}^{\mathrm{CTUR}}$ in Eq. (82) is shown in Fig. 7 together with the optimal $\mathcal{Q}^{\mathrm{CTUR}}_{\mathrm{opt}}$ and the quality factors of the current TUR $\mathcal{Q}^{J\mathrm{TUR}}_{(23)}$, density TUR $\mathcal{Q}^{\rho\mathrm{TUR}}_{12}$, and a CTUR $\mathcal{Q}^{\mathrm{CTUR}}_{c=10}$ with $c(t)=10$ for the calmodulin system (see Fig. 6c). We consider the transition weights $\kappa_{ij}^{(23)}=\delta_{i2}\delta_{j3}-\delta_{i3}\delta_{j2}$ and state function $V_{i}=(\delta_{i1}+\delta_{i2})/2$. Only the saturated quality factor remains finite for $t\to\infty$. As expected, the optimal $\mathcal{Q}^{\mathrm{CTUR}}_{\mathrm{opt}}$ gives the best inferred entropy production when $\bm{\kappa}(\tau)$ and $\bm{V}(\tau)$ do not saturate the TUR, whereas $\mathcal{Q}^{\mathrm{CTUR}}_{c=10}\leq\mathcal{Q}^{J\mathrm{TUR}}_{(23)}$ for $t\leq 3$. Hence, simply using an arbitrary $c(t)$ in the CTUR may in fact lead to a worse estimate for the entropy production compared to the current TUR. The non-trivial behavior of $c(t)$ is also highlighted in Fig. 8. The expressions of $c^{*}(t)$ and $c^{\dagger}(t)$ Eqs. (83) can be both positive and negative. Furthermore, they are in no particular relation to another, i.e., one cannot easily say if and when $c^{*}(t)<c^{\dagger}(t)$ and vice versa. A more detailed discussion is given in App. A.10.
Figure 7: Saturation of CTUR in the transient calmodulin folding dynamics (see Fig. 6c). We present the various CTUR quality factors $\mathcal{Q}$ arising from various combinations of density, $\rho$, and current, $J_{t}$, as a function of time. The black line denotes the saturation of the TUR, i.e., when $\bm{\kappa}(\tau)=z_{\tau}$ and $c(t)\rho_{t}=J_{t}^{\mathrm{II}}$, corresponding to the estimated entropy production being the pseudo-entropy production $\Sigma_{\mathrm{est}}=2\langle A_{t}^{2}\rangle$. Above the black line is the region $\Sigma_{\mathrm{est}}>2\langle A_{t}^{2}\rangle$ which is not accessible using the (C)TUR. The remaining quality factors shown use the current ($\mathcal{Q}_{(23)}^{J\mathrm{TUR}}$), density ($\mathcal{Q}_{12}^{\rho\mathrm{TUR}}$), or a combination of these ($\mathcal{Q}_{c=10}^{\mathrm{CTUR}}$ and $\mathcal{Q}_{\mathrm{opt}}^{\mathrm{CTUR}}$) in orange, green, purple, and red, respectively. For the CTURs, the proportionality $c(t)=10$ (purple) and $c(t)=c^{*}(t)$ (red) are chosen, the latter optimizing the quality factor for given $J_{t}$ and $\rho_{t}$. The rates can be found in Tab. 2.
Figure 8: Quality factor as a function of density weighting function $c(t)$ in the calmodulin system at a fixed $t=0.05$. The color correspond to different transition weights $\kappa_{ij}^{(kl)}=\delta_{ik}\delta_{jl}-\delta_{il}\delta_{jk}$ between two states, i.e., $\kappa_{ij}^{(12)}$ (blue), $\kappa_{ij}^{(56)}$ (orange), and $\kappa_{ij}^{(26)}$ (green). The state function entering the density is $V_{i}(\tau)=\delta_{i3}$ and the initial condition is $p_{i}(0)=\delta_{i6}$. The circles are the quality factors for most optimal $c(t)=c^{*}(t)$, while the squares denote the least optimal $c(t)=c^{\dagger}(t)$.
IV.3 Thermodynamic Transport Bound
The thermodynamic bound on the transport [or short “transport bound” (TB)] of a scalar observable $z_{\tau}=z(\bm{x}_{\tau},\tau)$ in continuous space was derived in [51, 151] and reads
$\frac{\langle z_{t}-z_{0}-\int_{0}^{t}\mathrm{d}\tau\partial_{\tau}z_{\tau}\rangle^{2}}{t\mathcal{D}_{t}^{\mathrm{c}}}\leq\frac{\Delta S_{\mathrm{tot}}^{\mathrm{c}}(t)}{2},$ (86)
where $\mathcal{D}_{t}^{\mathrm{c}}$ is the continuous-space fluctuation scale function
$\mathcal{D}_{t}^{\mathrm{c}}\equiv\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\langle[\nabla_{\bm{x}}z_{\tau}]^{T}\bm{\gamma}^{-1}(\tau)\nabla_{\bm{x}}z_{\tau}\rangle,$ (87)
that accounts for how much $z(\bm{x}_{\tau},\tau)$ “stretches” microscopic coordinates $\bm{x}_{\tau}$ [51]. Here, $\bm{\gamma}$ is a positive definite and symmetric friction matrix that is allowed to depend on time and the averages $\langle\cdot\rangle$ in Eqs. (86) and (87) are taken over transient trajectories. In Refs. [51, 113] it was further shown how $\mathcal{D}_{t}^{\mathrm{c}}$ relates to the short-time fluctuations $\mathrm{d}z_{\tau}=z_{\tau+\mathrm{d}\tau}-z_{\tau}$, i.e.,
$\mathcal{D}_{t}^{\mathrm{c}}=\int_{0}^{t}\mathrm{d}\tau\frac{\mathrm{var}(\mathrm{d}z_{\tau})}{2\mathrm{d}\tau}.$ (88)
To prove the discrete-space transport bound we use Eq. (67) and the stochastic integral
$B_{t}=\int_{\tau=0}^{\tau=t}\mathrm{Tr}[\bm{H}(\tau)\mathrm{d}\bm{\varepsilon}(\tau)],$ (89)
where $\bm{H}(\tau)]_{ij}=z_{i}(\tau)-z_{j}(\tau)$ is the difference of some observable state function $z_{\tau}\equiv z(x_{\tau},\tau)=\sum_{k}\delta_{k,x_{\tau}}z_{k}(\tau)$. The mean of $B_{t}$ vanishes while the second moment reads
$\langle B^{2}_{t}\rangle=\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\left[z_{i}(\tau)-z_{j}(\tau)\right]^{2}r_{ij}p_{i}(\tau),$ (90)
which we can identify with the short-time fluctuation function $\mathcal{D}_{t}=\langle B^{2}_{t}\rangle/t$. Namely, consider the infinitesimal difference $\mathrm{d}z(\tau)=z_{\tau+\mathrm{d}\tau}-z_{\tau}$ along the trajectory. Then for any $n\in\mathds{N}_{>0}$ we then get $\langle[\mathrm{d}z(\tau)]^{n}\rangle=\sum_{x,y}(z_{y}-z_{x})^{n}r_{xy}p_{x}(\tau)\mathrm{d}\tau+\mathcal{O}(\mathrm{d}\tau^{2})$. Thus, $\mathrm{var}[\mathrm{d}z(\tau)]=\sum_{x,y}(z_{y}-z_{x})^{2}r_{xy}p_{x}(\tau)\mathrm{d}\tau+\mathcal{O}(\mathrm{d}\tau^{2})$ and
$\mathcal{D}_{t}=\frac{1}{t}\int_{0}^{t}\mathrm{var}[\mathrm{d}z(\tau)]\,.$ (91)
After performing an integration by parts, the cross correlation between $A_{t}$ [given in Eq. (67)] and $B_{t}$ is
$\langle A_{t}B_{t}\rangle$ $=\sum_{i}\left[z_{i}(t)p_{i}(t)-z_{i}(0)p_{i}(0)-\int_{0}^{t}\mathrm{d}\tau\partial_{\tau}z_{i}(\tau)p_{i}(\tau)\right]$
$=\langle z_{t}-z_{0}-\int_{0}^{t}\mathrm{d}\tau\partial_{\tau}z_{\tau}\rangle.$ (92)
The average of $z_{t}$ (and of any vector) is given by $\langle z_{\tau}\rangle=\sum_{k}\sum_{l}\delta_{k,l}z_{k}(\tau)p_{l}(t)$. By combining the various averages entering the Cauchy-Schwarz inequality $\langle A_{t}B_{t}\rangle\leq\langle A_{t}^{2}\rangle\langle B_{t}^{2}\rangle$ we obtain the transport bound for MJP
$\frac{\langle z_{t}-z_{0}-\int_{0}^{t}\mathrm{d}\tau\partial_{\tau}z_{\tau}\rangle^{2}}{t\mathcal{D}_{t}}\leq\frac{\Delta S_{\mathrm{tot}}(t)}{2}.$ (93)
Upon identifying the dynamical activity $\int_{0}^{t}\mathrm{d}\tau\sum_{x,y\neq x}r_{xy}p_{x}(\tau)$ [111, 112], the fluctuation-scale function $\mathcal{D}_{t}$ can be seen as a $z_{y}(\tau)-z_{x}(\tau)]^{2}-$weighted dynamical activity, which justifies referring to $\mathcal{D}_{t}$ as the diffusion constant of $z_{\tau}$ on a graph.
IV.3.1 Saturation of the Transport Bound
While the continuous TB in Eq. (86) can be saturated, however not uniquely, with $\nabla_{\bm{x}}z_{\tau}=c\bm{\gamma}\bm{j}(\bm{x},\tau)/P(\bm{x},\tau)$ (see Ref. [51] for details), there is no general way to saturate the discrete TB. We provide a simple counterexample to show why this is the case. Consider some $a,b,c>0$ and let
$\bm{L}=\begin{pmatrix}-1-a&b&1\\ 1&-1-b&c\\ a&1&-1-c\end{pmatrix}$, (94)
be the generator for a fully connected three state MJP with states $\{1,2,3\}$. Saturation of Eq. (93) entails choosing $z_{\tau}$ such that $\bm{H}(\tau)^{T}\propto z_{\tau}$ to saturate the Cauchy-Schwarz inequality. Explicitly, for all $\tau$
$z_{2}(\tau)-z_{1}(\tau)$ $\overset{!}{=}Z_{12}(\tau)=\frac{p_{1}(\tau)-bp_{2}(\tau)}{p_{1}(\tau)+bp_{2}(\tau)}\,$,
$z_{3}(\tau)-z_{1}(\tau)$ $\overset{!}{=}Z_{13}(\tau)=\frac{ap_{1}(\tau)-p_{3}(\tau)}{ap_{1}(\tau)+p_{3}(\tau)}\,$,
$z_{3}(\tau)-z_{2}(\tau)$ $\overset{!}{=}Z_{23}(\tau)=\frac{p_{2}(\tau)-cp_{3}(\tau)}{p_{2}(\tau)+cp_{3}(\tau)}\,.$ (95)
A solution of Eqs. (95) involves $Z_{23}(\tau)=Z_{13}(\tau)-Z_{12}(\tau)$, i.e.,
$\frac{p_{2}(\tau)-cp_{3}(\tau)}{p_{2}(\tau)+cp_{3}(\tau)}=\frac{2p_{1}(\tau)(abp_{2}(\tau)-p_{3}(\tau))}{(ap_{1}(\tau)+p_{3}(\tau))(p_{1}(\tau)+bp_{2}(\tau))}.$ (96)
Even for $abc=1$, which corresponds to a system relaxing towards equilibrium, we can always choose an initial condition $\bm{p}(\tau=0)$ such that the equality in Eq. (96) is not achieved. For example, let $a=b=c=2$ and $p_{2}(0)=p_{3}(0)=1/2$. The l.h.s. of Eq. (96) yields 1/3-1/3 while the r.h.s. is 0, hence the equality is not achieved and the set of equations Eqs. (95) has no solution. Thus, there is no state function $z_{\tau}$ which saturates the TB.
IV.3.2 Unification of (C)TUR and TB
It turns out that the TB is a special case of the CTUR. Starting from the transient CTUR (80), we now choose $\rho_{t}$ such that $J_{t}-c(t)\rho_{t}=J^{\mathrm{I}}_{t}$, i.e., $V_{i}(\tau)=\sum_{j}\kappa_{ij}(\tau)r_{ij}/c(t)$. Then,
$t\partial_{t}\langle\rho_{t}\rangle$ $=t\sum_{i}V_{i}(t)p_{i}(t)$
$=t\sum_{i,j\neq i}\frac{\kappa_{ij}(\tau)r_{ij}}{c(t)}p_{i}(t)$
$=\frac{1}{c(t)}t\partial_{t}\langle J_{t}\rangle,$ (97)
and similarly
$\langle\tilde{\rho}_{t}\rangle$ $=\int_{0}^{t}\mathrm{d}\tau\sum_{i}\tau\partial_{\tau}V_{i}(\tau)p_{i}(\tau)$
$=\int_{0}^{t}\mathrm{d}\tau\sum_{i,j\neq i}\frac{\tau\partial_{\tau}\kappa_{ij}(\tau)r_{ij}}{c(t)}p_{i}(\tau)$
$=\frac{1}{c(t)}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j\neq i}\tau\partial_{\tau}\kappa_{ij}(\tau)r_{ij}p_{i}(\tau)$
$=\frac{1}{c(t)}\langle\tilde{J}_{t}\rangle.$ (98)
An analogous calculation reveals $\langle\rho_{t}\rangle=\langle J_{t}\rangle/c(t)$ and hence using this choice of $\rho_{t}$ in Eq. (80) yields the following form of the transient CTUR
$\frac{\langle J_{t}\rangle^{2}}{\mathrm{var}(J_{t}-c(t)\rho_{t})}=\frac{\langle J_{t}\rangle^{2}}{\mathrm{var}(J_{t}^{\mathrm{I}})}\leq\frac{\Delta S_{\mathrm{tot}}(t)}{2}\,.$ (99)
We call Eq. (99) the generalized TB (gTB) because the choice $\kappa_{ij}(\tau)=z_{j}(\tau)-z_{i}(\tau)$ yields the transport bound. Explicitly, the mean current is
$\langle J_{t}\rangle$ $=\int_{0}^{t}\mathrm{d}s\sum_{i,j\neq i}[z_{j}(s)-z_{i}(s)]r_{ij}p_{i}(s)$
$=\int_{0}^{t}\mathrm{d}s\sum_{i}z_{i}(s)\underbrace{\sum_{j\neq i}[r_{ji}p_{j}(s)-r_{ij}p_{i}(s)]}_{=\partial_{s}p_{i}(s)}$ (100)
$=\sum_{i}\left(z_{i}(t)p_{i}(t)-z_{i}(0)p_{i}(0)-\int_{0}^{t}\mathrm{d}s[\partial_{s}z_{i}(s)]p_{i}(s)\right),$
and the variance of the dissipative current reads
$\mathrm{var}(J_{t}^{\mathrm{I}})$ $=\langle(J_{t}^{\mathrm{I}})^{2}\rangle-\underbrace{\langle J_{t}^{\mathrm{I}}\rangle^{2}}_{=0}$
$=\int_{0}^{t}\mathrm{d}s\sum_{i,j\neq i}[z_{j}(s)-z_{i}(s)]^{2}r_{ij}p_{i}(s)$
$=t\mathcal{D}_{t}.$ (101)
With the specific choice $\kappa_{ij}(\tau)=z_{j}(\tau)-z_{i}(\tau)$, i.e., using Eqs. (100) and (101), Eq. (99) becomes the TB
$\frac{\langle z_{t}-z_{0}-\int_{0}^{t}\mathrm{d}\tau\partial_{\tau}z_{\tau}\rangle^{2}}{t\mathcal{D}_{t}}\leq\frac{\Delta S_{\mathrm{tot}}(t)}{2}.$ (102)
Equation (99) therefore naturally relates the TB to the CTUR as a special case. Indeed, the gTB is efficient, i.e. yields a non-zero bound in both, transient and stationary systems, while the TB only yields nontrivial bounds in transient systems. A similar connection can be made in continuous space, where a recent TUR for underdamped Langevin dynamics [126] can be seen as a generalization of the continuous TB [51]. It should be noted that, while the TB in general cannot be saturated, the gTB can be saturated with the choice of weight $\bm{\kappa}(\tau)\propto z_{\tau}$. Moreover, in some cases the transient TB of one system may be identical to the stationary gTB of another system. The following example, an extension of the SAT model (see Fig. 6b), shows how the TB may give rise to the gTB in a special case.
Figure 9: Comparison of TUR and transport bound with the periodic extension of the SAT model in Fig. 6b. The mean of ${J}_{t}$ (orange) and $z_{x,n}$ (blue) from $N=50000$ trajectories sampled with the Gillespie algorithm is shown in (a) as a function of $t$, compared to the analytic $\langle{J}_{t}\rangle$ [see Eq. (73)], for $e^{\mathrm{out}}_{B}\approx 40$. The inset shows the fraction $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}$ for different values of $e^{\mathrm{out}}_{B}$. The dashed orange and red lines indicate the values of $e^{\mathrm{out}}_{B}$ where the antiporter regime begins ($e^{\mathrm{out}}_{B}\approx 27$) and where $\mathcal{A}_{\mathcal{C}_{3}}=0$ ($e^{\mathrm{out}}_{B}=6$), respectively, while the black line is the upper limit $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}$ can have so that the system is an antiporter. The mean current is shown in (b) for various values of $e^{\mathrm{out}}_{B}$. The steady-state quality factors for TUR (solid) and TB (dashed) are shown in (c) as a function of $t$ for various values of $e^{\mathrm{out}}_{B}$ in the antiporter regime. The parameters are listed in Tab. 3.
The SAT describes the transport of a large molecule (denoted molecule $A$) through a cell membrane against its chemical gradient using the transport of helper molecule (denoted molecule $B$) [145, 146, 147], see Fig. 6a-b. The reduction to a six-state Markov model, $\mathcal{N}=\{1,2,3,4,5,6\}$, is commonly used to describe this secondary active transport [145, 146, 147]. The efficiency of the process is related to the number of molecules $A$ transported given a concentration of $B$. The transport can be split into two categories: symporters and antiporters. These refer to molecule $A$ being transported parallel or anti-parallel to molecule $B$, respectively. To effectively describe the total number of transported molecules, we extend the six-state model to a chain of six-state models, i.e., $\tilde{\mathcal{N}}=\mathcal{N}\times\mathds{Z}$, where the additional degree of freedom corresponds to the number of $A$ molecules transported through the membrane. We are interested in the fraction of the total steady-state entropy production that we can estimate via the thermodynamic bounds for various values of transition rates ${\rm e}^{\mathrm{out}}_{B}$. These transition rates are assumed proportional to the concentration of $B$ outside the cell [145].
To make reasonable comparisons, we either have to consider symporter or antiporter systems. We are going to focus on the latter. We first vary ${\rm e}^{\mathrm{out}}_{B}$ to assess which values of transition rates give rise to antiporter behaviour before adressing thermodynamic inferrence.
We start with the cycle affinities of the model. The six-state SAT model consist of three cycles $\mathcal{C}_{1}=\{1\to 2,2\to 4,4\to 3,3\to 1\}$, $\mathcal{C}_{2}=\{3\to 4,4\to 6,6\to 5,5\to 3\}$, and $\mathcal{C}_{3}=\{1\to 2,2\to 4,4\to 6,6\to 5,5\to 3,3\to 1\}$. The corresponding cycle affinities are $\mathcal{A}_{\mathcal{C}_{1}}=\log(e^{\mathrm{out}}_{A}/e^{\mathrm{in}}_{A})$, $\mathcal{A}_{\mathcal{C}_{2}}=\log(e^{\mathrm{in}}_{B}/e^{\mathrm{out}}_{B})$, and $\mathcal{A}_{\mathcal{C}_{3}}=\mathcal{A}_{\mathcal{C}_{1}}+\mathcal{A}_{\mathcal{C}_{2}}=\log(e^{\mathrm{out}}_{A}e^{\mathrm{in}}_{B}/e^{\mathrm{in}}_{A}e^{\mathrm{out}}_{B})$. Since the transport of molecule $A$ with help of $B$ is of interest, one could naively assume that the transport relates to $\mathcal{A}_{\mathcal{C}_{3}}$. However, $\mathcal{A}_{\mathcal{C}_{3}}$ does not quantify the direction of transport, as $\mathcal{A}_{\mathcal{C}_{3}}$ does not give any information on the transport of type $A$ alone. One can also convince oneself that $\mathcal{A}_{\mathcal{C}_{1}}$ is insufficient as well, due to the cycle affinity being independent of $e^{\mathrm{out}}_{B}$.
However, the current (entering the stationary TUR) in the six-state model gives the insight we require. The relevant transition weight is $\kappa_{ij}^{\mathrm{SAT}}=\delta_{i1}\delta_{j3}-\delta_{i3}\delta_{j1}$, effectively counting the total number of molecules $A$ being transported towards the exterior of the cell. The mean current is $\langle J_{t}^{\mathrm{SAT}}\rangle_{\mathrm{s}}=\sum_{x,y}\kappa_{xy}^{\mathrm{SAT}}r_{xy}p_{x}^{\mathrm{s}}$. If $\langle J_{t}^{\mathrm{SAT}}\rangle_{\mathrm{s}}>0$, the average total number of molecules $A$ transported towards the outside of the cell is positive, corresponding to an average transport of $A$ from the inside to the outside of the cell. Thus, the condition for an antiporter behavior is $l_{A}/e^{\mathrm{out}}_{A}>p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}$, which depends on ${\rm e}^{\mathrm{out}}_{B}$ via the steady-state distribution. With the parameters we use, see Tab. 3, we need $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}<2/3$. The value $2/3$ is marked as a black line in the inset of Fig. 9a, where the green line corresponds to the value of $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}$ as a function of ${\rm e}^{\mathrm{out}}_{B}$. The dashed red line marks the value of ${\rm e}^{\mathrm{out}}_{B}$ s.t. $\mathcal{A}_{\mathcal{C}_{3}}=0$, while the dashed orange line is where $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}=2/3$. The same can be readily seen in Fig. 9b, where the mean current is presented for various values of ${\rm e}^{\mathrm{out}}_{B}$. Clearly, the condition $\mathcal{A}_{\mathcal{C}_{3}}=0$ is not sufficient for a positive current.
Additionally, we can consider the transient extended model and apply the TB. For the extended model, consider $p_{x,n}(\tau)$ the probability of being in state $x$ in system $n$ at time $\tau$, where $n\in\mathds{Z}$ corresponds to the number of $A$ molecules transported through the membrane. When marginalizing over $\mathds{Z}$ we recover the steady-state probability $p_{x}^{\mathrm{s}}=\sum_{n\in\mathds{Z}}p_{x,n}(\tau)$. As a state function we choose the number of $A$ molecules $z_{\tau}=n$, such that
$\langle z_{\tau}\rangle=\sum_{n\in\mathds{Z}}n\sum_{x=1}^{6}p_{x,n}(\tau)=\sum_{n\in\mathds{Z}}n\tilde{p}_{n}(\tau),$ (103)
where $\tilde{p}_{n}(\tau)$ is the probability distribution marginalized over $\mathcal{N}$. Without loss of generality, we use $\langle z_{0}\rangle=0$. Equation (103) therefore describes the average number of net-transitions $\langle n_{13}-n_{31}\rangle$ between states 1 and 3 in the original system until time $\tau$, as $z_{\tau}$ the net number of transitions. Hence, one can immediately recognize that $\langle z_{\tau}\rangle=\langle J_{t}^{\mathrm{SAT}}\rangle_{\mathrm{s}}$, which can be readily seen in Fig. 9a.
The fluctuation-scale function is recognized to be
$\mathcal{D}_{t}^{\mathrm{ext}}=\frac{1}{t}\sum_{n\in\mathds{Z}}\sum_{x,y\in\mathcal{N}}\int_{0}^{t}\mathrm{d}\tau$ $\left[(z_{y,n}-z_{x,n})^{2}\delta_{x\to y,n}+(z_{y,n+1}-z_{x,n})^{2}\delta_{x\to y,n\to n+1}\right$.
$+\left.(z_{y,n-1}-z_{x,n})^{2}\delta_{x\to y,n\to n-1}\right]r_{xy}p_{x,n}(\tau)\,.$ (104)
Note that $\delta_{x\to y,n\to n+1}$ is one if the transition $x\to y$ is observed while the system goes from $n$ to $n+1$ and zero otherwise. We can simplify Eq. (104) to find
$\mathcal{D}_{t}^{\mathrm{ext}}=\frac{1}{t}\sum_{n\in\mathds{Z}}\int_{0}^{t}\mathrm{d}\tau\left[r_{13}p_{1,n}(\tau)+r_{31}p_{3,n}(\tau)\right],$ (105)
because $z_{x,n}-z_{y,n}=0$ and $z_{x,n}-z_{y,n\pm 1}=\pm 1$. Marginalizing over $n$ finally gives
$\mathcal{D}_{t}^{\mathrm{ext}}$ $=\left[r_{13}p_{1}^{\mathrm{s}}+r_{31}p_{3}^{\mathrm{s}}\right]=\sum_{x,y\in\mathcal{N}}(\kappa_{xy}^{\mathrm{SAT}})^{2}r_{xy}p_{x}^{\mathrm{s}}.$ (106)
Upon comparing Eq. (106) to Eq. (60) it becomes apparent that $\mathcal{D}_{t}^{\mathrm{ext}}=\lim_{t\to 0}\mathrm{var}_{\mathrm{s}}(J_{t}^{\mathrm{SAT}})/t$. Furthermore, using the TB on the transient extended SAT model yields the stationary gTB of the original six-state SAT model, as the TB reduces to Eq. (99).
The quality factors of the TUR $\mathcal{Q}^{\mathrm{TUR}}$ and TB $\mathcal{Q}^{\mathrm{TB}}$ are shown in Fig. 9c (solid and dashed line, respectively) for ${\rm e}^{\mathrm{out}}_{B}$ in the antiporter regime. Because $t\mathcal{D}_{t}^{\mathrm{ext}}>\mathrm{var}_{\mathrm{s}}(J_{t}^{\mathrm{SAT}})$ for all $t>0$, the TUR infers a higher fraction of the entropy production, i.e., $\mathcal{Q}^{\mathrm{TUR}}\geq\mathcal{Q}^{\mathrm{TB}}$. Additionally, one may show that taking ${\rm e}^{\mathrm{out}}_{B}\to\infty$ leads to the quality factors converging to a finite value [12, 152]. Hence, there is no improvement in the inference of dissipation at higher concentrations of $B$ outside the cell beyond a certain point.
IV.4 Thermodynamic Correlation Bound
Thermodynamic bounds on correlation times were studied in the context of Langevin dynamics in [113] in stationary systems at large times. Recently such thermodynamic correlation inequalities were extended to finite times and transients [114]. Following the latter stochastic-calculus approach, we here develop analogous results for time-homogeneous MJP, in the general transient setting and for arbitrary times. That such results are possible was already anticipated in [113].
The proof of the correlation bound starts from the auxiliary stochastic integral
${B}_{t}$ $=\frac{1}{\sqrt{t}}\int_{\tau=0}^{\tau=t}\mathrm{Tr}\left[\left(\bm{H}(\tau)-\tilde{\bm{Z}}(\tau)^{T}\right)\mathrm{d}\bm{\varepsilon}(\tau)\right],$ (107)
where $\tilde{Z}_{xy}(\tau)=Z_{xy}(\tau)\left[z_{x}(\tau)+z_{y}(\tau)-2F(\tau)\right]$ for some time-dependent function $F(\tau)$. The second key ingredient is the shifted and rescaled density observable
${C}_{t}$ $=\frac{1}{\sqrt{t}}\int_{\tau=0}^{\tau=t}\left[\bm{V}-\langle{V}_{\tau}\rangle\bm{1}\right]^{T}\mathrm{d}\bm{\tau}(\tau),$ (108)
where $\bm{1}$ is the one-vector [i.e. $(\bm{1})_{i}=1$ for all $i\in\mathcal{N}$] and $\langle V_{\tau}\rangle=\sum_{i}V_{i}p_{i}(\tau)$. While the TUR and TB allow the state functions $z_{\tau}$ and $\bm{V}(\tau)$ to be explicitly time-dependent, the correlation bound requires these to be strictly time-independent. Furthermore, we write, e.g., $\langle z_{\tau}\rangle=\sum_{i}z_{i}p_{i}(\tau)$ to explicitly highlight that the average depends on time through $p_{i}(\tau)$. Additionally, we need to introduce the re-weighted two-point probability
$p_{xy}^{\mathrm{ps}}(\tau)=\frac{p_{x}(\tau)r_{xy}Z_{xy}(\tau)^{2}}{\Sigma_{\mathrm{ps}}}\,$, (109)
with the renormalization constant $\Sigma_{\mathrm{ps}}=\sum_{x,y}p_{x}(\tau)r_{xy}Z_{xy}(\tau)^{2}$ which is the pseudo-entropy production rate (pseudo-EPR). Expectation values w.r.t. $p_{xy}^{\mathrm{ps}}(\tau)$ will be denoted by $\langle\cdot\rangle_{\mathrm{ps}}$. We will also introduce the pseudo variance $\mathrm{pvar}_{\mathrm{ps}}^{F}(z_{\tau})=\sum_{x,y}(z_{x}+z_{y}-2F(\tau))^{2}p_{xy}^{\mathrm{ps}}(\tau)$, which is the “true” variance only for $F(\tau)=\langle z_{\tau}\rangle_{\mathrm{ps}}/2=\sum_{x,y}(z_{x}+z_{y})p_{xy}^{\mathrm{ps}}(\tau)/2$ and some other positive time-dependent quantity otherwise.
We immediately see that
$\langle C_{t}^{2}\rangle=\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\int_{0}^{t}\mathrm{d}\tau^{\prime}\mathrm{cov}\left({V}_{\tau},{V}_{\tau^{\prime}}\right)\,$, (110)
where $V_{\tau}\equiv V(x_{\tau})=\sum_{k}\delta_{k,x_{\tau}}V_{k}$, and introducing the time-averaged density
$\overline{V}_{t}$ $\equiv\frac{1}{t}\int_{\tau=0}^{\tau=t}V_{\tau}\mathrm{d}\tau\,$,
$\overline{z}_{t}$ $\equiv\frac{1}{t}\int_{\tau=0}^{\tau=t}z_{\tau}\mathrm{d}\tau,$ (111)
Eq. (110) can be identified as
$t\mathrm{var}(\overline{V}_{t})=\langle C_{t}^{2}\rangle.$ (112)
With the shorthand notations $\bm{V}^{\Delta}(\tau)=\bm{V}-\langle{V}(\tau)\rangle\bm{1}$ and $z^{F}_{\tau}=z_{\tau}-\langle F(\tau)\rangle\bm{1}$, the cross term is computed to be
$t\langle B_{t}C_{t}\rangle$ $=\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{x,y,i}\mathds{1}_{\tau<\tau^{\prime}}V^{\Delta}_{i}(\tau^{\prime})\left(H_{yx}-Z_{xy}(\tau)\left[z_{x}^{F}(\tau)+z_{y}^{F}(\tau)\right]r_{xy}p_{x}(\tau)\right)\left[P(i,\tau^{\prime}|y,\tau)-P(i,\tau^{\prime}|x,\tau)\right],$ (113)
where the sum over $x,y$ can be simplified because of symmetries (we suppress the time indices here) to get
$\sum_{x,y}H_{yx}r_{xy}p_{x}[P(i|y)-P(i|x)]$
$=-\sum_{x,y}z_{x}^{F}(r_{xy}p_{x}+r_{yx}p_{y})[P(i|y)-P(i|x)],$ (114)
and
$\sum_{x,y}Z_{xy}(z_{x}^{F}+z_{y}^{F})r_{xy}p_{x}\left[P(i|y)-P(i|x)\right]$
$=\sum_{x,y}z_{x}^{F}(r_{xy}p_{x}-r_{yx}p_{y})\left[P(i|y)-P(i|x)\right].$ (115)
Using time-homogeneity of the semi-group $\bm{P}(\tau^{\prime}|\tau)=\bm{P}(\tau^{\prime}-\tau)\equiv{\rm e}^{\bm{L}(\tau^{\prime}-\tau)}$ which obeys the master equation Eq. (32), i.e. $\partial_{\tau}\bm{P}(\tau)=\bm{L}\bm{P}(\tau)=\bm{P}(\tau)\bm{L}$, and $\partial_{\tau}\bm{P}(\tau^{\prime}-\tau)=-\partial_{\tau^{\prime}}\bm{P}(\tau^{\prime}-\tau)$, the integrand in Eq. (113) simplifies to $2\sum_{x,i}\mathds{1}_{\tau<\tau^{\prime}}V^{\Delta}_{i}(\tau^{\prime})z^{F}_{x}(\tau)p_{x}(\tau)\partial_{\tau^{\prime}}P(i,\tau^{\prime}|x,\tau)$. Since
$\sum_{x,i}\left[\partial_{\tau^{\prime}}V^{\Delta}_{i}(\tau^{\prime})\right]z^{F}_{x}(\tau)p_{x}(\tau)P(i,\tau^{\prime}|x,\tau)$
$=-\partial_{\tau^{\prime}}\langle{V}_{\tau^{\prime}}\rangle\underbrace{\sum_{x}z^{F}_{x}(\tau)p_{x}(\tau)\underbrace{\sum_{i}P(i,\tau^{\prime}|x,\tau)}_{=1}}_{=\langle z_{\tau}\rangle-F(\tau)},$ (116)
the particular choice $F(\tau)=\langle z_{\tau}\rangle$ yields
$t\langle B_{t}C_{t}\rangle$ $=2\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\mathds{1}_{\tau<\tau^{\prime}}\partial_{\tau^{\prime}}\mathrm{cov}({V}_{\tau^{\prime}},z_{\tau}).$ (117)
An integration by parts in $\tau^{\prime}$ with $\partial_{\tau^{\prime}}\mathds{1}_{\tau<\tau^{\prime}}=\delta(\tau^{\prime}-\tau)$ further yields
$\langle B_{t}C_{t}\rangle$ $=2\mathrm{cov}\left[{V}_{t},\overline{{z}}_{t}\right]-\frac{2}{t}\int_{0}^{t}\mathrm{d}\tau\mathrm{cov}\left[{V}_{\tau},z_{\tau}\right].$ (118)
Lastly, the second moment of $B_{t}$ reads
$\langle B_{t}^{2}\rangle=$ $\mathcal{D}_{t}+\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\Sigma^{\mathrm{ps}}(\tau)\mathrm{pvar}_{\mathrm{ps}}^{F}\left(z_{\tau}\right)$
$-\frac{2}{t}\int_{0}^{t}\mathrm{d}\tau\sum_{x,y}\left(\left[z_{y}-F(\tau)\right]^{2}-\left[z_{x}-F(\tau)\right]^{2}\right)$
$\times Z_{xy}(\tau)r_{xy}p_{x}(\tau)\,.$ (119)
The integrand of the last term in Eq. (119) reduces to
$\sum_{x,y}\left(\left[z_{y}-F(\tau)\right]^{2}-\left[z_{x}-F(\tau)\right]^{2}\right)Z_{xy}(\tau)r_{xy}p_{x}(\tau)$
$=\sum_{y}\left[z_{y}-F(\tau)\right]^{2}\partial_{\tau}p_{y}(\tau)\,.$ (120)
Choosing $F(\tau)=\langle z_{\tau}\rangle$ again simplifies Eq. (119) to
$\langle B_{t}^{2}\rangle=$ $\mathcal{D}_{t}+\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\Sigma^{\mathrm{ps}}(\tau)\mathrm{pvar}_{\mathrm{ps}}^{F}\left(z_{\tau}\right)$
$-\frac{2}{t}\left[\mathrm{var}\left(z_{t}\right)-\mathrm{var}\left(z_{0}\right)\right]\,.$ (121)
Combining Eqs. (112), (118), and (121) yields the transient finite-time correlation bound
$\frac{\left[2\mathrm{cov}\left({V}_{t},\overline{{z}}_{t}\right)-\frac{2}{t}\int_{0}^{t}\mathrm{d}\tau\mathrm{cov}\left({V}_{\tau},z_{\tau}\right)\right]^{2}}{t\mathrm{var}(\overline{V}_{t})}-\mathcal{D}_{t}+\frac{2}{t}\left[\mathrm{var}\left(z_{t}\right)-\mathrm{var}\left(z_{0}\right)\right]\leq\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\Sigma^{\mathrm{ps}}(\tau)\mathrm{pvar}_{\mathrm{ps}}^{F}\left(z_{\tau}\right).$ (122)
In NESS, Eq. (122) simplifies to [we suppress time indices for stationary single-time (pseudo/co-)variances]
$\frac{4\left[\mathrm{cov}({V}_{t},\overline{{z}}_{t})-\mathrm{cov}_{\mathrm{s}}({V},{z})\right]^{2}}{t\mathrm{var}(\overline{V}_{t})}-\mathcal{D}_{t}$
$\leq\frac{\langle A_{t}^{2}\rangle_{\mathrm{s}}}{t}\mathrm{pvar}_{\mathrm{ps}}^{F}({z})\leq\frac{\Delta S_{\mathrm{tot}}(t)}{2t}\mathrm{pvar}_{\mathrm{ps}}^{F}({z}).$ (123)
The two point probability density $p_{xy}(\tau)$ is not accessible, since it requires knowing $\bm{L}$, hence the pseudo variance in Eqs. (122) and (123) needs to be further bounded. In the former case, it was assumed that $F(\tau)=\langle z_{\tau}\rangle$, while in the stationary case the choice of $F(\tau)=F$ is free. Since we consider bounded $z_{\tau}$ on finite state spaces, there exist bounds $z_{\max/\min}$ such that for all $x\neq y$  [^2] It should be clear that $p_{xx}(\tau)=0$ for all $x$ and $\tau$, see Eq. (109). the function entering the pseudo variance satisfies $z_{\min}\leq z_{x}+z_{y}\leq z_{\max}$. In steady state with $F=\langle z_{\tau}\rangle_{\mathrm{ps}}/2$, $\mathrm{pvar}_{\mathrm{ps}}^{F}({z})=\mathrm{var}_{\mathrm{ps}}({z})\leq(z_{\max}-z_{\min})^{2}/4$, where we used the Popoviciu’s inequality.
It should also be clear from Eqs. (122) and (123) that the quality factor for the CB, $\mathcal{Q}^{\mathrm{CB}}$, can become negative (and hence the bound is uninformative). To be precise, let
$G(t)\equiv\frac{\left[2\mathrm{cov}\left({V}_{t},\overline{{z}}_{t}\right)-\frac{2}{t}\int_{0}^{t}\mathrm{d}\tau\mathrm{cov}\left({V}_{\tau},z_{\tau}\right)\right]^{2}}{t\mathrm{var}(\overline{V}_{t})},$ (124)
then clearly $G(t),\mathcal{D}_{t},\mathrm{var}\left(z_{t}\right)\geq 0$ for all $t\geq 0$. Therefore, if $G(t)-\mathcal{D}_{t}+\frac{2}{t}\left[\mathrm{var}\left(z_{t}\right)-\mathrm{var}\left(z_{0}\right)\right]<0$, the inferred entropy production is negative. The fact that the quality factor may become negative, giving a lower bound on the total entropy production that is worse than the second law of thermodynamics, is not particularly appealing. However, the CB can be applied in situations where many established methods fail to infer entropy production. In the following we provide a simple four-state example where CB can be applied successfully but other methods fail.
Example II: 4-State Toy model.—Consider a four-state model $\{1,2,3,4\}$ on a ring, see Fig. 6e. The rates clockwise are unity and the counterclockwise rates are controlled by the parameter $\epsilon$. The steady-state distribution $p_{i}^{\mathrm{s}}=p=1/4$ is independent of $\epsilon$. Suppose that we can only distinguish between a pair of mesostates A and B it can only be distinguished whether the states are in either states 1 and 2 or states 3 and 4. These we refer to mesostates A (containing $\{1,2\}$) and B (containing $\{3,4\}$), see Fig. 6e.
Inferring the total steady-state entropy production $\Delta S_{\mathrm{tot}}(t)=t(\epsilon-1)\log\epsilon$ by only observing transitions between states A and B is a challenge where surprisingly few bounds provide non-trivial estimates. For example, the TUR (only) gives the second law, as
$\langle J_{t}\rangle=p\left(\underbrace{\epsilon\kappa_{\mathrm{AB}}}_{2\to 3}+\underbrace{\kappa_{\mathrm{BA}}}_{3\to 2}+\underbrace{\epsilon\kappa_{\mathrm{BA}}}_{4\to 1}+\underbrace{\kappa_{\mathrm{AB}}}_{1\to 4}\right)=0\,.$ (125)
Including a density, i.e., using the CTUR Eq. (80), yields the same bound, because the density only contributes in transient dynamics. Similarly, the TB is only non-trivial for transient systems. Using the gTB is not profitable either, due to Eq. (125).
However, the CB can be applied to yield a positive bound on the EP for sufficiently large $t$ and $\epsilon$. The fluctuation scale function is $\mathcal{D}_{t}=(1+\epsilon)/2$ and for $\mathrm{X}=\mathrm{A},\mathrm{B}$ the steady-state (co)variance is $\mathrm{cov}_{\mathrm{s}}(V^{\mathrm{A}},V^{\mathrm{X}})=(\delta_{\mathrm{X},\mathrm{A}}-\delta_{\mathrm{X},\mathrm{B}})/4$. If $\lambda_{k}$, $k\in\{1,2,3,4\}$, are the eigenvalues of the generator
$\lambda_{1}$ $=0,$
$\lambda_{2}$ $=-2(\epsilon+1),$
$\lambda_{3/4}$ $=-1-\epsilon\pm(\epsilon-1)i,$ (126)
then for times $t\gg\max_{k\in\{2,3,4\}}\left(-1/\mathrm{Re}(\lambda_{k})\right)=1/(1+\epsilon)$ we can use
$t\mathrm{var}(\overline{V}_{t}^{\mathrm{A}})=2\int_{0}^{t}\mathrm{d}\tau\mathrm{cov}(V^{\mathrm{A}}_{\tau},V^{\mathrm{A}}_{0})+\mathcal{O}(t^{-1}),$ (127)
and is therefore similar to
$\mathrm{cov}(V^{\mathrm{A}}_{t},\overline{V}^{\mathrm{B}}_{t})$ $=-\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\mathrm{cov}(V^{\mathrm{A}}_{t},V^{\mathrm{A}}_{\tau})$
$=-\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\mathrm{cov}(V^{\mathrm{A}}_{\tau},V^{\mathrm{A}}_{0}).$ (128)
Thus, the l.h.s. of Eq. (123) for $t\gg 1/(1+\epsilon)$ reads
$\frac{4\left[\mathrm{cov}({V}_{t},\overline{{z}}_{t})-\mathrm{cov}_{\mathrm{s}}({V},{z})\right]^{2}}{t\mathrm{var}(\overline{V}_{t})}-\mathcal{D}_{t}$
$=\frac{1}{8\int_{0}^{t}\mathrm{d}\tau\mathrm{cov}(V^{\mathrm{A}}_{\tau},V^{\mathrm{A}}_{0})}-\frac{1+\epsilon}{2}+\mathcal{O}(t^{-1})$ (129)
The first term can be simplified further to be $(\epsilon^{2}+1)/(\epsilon+1)$ (see App. A.6), so that the large $t$ quality factor becomes
$\mathcal{Q}$ $\xrightarrow{t\to\infty}\frac{1}{\log\epsilon}\left[\frac{(\epsilon+1)^{2}+(\epsilon-1)^{2}}{(\epsilon^{2}-1)}-\frac{\epsilon+1}{\epsilon-1}\right]\frac{1}{\mathrm{pvar}_{\mathrm{ps}}^{F}(V^{\mathrm{A}})}$
$=\frac{(\epsilon-1)}{(\epsilon+1)\log\epsilon}\frac{1}{\mathrm{pvar}_{\mathrm{ps}}^{F}(V^{\mathrm{A}})}\,.$ (130)
Recall that the choice of $F(\tau)$ is free for stationary systems, hence we can choose $F(\tau)=\langle V^{\mathrm{A}}_{x}+V^{\mathrm{A}}_{y}\rangle_{\mathrm{ps}}/2$ s.t. $\mathrm{pvar}_{\mathrm{ps}}^{F}(V^{\mathrm{A}})=\mathrm{var}_{\mathrm{ps}}(V^{\mathrm{A}})$. Using Popoviciu’s inequality, the pseudo variance (i.e., variance) with the choice $F(\tau)=\langle V^{\mathrm{A}}_{x}+V^{\mathrm{A}}_{y}\rangle_{\mathrm{ps}}/2$ (i.e., the two-point variance of $V^{\mathrm{A}}_{x}+V^{\mathrm{A}}_{y}$) can be bounded by $\mathrm{var}_{\mathrm{ps}}^{F}(V^{\mathrm{A}})=1/2\leq 1$ (see Appendix A.6). Thus, the estimated entropy production using Popoviciu’s inequality is half of that estimated with the exact pseudo variance, resulting in $2\mathcal{Q}^{\mathrm{Pop}}=\mathcal{Q}$, which is shown in Fig. 10a, where the exact EPR is shown together with the long-time limit of estimated EPR with and without Popoviciu’s inequality. The inset in Fig. 10a shows the corresponding quality factors Eq. (130) in the limit $t\to\infty$. As a sanity check, theoretical and simulation results obtained with the celebrated Gillespie algorithm [143, 144]) of $G(t)-\mathcal{D}_{t}$ are compared in Fig. 10b, where a good agreement is found. In addition, the leading order correction is shown in Fig. 10c by considering
$\delta(G(t)-\mathcal{D}_{t})=G(t)-\mathcal{D}_{t}-\frac{(\epsilon-1)^{2}}{2(\epsilon+1)}\,$, (131)
i.e., subtracting the $t\to\infty$ limit. As expected, the leading order correction decays as $t^{-1}$.
Figure 10: Correlation bound applied to the driven ring, see Fig. 6e. In (a), the exact EPR and long-time estimated EPR with and without bounding the (pseudo) variance are shown in dependence of $\epsilon$. Specifically, the choice $F(\tau)=\langle V_{x}^{\mathrm{A}}+V_{y}^{\mathrm{A}}\rangle_{\mathrm{ps}}/2$ is used. The inset shows the resulting quality factors. The l.h.s. of the stationary CB, $G(t)-\mathcal{D}_{t}$, is shown in (b) evaluated analytically and with simulations. Additionally, the long-time limit is included. The deviation from the long-time limit $\delta(G(t)-\mathcal{D}_{t})$ is presented in (c) correctly displaying the $\mathcal{O}(t^{-1})$ decay. The simulations are performed with the Gillespie algorithm using $N=50000$ trajectories.
V Response to General Perturbation
Numerous works addressed the response of Markovian dynamics to perturbations by means of various methods in the setting of equilibrium systems [154, 155, 156], nonequilibrium steady-states [157, 158, 18, 159, 160, 69], as well as general perturbations of single-time observables [161] which was recently generalized to path observables of diffusions [136]. Here we develop the counterpart of [136] for MJP using the developed stochastic calculus.
V.1 Derivation of Response Formula for MJP
We are interested in the response of a (generally path-wise) observable $O(t)=O[(x_{\tau})_{0\leq\tau\leq t}]$ to a perturbation of a control parameter $\chi\to\chi+\delta$ (e.g., a temperature change $T\to T+\delta$). We seek for an exact result for
$\partial_{\delta}O_{t}=\lim_{\delta\to 0}\frac{1}{\delta}\left(\langle O(t)\rangle_{\delta}-\langle O(t)\rangle\right),$ (132)
where $\langle O(t)\rangle$ is the expectation w.r.t. the unperturbed process in Eq. (39), while $\langle O(t)\rangle_{\delta}$ is the average w.r.t. the perturbed process evolving according to
$\mathrm{d}\bm{n}^{\delta}(\tau)=\bm{R}^{\delta}(x^{\delta}_{\tau})\mathrm{d}\tau+\mathrm{d}\bm{\varepsilon}^{\delta}(\tau)\,.$ (133)
Depending on the observable, the averages $\langle O(t)\rangle$ and $\langle O(t)\rangle_{\delta}$ can either be understood as single-time or path-wise averages. For example, if $O(t)$ measures some instantaneous quantity at time $t$, e.g., the instantaneous internal energy of the system, the averages are to be understood as $\langle O(t)\rangle_{(\delta)}=\sum_{x}O_{x}(t)p^{(\delta)}_{x}(t)$, where $p^{(\delta)}_{x}(t)$ is the single-time probability evolving according to the master equation of the (perturbed) system. Conversely, if $O(t)=O[(x_{\tau})_{0\leq\tau\leq t}]$ is a path-wise observable depending on the entire trajectory, e.g., a time-integrated current $J_{t}$, the average $\langle\cdot\rangle$ is over all paths evolving according to Eq. (39) [and $\langle\cdot\rangle_{(\delta)}$ over all paths evolving according Eq. (133)]
$\langle O(t)\rangle_{(\delta)}=\sum_{(x_{\tau})_{0\leq\tau\leq t}}O[(x_{\tau})_{0\leq\tau\leq t}]\mathds{P}^{(\delta)}[(x_{\tau})_{0\leq\tau\leq t}].$ (134)
The path measures Eq. (33) of processes in Eqs. (39) and (133) can be conveniently written in terms of the stochastic differentials as
$\mathds{P}^{\delta}[(x_{\tau})_{0\leq\tau\leq t}]=p_{x_{0}}(0)\exp\left[\int_{s=0}^{s=t}\sum_{x}r^{\delta}_{xx}\mathrm{d}\tau_{x}(s)+\int_{s=0}^{s=t}\sum_{x,y\neq x}\log r_{xy}^{\delta}\mathrm{d}n_{xy}(s)\right],$ (135)
where $p_{x_{0}}(0)$ is the distribution of the initial state $x_{\tau=0}=x_{0}$. Note that a given path realization defines a unique sequence of transitions and hence $\mathrm{d}\tau_{x}(s)$ and $\mathrm{d}n_{xy}(s)$, which in turn enter equally $\mathds{P}$ in Eq. (33) and $\mathds{P}^{\delta}$ in Eq. (135). However, because of the different rates ($r^{\delta}_{xy}$ versus $r_{xy}$), the measure of the same path is generally different under $\mathds{P}$ and $\mathds{P}^{\delta}$. Moreover, writing the path measure in terms of stochastic differentials as in Eq. (135) further has the advantage of allowing for time-dependent rates.
Consider the perturbed rates
$r_{xy}^{\delta}=r_{xy}(\chi+\delta)=r_{xy}(\chi)+\delta\tilde{r}_{xy}(\chi)+\mathcal{O}(\delta^{2}).$ (136)
Note that we assume that the rates $r_{xy}=r_{xy}(\chi)$ and $r_{xy}^{\delta}$ are functions of $\chi$ and $\chi+\delta$, respectively, ($\chi$ is a placeholder for the perturbed quantity, e.g., temperature $T$) and that the function is analytic at $\delta=0$, i.e., that it has a well-defined Taylor expansion at that point. Therefore, $\mathds{P}^{\delta}$ is absolutely continuous w.r.t. $\mathds{P}$ for sufficiently small $\delta$  [^3] For MJP, the perturbed measure $\mathds{P}^{\delta}\ll\mathds{P}$ is absolutely continuous w.r.t. the unperturbed measure $\mathds{P}$ as long as for all $x\neq y$ $r_{xy}>0\implies r_{xy}^{\delta}>0$, i.e., as long as the underlying topology of the MJP remains unchanged. and the perturbed average $\langle O(t)\rangle_{\delta}$ can be written in terms of the unperturbed process in Eq. (39) using the Cameron-Martin-Girsanov theorem with the Radon-Nikodym derivative [163]
$\langle O(t)\rangle_{\delta}=\left\langle O(t)\frac{\mathrm{d}\mathds{P}^{\delta}}{\mathrm{d}\mathds{P}}\right\rangle$ (137)
where we assumed that the processes start from the same initial distribution. The Radon-Nikodym derivative reads
$\frac{\mathrm{d}\mathds{P}^{\delta}}{\mathrm{d}\mathds{P}}\left[(x_{\tau})_{0\leq\tau\leq t}\right]$ $=\exp\left[\int_{s=0}^{s=t}\sum_{x}(r_{xx}^{\delta}-r_{xx})\mathrm{d}\tau_{x}(s)+\int_{s=0}^{s=t}\sum_{x,y\neq x}\log\frac{r_{xy}^{\delta}}{r_{xy}}\mathrm{d}n_{xy}(s)\right]$
$=\exp\left[-\delta\int_{s=0}^{s=t}\sum_{x,y\neq x}\tilde{r}_{xy}\mathrm{d}\tau_{x}(s)+\delta\int_{s=0}^{s=t}\sum_{x,y\neq x}\frac{\tilde{r}_{xy}}{r_{xy}}\mathrm{d}n_{xy}(s)+\mathcal{O}(\delta^{2})\right]$
$=1-\delta\int_{s=0}^{s=t}\sum_{x,y\neq x}\frac{\tilde{r}_{xy}}{r_{xy}}\left[r_{xy}\mathrm{d}\tau_{x}(s)-\mathrm{d}n_{xy}(s)\right]+\mathcal{O}(\delta^{2})$
$=1+\delta\int_{s=0}^{s=t}\sum_{x,y\neq x}\frac{\tilde{r}_{xy}}{r_{xy}}\mathrm{d}\varepsilon_{xy}(s)+\mathcal{O}(\delta^{2}).$ (138)
The response to a perturbation in $\chi$ (e.g., the temperature) is therefore evaluated to be
$\partial_{\delta}O_{t}=\left\langle O(t)\int_{s=0}^{s=t}\sum_{x,y\neq x}\frac{\tilde{r}_{xy}}{r_{xy}}\mathrm{d}\varepsilon_{xy}(s)\right\rangle.$ (139)
Note that we did not use any particular property of $O(t)$ that would distinguish between single-time and path-wise observables. Nevertheless, for the sake of completeness we provide in Appendix A.8 an explicit derivation of path-wise observables.
While elegant, Eq. (139) (similarly to the continuous-space result [136]) cannot generally be applied to experimental data. This is because of the integration over the noise (which is never known in experiments) and the effective drift, including what defines the rates, which renders Eq. (139) operationally inaccessible. However, Eq. (139) may be useful in simulations of high-dimensional systems, where the rates and noise are known but the generator $\bm{L}$ cannot be easily diagonalized. Note that $\mathrm{d}\varepsilon_{xy}(s)$ depends on $p_{x}(s)$ through Eq. (40).
The stochastic-analysis route to Eq. (139) can be seen as an alternative way of evaluating responses to perturbations of observables complementing well established existing methods [161]. The limit in Eq. (132) can in fact be evaluated directly using a type of Dyson identity [161] (see App. A.8 for details). However, Eq. (139) allows for the same result to be obtained using correlations of stochastic differentials without invoking perturbative calculations.
V.2 Equilibrium Response to Temperature Perturbations
For MJP in equilibrium we have $p_{i}^{\mathrm{eq}}\propto\mathrm{e}^{-E_{i}/T}$ in terms of free energies $E_{i}$. In addition, we have detailed balance, which, however, does not uniquely specify the rates. We choose the rates to be
$r_{xy}=D\mathrm{e}^{\left(\lambda E_{x}-(1-\lambda)E_{y}\right)/T}$ (140)
for any $\lambda\in[0,1]$ (commonly used are $\lambda=1$ and $\lambda=1/2$) and $D$ may scale with $T$ ($D=\tilde{D}T$ with $\tilde{D}=\mathrm{const.}$). The rates (140) satisfy detailed balance.
The aim here is to relate the response in Eq. (139) of a single-time observable $A(t)$ to the equilibrium correlation function,
$C_{OE}(t)=\left(\langle OE\rangle_{\mathrm{eq}}-\langle O(t)E(0)\rangle_{\mathrm{eq}}\right).$ (141)
We start from Eq. (139) and use the noise-time lemma in Eq. (42) to get
$\left\langle O(t)\int_{s=0}^{s=t}\sum_{x,y\neq x}\frac{\tilde{r}_{xy}}{r_{xy}}\mathrm{d}\varepsilon_{xy}(s)\right\rangle_{\mathrm{eq}}=$
$\frac{1}{T^{2}}\!\!\int_{0}^{t}\!\!\mathrm{d}s\!\sum_{i}\!\!\!\sum_{x,y\neq x}\!\!\!O_{i}\left[P(i,t|y,s)-P(i,t|x,s)\right]\times$
$r_{xy}p_{x}^{\mathrm{eq}}\left(-\lambda E_{x}+(1-\lambda)E_{y}+{T}\right),$ (142)
since
$\frac{\tilde{r}_{xy}}{r_{xy}}$ $=\frac{\partial_{\delta}\tilde{D}(T+\delta)\mathrm{e}^{\left(\lambda E_{x}-(1-\lambda)E_{y}\right)/(T+\delta)}|_{\delta=0}}{\tilde{D}T\mathrm{e}^{\left(\lambda E_{x}-(1-\lambda)E_{y}\right)/T}}$
$=\frac{\tilde{D}(1-\frac{\lambda E_{x}-(1-\lambda)E_{y}}{T})\mathrm{e}^{\left(\lambda E_{x}-(1-\lambda)E_{y}\right)/T}}{\tilde{D}T\mathrm{e}^{\left(\lambda E_{x}-(1-\lambda)E_{y}\right)/T}}$
$=\frac{T-\lambda E_{x}+(1-\lambda)E_{y}}{T^{2}}.$ (143)
The last term in Eq. (142) including $T$ vanishes because
$\sum_{x,y\neq x}\left[P(i,t|y,s)-P(i,t|x,s)\right]r_{xy}p_{x}^{\mathrm{eq}}$
$=$ $\sum_{x,y\neq x}P(i,t|x,s)r_{yx}p_{y}^{\mathrm{eq}}-\sum_{x,y\neq x}P(i,t|x,s)r_{xy}p_{x}^{\mathrm{eq}}$
$=$ $\sum_{x,y\neq x}P(i,t|x,s)r_{xy}p_{x}^{\mathrm{eq}}-\sum_{x,y\neq x}P(i,t|x,s)r_{xy}p_{x}^{\mathrm{eq}}$
$=$ $0,$ (144)
where we use detailed balance in the third line. Similarly, the first terms in Eq. (142) simplifies to
$\!\sum_{x,y\neq x}\!\!\left[P(i,t|y,s)-P(i,t|x,s)\right]r_{xy}p_{x}^{\mathrm{eq}}\left(-\lambda E_{x}+(1-\lambda)E_{y}\right)$
$=$ $\sum_{x,y\neq x}P(i,t|x,s)r_{xy}p_{x}^{\mathrm{eq}}\times$
$\underbrace{\left(-\lambda E_{y}+(1-\lambda)E_{x}+\lambda E_{x}-(1-\lambda)E_{y}\right)}_{=E_{x}-E_{y}}$
$=$ $-\sum_{x}p_{x}^{\mathrm{eq}}E_{x}\underbrace{\sum_{y\neq x}\left[P(i,t|y,s)-P(i,t|x,s)\right]r_{xy}}_{=\sum_{y\neq x}P(i,t|y,s)r_{xy}+P(i,t|x,s)r_{xx}}$
$=-\sum_{x}p_{x}^{\mathrm{eq}}E_{x}\underbrace{\sum_{y}P(i,t|y,s)L_{yx}}_{=-\partial_{s}P(i,t|x,s)}.$ (145)
Plugging Eqs. (144) and (145) into Eq. (142), we finally get
$\left\langle O(t)\int_{s=0}^{s=t}\sum_{x,y\neq x}\frac{\tilde{r}_{xy}}{r_{xy}}\mathrm{d}\varepsilon_{xy}(s)\right\rangle_{\mathrm{eq}}=$ $\frac{1}{T^{2}}\int_{0}^{t}\mathrm{d}s\sum_{i}\sum_{x}O_{i}E_{x}p_{x}^{\mathrm{eq}}\partial_{s}P(i,t|x,s)$
$=$ $\frac{1}{T^{2}}\left(\sum_{i}\sum_{x}O_{i}E_{x}p_{x}^{\mathrm{eq}}\underbrace{P(i,t|x,t)}_{\delta_{ix}}-\sum_{i}\sum_{x}O_{i}E_{x}p_{x}^{\mathrm{eq}}P(i,t|x,0)\right)$
$=$ $\frac{1}{T^{2}}\left(\langle OE\rangle_{\mathrm{eq}}-\langle O(t)E(0)\rangle\right).$ (146)
Note that Eq. (146) does not explicitly depend on $\lambda$, only implicitly through the propagator, i.e., it holds for every parametrization of the rates in Eq. (63). Equation (146) is the linear response of an equilibrium MJP to a temperature perturbation.
V.3 Examples
To visualize the results in Eqs. (139) and (146), we consider the response of a single time observable $O(t)=(1,1,0,0)$ of a four-state ring in an equilibrium system (see Fig. 11a) and in a transient system with constant driving (see Fig. 11b), as well as of a current $J_{t}$ with weight $\kappa_{ij}=\delta_{i4}\delta_{j1}-\delta_{i1}\delta_{j4}+\delta_{i3}\delta_{j2}-\delta_{i2}\delta_{j3}$ in transient systems with equilibrium (see Fig. 11c) rates and constant driving (see Fig. 11d).
For simplicity we set $D=T$ in Eq. (140). In Fig. 11a and Fig. 11c we use Eq. (63) with the free energies listed in Tab. 4. In Fig. 11a, we compare $\lambda=0.3$ and $\lambda=0.9$, while in Fig. 11c we use $\lambda=0.5$. The rates used in Fig. 11b and Fig. 11d are $r_{12}=r_{23}=r_{34}=r_{41}=T\mathrm{e}^{-\log\epsilon/T}$ in one direction and $r_{21}=r_{32}=r_{43}=r_{14}=T\mathrm{e}^{-\log\alpha/T}$ in the opposite direction with $\epsilon=3$ and $\alpha=2$, respectively. The temperature is set to $T=1$ and equilibrium initial conditions $p^{\mathrm{init}}_{i}=p^{\mathrm{eq}}_{i}\propto e^{-E_{i}/T}$ are assumed in Fig. 11a, while $T=2$ and initial condition $p^{\mathrm{init}}=(0,0,1,0)$ is used in Fig. 11b-Fig. 11d. The diamonds on the x-axis denote the largest relaxation time of the respective generators.
As a visual aid, the value of the equilibrium covariance between $O$ and $E$, $\mathrm{cov}_{\mathrm{eq}}(O,E)=\langle OE\rangle_{\mathrm{eq}}-\langle O\rangle_{\mathrm{eq}}\langle E\rangle_{\mathrm{eq}}$, is included in Fig. 11a to show how, independently of $\lambda$, the response $\partial_{\delta}O_{t}\to\mathrm{cov}_{\mathrm{eq}}(O,E)$ for $t\to\infty$.
Figure 11: Perturbations of the equilibrium and driven rings. Analytical and numerical response curves of $O=(1,1,0,0)$ for equilibrium (a) and transient (b) dynamics on a four state ring. In (a), we compare the numerical simulations (solid lines) with theoretical prediction Eq. (146) (dashed lines) for $\lambda=0.3$ (orange) and $\lambda=0.9$ (green). The remaining parameters are $T=1$, $\delta=0.1$, $N=5\cdot 10^{6}$, and $\tilde{D}_{xy}=1$. In (b), the initial distribution is $p^{\mathrm{init}}=(0,0,1,0)$ and the rates are $r_{12}=r_{23}=r_{34}=r_{41}=T\mathrm{e}^{-\log\epsilon/T}$ and $r_{21}=r_{32}=r_{43}=r_{14}=T\mathrm{e}^{-\log\alpha/T}$ with $\epsilon=3$ and $\alpha=2$, respectively. The theoretical result Eq. (203) (dashed purple line) is compared with numerical simulations using $N=5\cdot 10^{5}$ (solid yellow line) and $N=5\cdot 10^{6}$ (solid green line) trajectories, respectively. The remaining parameters are $T=2$, $\delta=0.1$, and $\tilde{D}_{xy}=1$. The markers in (a) and (b) denote respective the largest relaxation time-scales in the systems. The dashed black lines are the expected $t\to\infty$ values $\mathrm{cov}_{\mathrm{eq}}(A,E)$ in (a) and 0 in (b). Figure (c) shows the response to a current with weight $\kappa_{ij}=\delta_{i4}\delta_{j1}-\delta_{i1}\delta_{j4}+\delta_{i3}\delta_{j2}-\delta_{i2}\delta_{j3}$ for the aforementioned equilibrium rates with $T=2$, $\delta=0.05$, $N=2\cdot 10^{6}$, $\lambda=0.5$, and initial condition as in (b). In (d), the same parameters as in (c) are used, with the same rates as in (b). The dashed orange lines in (c) and (d) are the theoretically predicted values, while the blue markers are values extracted from numerical simulations. In (a-d) the uncertainty bands are extrapolated from bootstrapping the trajectories into 2020 samples with $30\%$ of the respective NNs. The free energies, $E_{i}$, used in (a) and (c) are listed in Tab. 4. The simulations are performed with the celebrated Gillespie algorithm.
VI Continuum limit
While we treated continuous-space (Langevin) and discrete-state MJP dynamics separately so far, it is well known that the two descriptions should agree in a particular limit. Namely, a $d$-dimensional continuous description can be approximated by a jump process on a $d$-dimensional grid (here, we assume a hypercubic lattice) with grid spacing $\Delta x$}, where the MJP description reproduces the dynamics as ${\Delta x}\to 0$ or for observations on scales much larger than $\Delta x$} (see, e.g., Ref. [164]). For thermodynamic consistency, we here show that both, the entropy Eq. (66) and pseudo entropy Eq. (69) approach the continuous-space expression Eq. (64) as ${\Delta x}\to 0$ for any jump dynamics on a grid approaching the continuous-space dynamics in this limit. In particular, this implies that the pseudo entropy Eq. (69) and entropy Eq. (66) agree in this limit (as they do for continuous dynamics). This in turn implies that the bounds for jump processes can also become saturated in this continuum limit, exactly as they can be saturated for continuous-space dynamics.
We now introduce jump rates for a given $d$-dimensional continuous dynamics $\mathrm{d}{\bm{x}}_{\tau}=\bm{F}({\bm{x}}_{\tau})\mathrm{d}\tau+\sqrt{2D}\mathrm{d}\bm{W}_{\tau}$ and given $\Delta x$}. For simplicity, we first consider dynamics settling into an equilibrium for long times, i.e., $\bm{F}({\bm{x}})=-D\nabla U({\bm{x}})$ for some sufficiently confining potential $U({\bm{x}})$ (in principle the approach equally works for irreversible (i.e., $D^{-1}\bm{F}$ not a potential field) and even non-ergodic dynamics). Following Ref. [164], the rate from $\bm{x}$} to a neighboring state ${\bm{x}}^{\prime}={\bm{x}}+{\Delta{\bm{x}}}$ can be written for small $\Delta x$} as
$r_{{\bm{x}}{\bm{x}}^{\prime}}$ $=\frac{D}{{\Delta x}^{2}}\exp\left[-\frac{U({\bm{x}}^{\prime})-U({\bm{x}})}{2}\right]$
$\approx\frac{D}{{\Delta x}^{2}}\exp\left[-\frac{\nabla U({\bm{x}})\cdot{\Delta{\bm{x}}}}{2}\right]\,.$ (147)
[…]
$\!\!\!\!\!\!\ln\left(\frac{p_{\bm{x}}(\tau)}{p_{{\bm{x}}^{\prime}}(\tau)}\right)$ $\approx\ln[p({\bm{x}},\tau)]-\ln[p({\bm{x}},\tau)+\nabla p({\bm{x}},\tau)\cdot{\Delta{\bm{x}}}]$
$\approx-\nabla\ln[p({\bm{x}},\tau)]\cdot{\Delta{\bm{x}}}\,$,
[…]
$=-\nabla U({\bm{x}})\cdot{\Delta{\bm{x}}}\,$,
$\ln\left(\frac{p_{\bm{x}}(\tau)r_{{\bm{x}}{\bm{x}}^{\prime}}}{p_{{\bm{x}}^{\prime}}(\tau)r_{{\bm{x}}^{\prime}{\bm{x}}}}\right)$ $=[-\nabla U({\bm{x}})-\nabla\ln[p({\bm{x}},\tau)]\cdot{\Delta{\bm{x}}}$
$=\frac{\bm{j}({\bm{x}},\tau)}{Dp({\bm{x}},\tau)}\cdot{\Delta{\bm{x}}}\,.$ (148)
Similarly, we obtain
$p_{\bm{x}}(\tau)r_{{\bm{x}}{\bm{x}}^{\prime}}-p_{{\bm{x}}^{\prime}}(\tau)r_{{\bm{x}}^{\prime}{\bm{x}}}$
$\approx{\Delta x}^{d}p({\bm{x}},\tau)r_{{\bm{x}}{\bm{x}}^{\prime}}-{\Delta x}^{d}[p({\bm{x}},\tau)+\nabla p({\bm{x}},\tau)\cdot{\Delta{\bm{x}}}]r_{{\bm{x}}^{\prime}{\bm{x}}}$
$\approx{\Delta x}^{d}p({\bm{x}},\tau)\left(r_{{\bm{x}}{\bm{x}}^{\prime}}-r_{{\bm{x}}^{\prime}{\bm{x}}}-\nabla\ln[p({\bm{x}},\tau)]\cdot{\Delta{\bm{x}}}\right)$
$\approx{\Delta x}^{d-2}Dp(x,\tau)\left(-\nabla U({\bm{x}})-\nabla\ln[p({\bm{x}},\tau)]\right)\cdot{\Delta{\bm{x}}}$
$={\Delta x}^{d-2}\bm{j}({\bm{x}},\tau)\cdot{\Delta{\bm{x}}}\,$, (149)
and
$p_{\bm{x}}(\tau)r_{{\bm{x}}{\bm{x}}^{\prime}}+p_{{\bm{x}}^{\prime}}(\tau)r_{{\bm{x}}^{\prime}{\bm{x}}}$
$\approx{\Delta x}^{d}p({\bm{x}},\tau)r_{{\bm{x}}{\bm{x}}^{\prime}}+[p({\bm{x}},\tau)+\nabla p({\bm{x}},\tau)\cdot{\Delta{\bm{x}}}]r_{{\bm{x}}^{\prime}{\bm{x}}}$
$\approx{\Delta x}^{d-2}2Dp(x,\tau)\,.$ (150)
For the simple hypercubic lattice, the increment vectors $\Delta{\bm{x}}$} to neighboring states are ${\Delta{\bm{x}}}=\pm{\Delta x}\bm{e}_{i}$. Here, $\bm{e}_{i}$ with $i\in\{1,2,\dots,d\}$ is a unit vector in the $i$-th coordinate, projecting onto individual components as $\bm{j}\cdot{\Delta{\bm{x}}}=\pm j_{i}{\Delta x}$. The relevant terms from the definitions in Eqs. (69) and (66) for the pseudo-entropy and entropy turn out to be
$\frac{1}{2}[p_{\bm{x}}(\tau)r_{{\bm{x}}{\bm{x}}^{\prime}}-p_{{\bm{x}}^{\prime}}(\tau)r_{{\bm{x}}^{\prime}{\bm{x}}}]\ln\left[\frac{p_{\bm{x}}(\tau)r_{{\bm{x}}{\bm{x}}^{\prime}}}{p_{{\bm{x}}^{\prime}}(\tau)r_{{\bm{x}}^{\prime}{\bm{x}}}}\right]$ (151)
$\approx{\Delta x}^{d}\frac{j_{i}({\bm{x}},\tau)^{2}}{2Dp({\bm{x}},\tau)}\,$,
$\frac{[p_{\bm{x}}(\tau)r_{{\bm{x}}{\bm{x}}^{\prime}}-p_{{\bm{x}}^{\prime}}(\tau)r_{{\bm{x}}^{\prime}{\bm{x}}}]^{2}}{p_{\bm{x}}(\tau)r_{{\bm{x}}{\bm{x}}^{\prime}}+p_{{\bm{x}}^{\prime}}(\tau)r_{{\bm{x}}^{\prime}{\bm{x}}}}\approx{\Delta x}^{d}\frac{j_{i}({\bm{x}},\tau)^{2}}{2Dp({\bm{x}},\tau)}\,.$ (152)
A summation over $i$ yields $\sum_{i=1}^{d}j_{i}({\bm{x}},\tau)^{2}=\bm{j}({\bm{x}},\tau)^{2}$. Upon summation over ${\bm{x}},{\bm{x}}^{\prime}$ as done in Eqs. (69) and (66) for the pseudo-entropy and entropy production, respectively, we find that both quantities approach the continuous-space formula for the entropy production in Eq. (64). Following Ref. [164] one can directly extend the argument to non-isotropic diffusion and non-conservative dynamics by using different “pseudopotentials” for each entry of the vector $\bm{F}({\bm{x}})$ and thus the above statements about the limit generalize.
[…]
VII Time-Dependent Driving
We now show how most of the results shown so far generalize to systems with time-dependent driving (i.e., time-inhomogeneous systems), i.e., where the generator is explicitly time-dependent $\bm{L}(v\tau)$ with some (constant) protocol velocity $v$. Notably, this means that the propagator is no longer time-translation invariant, $P(i,\tau^{\prime}|x,\tau;v)\neq P(i,\tau^{\prime}-\tau|x,0;v)$, as it reads [4, 60]
$P(i,\tau^{\prime}|x,\tau;v)=\left[\mathcal{T}\exp\left\{\int_{\tau}^{\tau^{\prime}}\mathrm{d}s\bm{L}(vs)\right\}\right]_{ix},$ (153)
where $\mathcal{T}$ is the time ordering operator. A straightforward calculation reveals that the expressions for increment correlations (52), (54), and (59), observable covariances (62), and transport bound (93) remain unchanged apart form additional arguments for $v$. For instance, the observable covariances become
$\mathrm{cov}({\rho}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P(x,\tau;i,\tau^{\prime};v)-p_{x}(\tau;v)p_{i}(\tau^{\prime};v)\right],$
$\mathrm{cov}({J}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\mathds{1}_{\tau>\tau^{\prime}}r_{xy}P(x,\tau|i,\tau^{\prime};v)p_{i}(\tau^{\prime};v)+\mathds{1}_{\tau<\tau^{\prime}}r_{yx}(v\tau)P(i,\tau^{\prime}|y,\tau;v)p_{y}(\tau;v)-r_{xy}(v\tau)p_{x}(\tau;v)p_{i}(\tau^{\prime};v)\right],$
$\mathrm{cov}(J_{t}^{k},{J}_{t}^{l})=$ $\int_{0}^{t}\mathrm{d}\tau\sum_{x,y}\kappa_{xy}^{k}(v\tau)\kappa_{xy}^{l}(v\tau)r_{xy}(v\tau)p_{x}(\tau;v)$
$+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\mathds{1}_{\tau>\tau^{\prime}}r_{xy}(v\tau)r_{ji}(v\tau^{\prime})P(x,\tau|j,\tau^{\prime};v)p_{i}(\tau^{\prime};v)+\mathds{1}_{\tau<\tau^{\prime}}r_{ij}(v\tau^{\prime})r_{yx}(v\tau)P(i,\tau^{\prime}|y,\tau;v)p_{y}(\tau;v)\right$.
$-\left.r_{xy}(v\tau)p_{x}(\tau;v)r_{ij}(v\tau^{\prime})p_{i}(\tau^{\prime};v)\right],$ (154)
where the functions $V^{l/k}(v\tau)\delta$ and $\kappa^{l/k}(v\tau)$ inside the integration operator $\hat{{I}}^{t}_{\cdot,\cdot}$ are time dependent as well.
However, this does not hold for the CTUR as shown in Refs. [60, 59], since the TUR with time-dependent driving reads [60, 59]
$\frac{\left[\left(t\partial_{t}-v\partial_{v}\right)\langle J_{t}\rangle\right]^{2}}{\mathrm{var}(J_{t})}\leq\frac{\Delta S_{\mathrm{tot}}(t)}{2},$ (155)
and thus requires some special care. The observable current and entropy production entering Eq. (155) read
$\Delta S_{\mathrm{tot}}(t)$ $=\int_{0}^{t}\mathrm{d}\tau\sum_{x,y\neq x}r_{xy}(v\tau)p_{x}(\tau;v)\log\frac{r_{xy}(v\tau)p_{x}(\tau;v)}{r_{yx}(v\tau)p_{y}(\tau;v)},$
$J_{t}$ $=\int_{\tau=0}^{\tau=t}\mathrm{Tr}\left[\bm{\kappa}(v\tau)^{T}\mathrm{d}\bm{n}(\tau)\right].$ (156)
Note that although we write $\mathrm{d}\bm{n}(\tau)$, the differential is not independent of $v$. This can be seen, e.g., in the statistical properties, such as the mean $\langle\mathrm{d}n_{xy}(\tau)\rangle_{x_{\tau}=i}=\delta_{ix}r_{xy}(v\tau)\mathrm{d}\tau$. However, as the differential $\mathrm{d}\bm{n}(\tau)$ is “observed” at time $\tau$, regardless of $v$, we choose to suppress the $v$ argument. Central to the proof of the TUR for driven systems is the identity [60]
$\int_{0}^{\tau^{\prime}}\mathrm{d}\tau\sum_{x}\left[\partial_{\tau}P(i,\tau^{\prime}|x,\tau;v)\right]p_{x}(\tau;v)$
$=-(\tau^{\prime}\partial_{\tau^{\prime}}-v\partial_{v})p_{i}(\tau^{\prime};v)\,.$ (157)
With this identity, we can extend the time-dependent TUR in Eq. (155) to also include densities,
$\rho_{t}$ $=\int_{\tau=0}^{\tau=t}\bm{V}(v\tau)^{T}\mathrm{d}\bm{\tau}(\tau),$ (158)
as well. To do so, we require $\langle A_{t}\rho_{t}\rangle$, which after some simplification using Eq. (42) reads
$\langle A_{t}\rho_{t}\rangle=$ $\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i}V_{i}(v\tau^{\prime})\mathds{1}_{\tau<\tau^{\prime}}\sum_{x,y}\left[P(i,\tau^{\prime}|y,\tau;v)-P(i,\tau^{\prime}|x,\tau;v)\right]p_{x}(\tau;v)r_{xy}(v\tau)Z_{xy}(v\tau).$ (159)
Rearranging the inner sum ($x,y$), we find that Eq. (159) simplifies to
$\langle$ $A_{t}\rho_{t}\rangle$ (160)
$=$ $\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{\tau^{\prime}}\mathrm{d}\tau\sum_{i,x}V_{i}(v\tau^{\prime})P(i,\tau^{\prime}|x,\tau;v)\partial_{\tau}p_{x}(\tau;v)$
$=$ $-\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{\tau^{\prime}}\mathrm{d}\tau\sum_{i,x}V_{i}(v\tau^{\prime})\left[\partial_{\tau}P(i,\tau^{\prime}|x,\tau;v)\right]p_{x}(\tau;v),$
where we performed a partial integration over $\tau$ to obtain the third line. The boundary term vanishes because $\sum_{x}P(i,\tau^{\prime}|x,\tau^{\prime};v)p_{x}(\tau^{\prime};v)=\sum_{x}P(i,\tau^{\prime}|x,0;v)p_{x}(0;v)$. Using Eq. (157), the integral over $\tau$ and sum over $x$ vanish, such that
$\langle$ $A_{t}\rho_{t}\rangle=\int_{0}^{t}\mathrm{d}\tau^{\prime}\sum_{i}V_{i}(v\tau^{\prime})\left(\tau^{\prime}\partial_{\tau^{\prime}}-v\partial_{v}\right)p(\tau^{\prime};v).$ (161)
Another integration by parts and recognizing $\tau^{\prime}\partial_{\tau^{\prime}}V_{i}(v\tau^{\prime})=v\partial_{v}V_{i}(v\tau^{\prime})$ finally yields
$\langle$ $A_{t}\rho_{t}\rangle=\left(t\partial_{t}-v\partial_{v}-1\right)\langle\rho_{t}\rangle.$ (162)
Hence, the Cauchy-Schwarz inequality delivers (for the first time) the CTUR for time-dependent driving
$\frac{\left(\hat{\Lambda}\langle J_{t}\rangle-c(t)\left\{\left[\hat{\Lambda}-1\right]\langle\rho_{t}\rangle\right\}\right)^{2}}{\mathrm{var}(J_{t}-c(t)\rho_{t})}\leq\frac{\Delta S_{\mathrm{tot}}}{2},$ (163)
with differential operator $\hat{\Lambda}\equiv t\partial_{t}-v\partial_{v}$. Notably, we are still able to saturate the Cauchy-Schwarz inequality, although the practical inaccessibility of the required transition weights and state function remains. However, one may again optimize the CTUR Eq. (163) w.r.t. $c(t)$. The results are again given by Eq. (83) albeit with $a(t)=\hat{\Lambda}\langle J_{t}\rangle$ and $b(t)=(\hat{\Lambda}-1)\langle\rho_{t}\rangle$.
Conversely, the correlation bound is not easily generalizable to systems with time-dependent driving. Specifically, since $z_{x}^{F(\tau)}$ in general explicitly depends on time $\tau$ through $F(\tau)$, Eq. (157) cannot be applied to shift the time derivative of the propagator. Hence, the expectation $\langle B_{t}C_{t}\rangle$ [see Eq. (113)] does not reduce to an accessible quantity.
VIII Outlook
We developed a stochastic-calculus for path-wise observables (i.e., functionals) of Markov-jump processes that unifies the different approaches to Markov-jump dynamics, as well as the descriptions of diffusion and jump dynamics. We presented the approach as an exact parallelism with the continuous-space diffusion counterpart, starting from a “Langevin equation for Markov-jump processes” together with a central noise-time-correlation Lemma, defined general path-wise observables and determined their complete (co)variation structure, and identified (generalized) Green-Kubo relations. The approach includes general transients and time-inhomogeneous dynamics (e.g., time-dependent driving). We used the stochastic calculus to prove directly the known thermodynamic inequalities (TUR, correlation TUR, transport bounds, bounds on correlations etc.) in their most general form and, afforded by the directness of the approach, discussed their saturation conditions. We showed that these inequalities follow directly from the equations of motion, which establishes them as an inherent property of stochastic (Markov-jump) equations of motion.
We further derived the response of any (generally path-wise) observable to a general (incl. thermal) perturbation. Finally, we established the continuum limit to achieve the complete unification of diffusion and jump dynamics. Our results ultimately place functionals of diffusion and jump processes on an equal footing on the level of individual stochastic realizations, and hence achieve a “contraction” of two until now disjoint frameworks in time-average statistical mechanics.
The results enable, and are expected to inspire, new directions of research. In particular, one may extend the jump-dynamics results to account for dynamics in phase space (i.e., momentum exchange) in the spirit of [165] (see Chapter 11 therein) on the discrete-state side. Conversely, one may also deepen the results on underdamped dynamics (i.e., unify the transport bound and the TUR derived in [126]). Moreover, we expect our results to inspire research in the direction of discrete-state analogs of generative diffusion models as in Ref. [71, 166] and the learning of stochastic thermodynamics from fluctuating discrete-state trajectories as recently carried out for diffusions in Ref. [167]. There the power of stochastic calculus unfolds in a very prominent manner.
Uffbassa! Mir hen da wos g’scheid’s gmochdd!
IX Data availability
The data that support the findings of this article are openly available at [168].
Acknowledgments
The financial support from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program (Grant Agreement No. 101086182 to AG) and the German Research Foundation (DFG) through the Heisenberg Program (grant GO 2762/4-1 to AG) is gratefully acknowledged.
Appendix A Proofs and Derivations
In the Appendix we provide some technical results that are required to understand the detailed derivations but are not critical for a conceptual understanding of the results.
A.1 Proof of Noise-Time Correlation Lemma
To prove how the noise Eq. (37) and time spent in a state in Eq. (36) correlate, i.e., $\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$, we consider three cases: the transition from $x$ to $y$ happens (i) before $\tau^{\prime}$, (ii) at $\tau^{\prime}$, or (iii) after $\tau^{\prime}$, corresponding to $\tau<\tau^{\prime}$, $\tau=\tau^{\prime}$, and $\tau>\tau^{\prime}$, respectively. Specifically, we can write
$\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle=$ $\mathds{1}_{\tau\geq\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$
$+\mathds{1}_{\tau<\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle,$ (164)
where the first term can be written as
$\mathds{1}_{\tau\geq\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$
$=\underbrace{\sum_{\mathrm{d}\varepsilon_{xy}}\mathrm{d}\varepsilon_{xy}p(\mathrm{d}\varepsilon_{xy})}_{=0}P(x,\tau|i,\tau^{\prime})p_{i}(\tau^{\prime}).$ (165)
Since the system is Markovian, (ii) and (iii) do not contribute. Hence, without loss of generality we can write
$\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle=\mathds{1}_{\tau<\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle.$ (166)
For small $\mathrm{d}\tau\to 0$, the noise increment can only take two values
$\mathrm{d}\varepsilon_{xy}(\tau)=\mathds{1}_{x}(\tau)\begin{cases}1-r_{xy}\mathrm{d}\tau&\text{transition},\\ -r_{xy}\mathrm{d}\tau&\text{no transition}.\end{cases}$ (167)
These values occur with probability $r_{xy}\mathrm{d}\tau$ and $1-r_{xy}\mathrm{d}\tau$, respectively. Defining the state function
$\gamma(\mathrm{d}\varepsilon_{xy}(\tau))=\begin{cases}y&\text{transition},\\ x&\text{no transition},\end{cases}$ (168)
allows us to explicitly evaluate Eq. (166) by introducing an intermediate point (see Fig. 12), such that to leading order
$\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$ $=\mathds{1}_{\tau<\tau^{\prime}}\mathrm{d}\tau^{\prime}\sum_{\mathrm{d}{\varepsilon}_{xy}(\tau)}p(\mathrm{d}{\varepsilon}_{xy}(\tau))\mathrm{d}{\varepsilon}_{xy}(\tau)P(i,\tau^{\prime}|\gamma(\mathrm{d}{\varepsilon}_{xy}(\tau)),\tau+\mathrm{d}\tau)p_{x}(\tau)$
$=\mathds{1}_{\tau<\tau^{\prime}}\mathrm{d}\tau^{\prime}p_{x}(\tau)r_{xy}\mathrm{d}\tau\left(1-r_{xy}\mathrm{d}\tau\right)\left[P(i,\tau^{\prime}|y,\tau+\mathrm{d}\tau)-P(i,\tau^{\prime}|x,\tau+\mathrm{d}\tau)\right]$
$=\mathds{1}_{\tau<\tau^{\prime}}\mathrm{d}\tau^{\prime}\left[P(i,\tau^{\prime}|y,\tau)-P(i,\tau^{\prime}|x,\tau)\right]p_{x}(\tau)r_{xy}\mathrm{d}\tau\,.$ (169)
Figure 12: Visualization of the intermediate point considered in the proof of Noise-Time correlation lemma. The general idea to evaluate the noise in $x_{\tau}$ at $\tau$ and being in state $x_{\tau^{\prime}}$ at time $\tau^{\prime}>\tau$ is to include the intermediate point $x_{\tau+\mathrm{d}\tau}$.
A.2 Derivation of Increment Correlations
A.2.1 Jump-Time Correlations
We see for an explicit expression for $\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle$. We can decompose this expectation using Eq. (38) as
$\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle=\langle\mathrm{d}\varepsilon_{xy}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle+r_{xy}\langle\mathrm{d}\tau_{x}(\tau)\tau_{i}(\tau^{\prime})\rangle.$ (170)
Using the last two equations in Eqs. (41), we get
$\frac{\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle}{r_{xy}\mathrm{d}\tau\mathrm{d}\tau^{\prime}}=$ $\mathds{1}_{\tau<\tau^{\prime}}P(i,\tau^{\prime}|y,\tau)p_{x}(\tau)$
$+\mathds{1}_{\tau\geq\tau^{\prime}}P(x,\tau|i,\tau^{\prime})p_{i}(\tau^{\prime}).$ (171)
A.2.2 Jump-Jump Correlations
Similar to the jump-time correlation, we again decompose the jump-jump correlation as
$\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}n_{ij}(\tau^{\prime})\rangle=$ $\langle\mathrm{d}\varepsilon_{xy}(\tau)\mathrm{d}\varepsilon_{ij}(\tau^{\prime})\rangle$
$+r_{xy}\langle\mathrm{d}\tau_{x}(\tau)\mathrm{d}\varepsilon_{ij}(\tau^{\prime})\rangle$
$+r_{ij}\langle\mathrm{d}\varepsilon_{xy}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle$
$+r_{xy}r_{ij}\langle\mathrm{d}\tau_{x}(\tau)\tau_{i}(\tau^{\prime})\rangle.$ (172)
Using Eqs. (41) therefore yields
$\frac{\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}n_{ij}(\tau^{\prime})\rangle}{r_{xy}r_{ij}\mathrm{d}\tau\mathrm{d}\tau^{\prime}}=$ $\delta_{ix}\delta_{jy}\delta(\tau-\tau^{\prime})p_{x}(\tau)$
$+\mathds{1}_{\tau<\tau^{\prime}}P(i,\tau^{\prime}|y,\tau)p_{x}(\tau)$
$+\mathds{1}_{\tau\geq\tau^{\prime}}P(x,\tau|j,\tau^{\prime})p_{i}(\tau^{\prime}).$ (173)
A.3 From Equations of Motion to Ensemble Description
A.3.1 Continuous Space: Langevin to Fokker-Planck
To highlight the analogy between continuous- and discrete space dynamics, we first recapitulate how the Fokker-Planck equation follows directly from the Langevin Equation in Eq. (31). The calculation is well ḱnown (see e.g., Ref. [132]). Let $K(x)$ be an arbitrary test function that obeys the boundary conditions. Considering the expectation $\langle K(\bm{x}(t))\rangle=\int\mathrm{d}\bm{x}K(\bm{x})P(\bm{x},t)$ for the dummy function at some time $t$, the temporal derivative can be evaluated using Itô’s lemma [132]
$\frac{\mathrm{d}}{\mathrm{d}t}\langle K(\bm{x}(t))\rangle=\langle\bm{F}(\bm{x})\cdot\nabla K(\bm{x})+\nabla\cdot\bm{D}\nabla K(\bm{x})\rangle$ (174)
$=\int\mathrm{d}\bm{x}\left[\bm{F}(\bm{x})\cdot\nabla K(\bm{x})+\nabla\cdot\bm{D}\nabla K(\bm{x})\right]P(\bm{x},t).$
Equivalently, the derivative can be written as
$\frac{\mathrm{d}}{\mathrm{d}t}\langle K(\bm{x}(t))\rangle=\int\mathrm{d}\bm{x}K(\bm{x})\frac{\partial}{\partial t}P(\bm{x},t).$ (175)
By equating Eqs. (174) and (175) we find
$\int$ $\mathrm{d}\bm{x}K(\bm{x})\frac{\partial}{\partial t}P(\bm{x},t)$
$=$ $\int\mathrm{d}\bm{x}\left[\bm{F}(\bm{x})\cdot\nabla K(\bm{x})+\nabla\cdot\bm{D}\nabla K(\bm{x})\right]P(\bm{x},t)$ (176)
$=$ $\int\mathrm{d}\bm{x}K(\bm{x})\left[-\nabla\left\{\bm{F}(\bm{x})P(\bm{x},t)\right\}\right.+\left.\nabla\cdot\bm{D}\nabla P(\bm{x},t)\right]\,$,
where we performed a partial integration to obtain the last line (the boundary terms vanish by the properties assumed on $K({\bm{x}})$). Since the above equation holds for any test function $K({\bm{x}})$ we derived the Fokker-Planck equation
$\partial_{t}P(\bm{x},t)=[\nabla\cdot\bm{D}\nabla-\nabla\bm{F}(\bm{x})]P(\bm{x},t).$ (177)
A.3.2 Discrete Space: Jumps to Master Equation
To show how the master equation emerges from the stochastic equations of motion in Eq. (38), we first need to consider the indicator function $\mathds{1}_{x}(\tau)$ of being in state $x$ at time $\tau$. Consider $\mathds{1}_{x}(\tau+\mathrm{d}\tau)=\mathds{1}_{x}(\tau)+\frac{\mathrm{d}}{\mathrm{d}\tau}\mathds{1}_{x}(\tau)\mathrm{d}\tau+\mathcal{O}(\mathrm{d}\tau^{2})$, where the derivative should be understood as
$\frac{\mathrm{d}}{\mathrm{d}\tau}\mathds{1}_{x}(\tau)=\lim_{\mathrm{d}\tau\to 0^{+}}\frac{\mathds{1}_{x}(\tau+\mathrm{d}\tau)-\mathds{1}_{x}(\tau)}{\mathrm{d}\tau}.$ (178)
By the non-explosive property of MJP [169] a finite number of jumps occur in any finite time interval with probability 1. Therefore, we do not need to consider any higher order correction, as the probability of a second transition occurring in an infinitesimal increment vanishes. The term $\frac{\mathrm{d}}{\mathrm{d}\tau}\mathds{1}_{x}(\tau)\mathrm{d}\tau$ can therefore take on values
[…]
As a consequence, we can write
$\frac{\mathrm{d}}{\mathrm{d}\tau}\mathds{1}_{x}(\tau)\mathrm{d}\tau=\sum_{y\neq x}\left(\mathrm{d}n_{yx}(\tau)-\mathrm{d}n_{xy}(\tau)\right),$ (180)
because at most one jump differential will be non-zero as $\mathrm{d}\tau\to 0$. The probability of being in a state $x$ at time $\tau$ is $p_{x}(\tau)=\langle\mathds{1}_{x}(\tau)\rangle$. Therefore,
$p_{x}(\tau+\mathrm{d}\tau)-p_{x}(\tau)$ $=\left\langle\mathds{1}_{x}(\tau+\mathrm{d}\tau)-\mathds{1}_{x}(\tau)\right\rangle$
$=\left\langle\frac{\mathrm{d}}{\mathrm{d}\tau}\mathds{1}_{x}(\tau)\mathrm{d}\tau\right\rangle$ (181)
$=\left\langle\sum_{y\neq x}(\mathrm{d}n_{yx}(\tau)-\mathrm{d}n_{xy}(\tau))\right\rangle$
$=\mathrm{d}\tau\sum_{y\neq x}\left(r_{yx}p_{y}(\tau)-r_{xy}p_{x}(\tau)\right)$
Rearranging Eq. (181) and taking the limit $\mathrm{d}\tau\to 0$ yields the master equation (32) in a way equivalent to the Fokker-Planck equation (177).
A.4 Derivation of Eq. (76)
Starting from Eq. (75), we can simplify the sum over $x,y$ by using the symmetry of $Z_{xy}(\tau)[P(i,\tau^{\prime}|y,\tau)-P(i,\tau^{\prime}|x,\tau)]$ to get
$\langle$ $A_{t}J_{t}^{\mathrm{II}}\rangle$
$=$ $\frac{1}{2}\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}\mathds{1}_{\tau<\tau^{\prime}}$
$\times\sum_{x,y}\left[P(i,\tau^{\prime}|y,\tau)-P(i,\tau^{\prime}|x,\tau)\right]\left[p_{x}(\tau)r_{xy}-p_{y}(\tau)r_{yx}\right]$
$=$ $-\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}\mathds{1}_{\tau<\tau^{\prime}}$ (182)
$\times\sum_{x}P(i,\tau^{\prime}|x,\tau)\sum_{y,y\neq x}\left[p_{x}(\tau)r_{xy}-p_{y}(\tau)r_{yx}\right].$
With Eq. (32) and an integration by parts, Eq. (182) reduces to
$\langle$ $A_{t}J_{t}^{\mathrm{II}}\rangle$
$=$ $\int_{0}^{t}\mathrm{d}\tau^{\prime}\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}\left[-\sum_{x}P(i,\tau^{\prime}|x,0)p_{x}(0)\right$.
$-\left.\int_{0}^{t}\mathrm{d}\tau\sum_{x}p_{x}(\tau)\partial_{\tau}\left\{P(i,\tau^{\prime}|x,\tau)\mathds{1}_{\tau<\tau^{\prime}}\right\}\right]$
$=$ $-\langle J_{t}\rangle+\int_{0}^{t}\mathrm{d}\tau^{\prime}\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}\partial_{\tau^{\prime}}\left[\tau^{\prime}p_{i}(\tau^{\prime})\right],$ (183)
where we use that $\partial_{\tau}P(i,\tau^{\prime}|x,\tau)=-\partial_{\tau^{\prime}}P(i,\tau^{\prime}|x,\tau)$, $\partial_{\tau}\mathds{1}_{\tau<\tau^{\prime}}=-\partial_{\tau^{\prime}}\mathds{1}_{\tau<\tau^{\prime}}$, and $\int_{0}^{\tau^{\prime}}\mathrm{d}\tau\mathds{1}_{\tau<\tau^{\prime}}=\tau^{\prime}$. With another integration by parts, the correlator finally becomes
$\langle A_{t}J_{t}^{\mathrm{II}}\rangle=$ $-\langle J_{t}\rangle+\underbrace{t\sum_{i,j}\kappa_{ij}r_{ij}p_{i}(t)}_{=t\partial_{t}\langle J_{t}\rangle}$ (184)
$-\int_{0}^{t}\mathrm{d}\tau^{\prime}\sum_{i,j}\partial_{\tau^{\prime}}\left[\kappa_{ij}(\tau^{\prime})\right]r_{ij}\partial_{\tau^{\prime}}\tau^{\prime}p_{i}(\tau^{\prime}).$
A.5 Indispensability of the Modified Current
Consider a three-state system with generator
$\bm{L}=\begin{pmatrix}-4&2&1\\ 1&-3&1\\ 3&1&-2\end{pmatrix}\,$, (185)
with eigenvalues 0,4,50,-4,-5. Let the current defining transition weights be $\kappa_{ij}(\tau)=\mathrm{e}^{4\tau}\left(\delta_{i1}\delta_{j2}-\delta_{i2}\delta_{j1}\right)$. Figure 13 shows various quality factors of the transient system with initial distribution $p_{i}(0)=(\delta_{i1}+2\delta_{i2})/3$. Explicitly, the generalized transport bound in Eq. (99), nonequilibrium steady state TUR, and the transient TUR in Eq. (78) with and without modified current ${\tilde{J}}_{t}$, see Eq. (77). As can be seen in Fig. 13, one needs ${\tilde{J}}_{t}$ in order to ensure that the bound $\mathcal{Q}\leq 1$ is not violated. To leading order $t\partial_{t}\langle{J}_{t}\rangle\sim t\mathrm{e}^{4t}$ for large $t$, while $t\partial_{t}\langle{J}_{t}\rangle-\langle{\tilde{J}}_{t}\rangle\sim\mathrm{e}^{4t}$. Additionally, $\mathrm{var}({J}_{t})\Delta S_{\mathrm{tot}}(t)\sim t\mathrm{e}^{8t}$ for large $t$, so that
$\frac{2\left(t\partial_{t}\langle{J}_{t}\rangle\right)^{2}}{\mathrm{var}({J}_{t})\Delta S_{\mathrm{tot}}(t)}$ $\sim t\,$, (186)
$\frac{2\left(t\partial_{t}\langle{J}_{t}\rangle-\langle{\tilde{J}}_{t}\rangle\right)^{2}}{\mathrm{var}({J}_{t})\Delta S_{\mathrm{tot}}(t)}$ $\sim\frac{1}{t}\,.$
Hence, the inclusion of the modified current is indeed necessary for the quality factor to remain bounded, as it otherwise diverges for $t\to\infty$, which obviously is a violation of the upper bound of the quality factor. Note that $t$ in Fig. 13 is not large enough to observe the $t^{-1}$ decay.
Figure 13: Validity and comparison of generalized transport bound and various TUR quality factors. A three-state system with generator Eq. (185) and initial condition $p_{i}(0)=(\delta_{i1}+2\delta_{i2})/3$. The transition weights is $\kappa_{ij}(\tau)=\mathrm{e}^{4\tau}\left(\delta_{i1}\delta_{j2}-\delta_{i2}\delta_{j1}\right)$. The black dashed lines are the theoretical predictions to the corresponding numerical quality factors of generalized transport bound Eq. (99) (blue), nonequilibrium steady-state TUR (green), transient TUR with ${\tilde{J}}_{t}$ (orange), and transient TUR without ${\tilde{J}}_{t}$ (red). A clear violation (shaded red region) of $\mathcal{Q}\leq 1$ is seen for the transient TUR without ${\tilde{J}}_{t}$. The inset shows the theory of the former three quality factors to better visualize their values. The numerical values are extracted from $N=5\cdot 10^{5}$ trajectories using the Gillespie algorithm.
A.6 Long-time limit of driven ring
Integrated Covariance
Here we present details on the large-$t$ limit of the stationary correlation bound in Eq. (123) applied to the driven ring Sec. IV.1 with the assumption that $\epsilon>1$. We start by realizing that $\Delta\bm{V}\equiv\bm{V}^{\mathrm{A}}-\langle{V}^{\mathrm{A}}\rangle\bm{1}=\frac{1}{2}(1,1,-1,-1)^{T}$. Hence, we can write
$\int_{0}^{t}\mathrm{d}\tau\mathrm{cov}_{\mathrm{s}}[V^{\mathrm{A}}_{\tau},V^{\mathrm{A}}_{0}]=\frac{1}{4}\int_{0}^{t}\mathrm{d}\tau\Delta\bm{V}^{T}\bm{P}(\tau)\Delta\bm{V},$ (187)
because $p_{i}^{\mathrm{s}}=1/4$. Writing the propagator as
$\bm{P}(\tau)=\begin{pmatrix}\bm{P}_{1}(\tau)&\bm{P}_{2}(\tau)\\ \bm{P}_{3}(\tau)&\bm{P}_{4}(\tau)\\ \end{pmatrix}$, (188)
with $\bm{P}_{i}(\tau)$ are $2\times 2$ matrices. Equation (187) therefore reads
$\int_{0}^{t}\mathrm{d}\tau\mathrm{cov}_{\mathrm{s}}[V^{\mathrm{A}}_{\tau},V^{\mathrm{A}}_{0}]$ (189)
$=\frac{1}{16}\int_{0}^{t}\mathrm{d}\tau\bm{1}^{T}\left[\bm{P}_{1}(\tau)+\bm{P}_{4}(\tau)-\bm{P}_{2}(\tau)-\bm{P}_{3}(\tau)\right]\bm{1}.$
All that remains is to evaluate the matrices $\bm{P}_{i}(\tau)$ and solve the integral. Consider the generator
$\bm{L}=\begin{pmatrix}-1-\epsilon&1&0&\epsilon\\ \epsilon&-1-\epsilon&1&0\\ 0&\epsilon&-1-\epsilon&1\\ 1&0&\epsilon&-1-\epsilon\end{pmatrix}$, (190)
which can be written as $\bm{L}=\bm{U}\bm{E}\bm{U}^{-1}$ with matrices
$\bm{U}$ $=\begin{pmatrix}1&i&-i&-1\\ 1&-1&-1&1\\ 1&-i&i&-1\\ 1&1&1&1\end{pmatrix}$, (191)
$\bm{E}$ $=\mathrm{diag}(0,\lambda_{4},\lambda_{3},\lambda_{2}),$ (192)
$\bm{U}^{-1}$ $=\frac{1}{4}\begin{pmatrix}1&1&1&1\\ -i&-1&i&1\\ i&-1&-i&1\\ 1&1&-1&1\end{pmatrix}$. (193)
The eigenvalues of $\bm{L}$ are listed in Eq. (126). Hence, the propagator is simply $\bm{P}(\tau)=\bm{U}\exp{\{\bm{E}\tau\}}\bm{U}^{-1}$ and Eq. (189) can explicitly be evaluated to be
$\int_{0}^{t}\mathrm{d}\tau\mathrm{cov}_{\mathrm{s}}[V^{\mathrm{A}}_{\tau},V^{\mathrm{A}}_{0}]=$ $\frac{4}{16}\int_{0}^{t}\mathrm{d}\tau\mathrm{e}^{-\tau(1+\epsilon)}\cos[\tau(\epsilon-1)]$
$=$ $\frac{1}{4}\frac{\mathrm{e}^{-\tau(1+\epsilon)}\left((1+\epsilon)\mathrm{e}^{\tau(1+\epsilon)}-(1+\epsilon)\cos[(\epsilon-1)t]+(\epsilon-1)\sin[(\epsilon-1)t]\right)}{2(1+\epsilon^{2})}$
$=$ $\frac{1}{8}\frac{1+\epsilon}{1+\epsilon^{2}}+\mathcal{O}(\mathrm{e}^{-(1+\epsilon)t}).$ (194)
(Pseudo-)Variance
Consider $\mathrm{pvar}_{\mathrm{ps}}^{F}(V^{\mathrm{A}})$ with the choice $F(\tau)=\langle V^{\mathrm{A}}_{x}+V^{\mathrm{A}}_{y}\rangle_{\mathrm{ps}}/2$, i.e., the variance $\mathrm{var}_{\mathrm{ps}}(V^{\mathrm{A}})$, for the stationary driven four-state ring. The two-point probability is
$\bm{p}^{\mathrm{ps}}=\frac{1}{4(\epsilon+1)}\begin{pmatrix}0&\epsilon&0&1\\ 1&0&\epsilon&0\\ 0&1&0&\epsilon\\ \epsilon&0&1&0\end{pmatrix}$, (195)
since $\Sigma_{\mathrm{ps}}=(\epsilon-1)^{2}/(\epsilon+1)$, $p_{x}^{\mathrm{s}}=1/4$, and, for $x\neq y$, $Z_{xy}^{2}=(\epsilon-1)^{2}/(\epsilon+1)^{2}$ Hence,
$2F=$ $\frac{1}{4(1+\epsilon)}\left(\underbrace{2\epsilon}_{x=1,y=2}+\underbrace{1}_{x=1,y=4}+\underbrace{2}_{x=2,y=1}\right$.
$+\left.\underbrace{\epsilon}_{x=2,y=3}+\underbrace{1}_{x=3,y=2}+\underbrace{\epsilon}_{x=4,y=1}\right)$
$=$ $1,$ (196)
and similarly
$\left\langle\left(V^{\mathrm{A}}_{x}+V^{\mathrm{A}}_{y}\right)^{2}\right\rangle_{\mathrm{ps}}=$ $\frac{1}{4(1+\epsilon)}\left(\underbrace{4\epsilon}_{x=1,y=2}+\underbrace{1}_{x=1,y=4}+\underbrace{4}_{x=2,y=1}\right$.
$+\left.\underbrace{\epsilon}_{x=2,y=3}+\underbrace{1}_{x=3,y=2}+\underbrace{\epsilon}_{x=4,y=1}\right)$
$=$ $1.5,$ (197)
so that
$\mathrm{var}_{\mathrm{ps}}(V^{\mathrm{A}})$ $=\left\langle\left(V^{\mathrm{A}}_{x}+V^{\mathrm{A}}_{y}\right)^{2}\right\rangle_{\mathrm{ps}}-4F^{2}=0.5.$ (198)
We may immediately recognize that $0\leq V_{x}^{\mathrm{A}}+V_{y}^{\mathrm{A}}\leq 2$, the former inequality saturates when $x,y\in\{3,4\}$ and the latter inequality saturates when $x,y\in\{1,2\}$. Thus, Popoviciu’s inequality yields
$\mathrm{var}_{\mathrm{ps}}(V^{\mathrm{A}})\leq\frac{(2-0)^{2}}{4}=1.$ (199)
A.7 Physically Motivated Rates - Pseudo Potential Form
We specifically assume the transition rates to be of the form [164]
$r_{xy}=D_{xy}\mathrm{e}^{-V_{xy}/T}=\tilde{D}_{xy}T\mathrm{e}^{-V_{xy}/T},$ (200)
where $D_{xy}$ is an edge specific “diffusion coefficient” and $V_{xy}$ is a pseudo-potential on the edge. This pseudo-potential allows for general nonequilibrium systems to be implemented and therefore allows, under given conditions, for taking the continuum limit. Temperature perturbations enter the rates as
$r_{xy}^{\delta}=\tilde{D}_{xy}(T+\delta)\mathrm{e}^{-V_{xy}/(T+\delta)}$ (201)
By expanding Eq. (201) around $\delta=0$ we get
$r_{xy}^{\delta}=r_{xy}+\tilde{D}_{xy}^{\mathrm{eff}}\mathrm{e}^{-V_{xy}/T}\delta+\mathcal{O}(\delta^{2}),$ (202)
where $\tilde{D}_{xy}^{\mathrm{eff}}\equiv\tilde{D}_{xy}(1+V_{xy}/T)$, so that $r_{xy}^{\delta}-r_{xy}=\tilde{D}_{xy}^{\mathrm{eff}}\mathrm{e}^{-V_{xy}/T}\delta+\mathcal{O}(\delta^{2})$. We can plug everything into Eq. (139) to obtain
$\partial_{\delta}O_{t}=$ $\lim_{\delta\to 0}\frac{1}{\delta}\left(\left\langle O(t)\frac{\mathrm{d}\mathds{P^{\delta}}}{\mathrm{d}\mathds{P}}\right\rangle-\langle O(t)\rangle\right)$
$=$ $\frac{1}{T^{2}}\left\langle O(t)\int_{s=0}^{s=t}\sum_{x,y\neq x}\left(V_{xy}+T\right)\mathrm{d}\varepsilon_{xy}(s)\right\rangle$ (203)
Note that we can remove the $T$ in the bracket in the last line if we assume the $D_{xy}$ not to scale with temperature, i.e., only a perturbation in the exponential in Eq. (201). By identifying the effective drift $(V_{xy}+T)$, the result Eq. (203) corresponds to the result from Ref. [136].
A.8 Perturbations of Path Observables
We already stated that Eq. (203) is valid for both, single-time and path observables. The latter, however, perhaps need some clarification in the context of evaluating expectations over path ensembles. To clarify this, consider some hollow matrix $\bm{b}$,
$O(t)=\int_{s=0}^{s=t}\mathrm{Tr}[\bm{b}^{T}\mathrm{d}\bm{\varepsilon}(s)].$ (204)
In this case, Eq. (139) simplifies to
$\partial_{\delta}O_{t}=\int_{0}^{t}\mathrm{d}s\sum_{x,y\neq x}b_{xy}\tilde{r}_{xy}p_{x}(s),$ (205)
which in general is non-zero. However, the definition of the derivative in Eq. (132) contains averages of Eq. (204). One may, falsely, assume both averages to vanish (recall that $\langle\mathrm{d}\varepsilon_{xy}(t)\rangle=0$ for all $x\neq y$ and $t\geq 0$). Only $\langle O(t)\rangle=0$, while $\langle O(t)\rangle_{\delta}\neq 0$ in general. To further understand this, we can decompose Eq. (204) into jump and dwell-time integrals using Eq. (39)
$O(t)=$ $\int_{s=0}^{s=t}\sum_{x,y\neq x}b_{xy}\mathrm{d}n_{xy}(s)$
- $\int_{s=0}^{s=t}\sum_{x,y\neq x}b_{xy}r_{xy}\mathrm{d}\tau_{x}(s).$ (206)
Taking the average w.r.t. the perturbed system, it can be recognized that
$\langle O(t)\rangle_{\delta}=$ $\left\langle\int_{s=0}^{s=t}\sum_{x,y\neq x}b_{xy}\mathrm{d}n_{xy}(s)\right\rangle_{\delta}$
$-\left\langle\int_{s=0}^{s=t}\sum_{x,y\neq x}b_{xy}r_{xy}\mathrm{d}\tau_{x}(s)\right\rangle_{\delta}$
$=$ $\int_{0}^{t}\mathrm{d}s\sum_{x,y\neq x}b_{xy}r_{xy}^{\delta}p^{\delta}_{x}(s)$
$-\int_{0}^{t}\mathrm{d}s\sum_{x,y\neq x}b_{xy}r_{xy}p^{\delta}_{x}(s)$
$=$ $\int_{0}^{t}\mathrm{d}s\sum_{x,y\neq x}b_{xy}\underbrace{(r_{xy}^{\delta}-r_{xy})}_{\delta\tilde{r}_{xy}}p^{\delta}_{x}(s)$
$=$ $\delta\int_{0}^{t}\mathrm{d}s\sum_{x,y\neq x}b_{xy}\tilde{r}_{xy}p_{x}(s)+\mathcal{O}(\delta^{2}),$ (207)
where $p^{\delta}_{x}(s)$ is the probability of $x^{\delta}_{s}=x$ in the perturbed system. The discrepancy of probabilities $\Delta p_{x}(s)=p_{x}^{\delta}(s)-p_{x}(s)$ in any state $x$ at time $s$ is $\mathcal{O}(\delta)$ [136], so that $p^{\delta}_{x}(s)=p_{x}(s)+\Delta p_{x}(s)=p_{x}(s)+\mathcal{O}(\delta)$.
Dividing Eq. (207) by $\delta$ and taking the limit $\delta\to 0$ delivers Eq. (205). In other words, it is important to account for the perturbation in the stochastic differentials only, since the observable is nominally independent of the perturbation. i.e., $b_{xy}$ and $b_{xy}r_{xy}$ in the first and second term of Eq. (206), respectively, are invariant under perturbations (even though the unperturbed rates enter the latter term).
Now let the observable $O(t)$ be a dwell-time integral
$O(t)=\int_{s=0}^{s=t}\bm{g}^{T}\mathrm{d}\bm{\tau}(s),$ (208)
with some state function vector $\bm{g}$. The response is
$\partial_{\delta}O_{t}$ $=\lim_{\delta\to 0}\frac{1}{\delta}\int_{0}^{t}\mathrm{d}s\sum_{x}g_{x}\left[p^{\delta}_{x}(s)-p_{x}(s)\right]$ (209)
Using the Dyson identity for the perturbed generator $\bm{L}^{\delta}=\bm{L}+\delta\tilde{\bm{L}}+\mathcal{O}(\delta^{2})$ with $(\tilde{\bm{L}})_{ij}=\tilde{r}_{ji}$ [161] yields
$\mathrm{e}^{\bm{L}^{\delta}t}=\mathrm{e}^{\bm{L}t}+\delta\int\mathrm{d}s\mathrm{e}^{\bm{L}(t-s)}\tilde{\bm{L}}\mathrm{e}^{\bm{L}^{\delta}s}+\mathcal{O}(\delta^{2}).$ (210)
To linear order in $\delta$, the probability is [161]
$p^{\delta}_{x}(t)$ $=p_{x}(t)+\delta\sum_{i,j,n}\int_{0}^{t}\mathrm{d}sP(x,t|i,s)\tilde{r}_{ji}P(j,s|n,0)p_{n}(0)$
$=p_{x}(t)+\delta\sum_{i,j}\int_{0}^{t}\mathrm{d}sP(x,t|i,s)\tilde{r}_{ji}p_{j}(s).$ (211)
Hence, the limit in Eq. (209) reduces to
$\lim_{\delta\to 0}\frac{1}{\delta}\left[p_{x}^{\delta}(s)-p_{x}(s)\right]=\sum_{i,j}\int_{0}^{s}\mathrm{d}zP(x,s|i,z)\tilde{r}_{ji}p_{j}(z),$ (212)
so that Eq. (209) becomes
$\partial_{\delta}O_{t}$ $=\int_{0}^{t}\mathrm{d}s\sum_{x}g_{x}\sum_{i,j}\int_{0}^{s}\mathrm{d}zP(x,s|i,z)\tilde{r}_{ji}p_{j}(z)$
$=\int_{0}^{t}\mathrm{d}s\sum_{x}g_{x}\int_{0}^{s}\mathrm{d}z\left[\sum_{i,j\neq i}P(x,s|i,z)\tilde{r}_{ji}p_{j}(z)+\sum_{i}P(x,s|i,z)\tilde{r}_{ii}p_{i}(z)\right]$
$=\int_{0}^{t}\mathrm{d}s\sum_{x}g_{x}\int_{0}^{s}\mathrm{d}z\sum_{i,j\neq i}\left[P(x,s|i,z)\tilde{r}_{ji}p_{j}(z)-P(x,s|i,z)\tilde{r}_{ij}p_{i}(z)\right]$
$=\int_{0}^{t}\mathrm{d}s\sum_{x}g_{x}\int_{0}^{s}\mathrm{d}z\sum_{i,j\neq i}\left[P(x,s|j,z)-P(x,s|i,z)\right]\tilde{r}_{ij}p_{i}(z).$ (213)
Conversely, employing Eqs. (139) and (42) yields
$\partial_{\delta}O_{t}=\sum_{i,x,y\neq x}\int_{0}^{t}\mathrm{d}\tau$ $\int_{0}^{t}\mathrm{d}s\mathds{1}_{s<\tau}g_{i}\tilde{r}_{xy}p_{x}(s)$ (214)
$\times\left[P(i,\tau|y,s)-P(i,\tau|x,s)\right],$
which agrees with the Eq. (213).
Since the jump increments $\mathrm{d}\bm{n}$ are a linear combination of the dwell-time and noise increments, it follows immediately that Eq. (139) applies for jump-integrated observables, e.g., currents, as well.
Lastly, although not explicitly written, the matrix $\bm{b}$ and vector $\bm{g}$ may also depend on time. Equations (205) and (213) are written to readily accommodate for this.
A.9 Time-Dependent Rates
The result in Eq. (139) easily allows for time-dependent rates to be included. In fact, the Radon-Nikodym derivative in Eq. (138) is already written in a form that allows to accommodate time-inhomogeneous dynamics. If $\tilde{r}_{xy}(t)$ depends on $t$, then it may even be seen as a time-dependent perturbation.
A.10 Optimization of CTUR
Figure 14: Shown are quality factor of the CTUR in (a-c) using values of $c(t)$ shown in (d-f) in the calmodulin model. The values of $c(t)=c^{*}(t)+\frac{\gamma}{5}(c^{*}(t)-c^{\dagger}(t))$ are shown for values $10\leq\gamma\leq 10$, where $\gamma=-10$ is black and $\gamma=10$ is yellow. The initial condition is $p_{i}(0)=\delta_{i6}$ and the density uses $V_{i}(\tau)=\delta_{i3}$ in each panel. The transition weights are $\bm{\kappa}^{(12)}$ in (a) and (d), $\bm{\kappa}^{(56)}$ in (b) and (e), and $\bm{\kappa}^{(26)}$ in (c) and (f). The most optimal $c(t)=c^{*}(t)$ is marked by diamonds and the least optimal $c(t)=c^{\dagger}(t)$ is marked by squares. In (a-c) the $\mathcal{Q}$ for $c=0$ is included as a blue line.
In Fig. 14 we further elaborate on the influence of $c(t)$ in the CTUR. The quality factor $\mathcal{Q}$ for the calmodulin system (see Fig. 6) for currents with $\bm{\kappa}^{(12)}$, $\bm{\kappa}^{(56)}$, and $\bm{\kappa}^{(26)}$ (recall the notation $\kappa_{ij}^{(kl)}=\delta_{ik}\delta_{jl}-\delta_{il}\delta_{jk}$) and density with $V_{i}(\tau)=\delta_{i3}$ is shown in Fig. 14a-Fig. 14c. Various values of $c(t)$ are used, including $c(t)=0$, $c^{\dagger}(t)$, and $c^{*}(t)$. The absolute values of the considered $c(t)$ are shown in Fig. 14d-Fig. 14f. As can be easily seen, the behavior of $c^{\dagger}(t)$ and $c^{*}(t)$ is highly nontrivial and changes greatly in magnitude.
A.11 Simulation Parameters
Transition $i\to j$ Rate $r_{ij}$
$1\to 2$ 5.997
$2\to 1$ 0.774
$1\to 4$ 13.439
$4\to 1$ 127.968
$1\to 5$ 15.330
$5\to 1$ 0.121
$5\to 6$ 3.749
$6\to 5$ 13.326
$2\to 3$ 1514.820
$3\to 2$ 53.0661
$2\to 6$ 13.441
$6\to 2$ 2.922
Table 2: Transition rates used in the calmodulin example from Ref. [149] which are adapted from Ref. [88].
Symbol Value
$\gamma$ 5
$x$ 2
$l_{A}$ 20
$l_{B}$ 1
$e^{\mathrm{out}}_{A}$ 30
$e^{\mathrm{in}}_{A}$ 10
$e^{\mathrm{in}}_{B}$ 2
Table 3: Values of fixed parameters in SAT simulations which are adapted from Ref. [146].
State $E_{i}$
11 $0.1$
22 $0.3$
33 $0.5$
44 $0.2$
Table 4: Overview of free energies for the four state ring with equilibrium rates.

### v2

@@
 We determine the response of path-wise observables to general (incl.
-thermal) perturbations and carry out the continuum limit to achieve the complete unification of diffusion and jump dynamics.
+thermal) perturbations and introduce a corresponding response-function formalism.
+We carry out the continuum limit to achieve the complete unification of diffusion and jump dynamics.
+In addition, we connect the framework to quantum unraveling and the Belavkin equation for open quantum systems, associating quantum and classical descriptions of thermal systems.
 Our results open new avenues in the direction of discrete-state analogs of generative diffusion models and the learning of stochastic thermodynamics from fluctuating trajectories.
@@
 Following this, we show in Sec.
-V how the stochastic calculus can be applied to derive results for the response to general perturbations for discrete-state systems and how this relates to the fluctuation-dissipation theorem in equilibrium.
+V how the stochastic calculus can be applied to derive results for the response to general perturbations for discrete-state systems, establish the corresponding response-function formalism, and show how this relates to the fluctuation-dissipation theorem in equilibrium.
 In Sec.
@@
 VII we extend the results to time-inhomogeneous dynamics.
+Subsequently, in Sec.
+VIII, we discuss how the classical stochastic-calculus approach is connected to quantum unraveling and the (discrete) Belavkin equation and state its classical analogue.
 We conclude in Sec.
-VIII with an outlook.
+IX with an outlook.
 The mathematical proofs of the noise-time correlation Lemma and increment correlations, and other technical details are given in the Appendix A.
@@
 $\mathrm{d}\bm{x}_{\tau}=\bm{F}(\bm{x}_{\tau})\mathrm{d}\tau+\bm{\sigma}\mathrm{d}\bm{W}_{\tau},$ (1)
-for the $d$-dimensional position $\bm{x}_{\tau}$ in a drift field $\bm{F}(\bm{x}_{\tau})$ subject to thermal noise $\bm{\sigma}\mathrm{d}\bm{W}_{\tau}$, where $\mathrm{d}\bm{W}_{\tau}$ are increments of the $d$-dimensional Wiener process (i.e., $d$-dimensional “white noise”).
+for the $d$-dimensional position $\bm{x}_{\tau}$ in a drift field $\bm{F}(\bm{x}_{\tau})$ subject to thermal noise $\bm{\sigma}\mathrm{d}\bm{W}_{\tau}$, where $\mathrm{d}\bm{W}_{\tau}$ are increments of the $d$-dimensional Wiener process (i.e., $d$-dimensional “white noise”), with $\langle\mathrm{d}\bm{W}_{\tau}\rangle=0$ and $\langle(\mathrm{d}\bm{W}_{\tau})_{i}(\mathrm{d}\bm{W}_{\tau^{\prime}})_{j}\rangle=\delta_{ij}\delta(\tau-\tau^{\prime})\mathrm{d}\tau\mathrm{d}\tau^{\prime}$, and $\bm{\sigma}$ is the noise amplitude assumed to have full rank (nonzero determinant).
 In Sec.
@@
 where the displacements have elements $(\mathrm{d}\bm{n}(\tau))_{xy}=\mathrm{d}n_{xy}(\tau)$ describing jumps from state $x$ to $y$ in $\tau,\tau+\mathrm{d}\tau]$.
-The matrix $(\bm{R}(x_{\tau},\tau))_{xy}\mathrm{d}\tau=\mathds{1}_{x_{\tau}=x}r_{xy}\mathrm{d}\tau$ encodes the expected (average) drift in $\tau,\tau+\mathrm{d}\tau]$ with time-independent transition rate $r_{xy}$ (following from the corresponding master equation with generator $(\bm{L})_{xy}=-\delta_{xy}\sum_{y\neq x}r_{xy}+(1-\delta_{xy})r_{yx}$), and the noise is encoded in the time-inhomogeneous shifted Poisson process $(\mathrm{d}\bm{\varepsilon}(\tau))_{xy}=\mathrm{d}\varepsilon_{xy}(\tau)$ [10, 134, 61, 60].
+The matrix $(\bm{R}(x_{\tau},\tau))_{xy}\mathrm{d}\tau={\color[rgb]{0,0,0}\delta_{x_{\tau}x}}r_{xy}\mathrm{d}\tau$ encodes the expected (average) drift in $\tau,\tau+\mathrm{d}\tau]$ with time-independent transition rate $r_{xy}$ (following from the corresponding master equation with generator $(\bm{L})_{xy}=-\delta_{xy}\sum_{y\neq x}r_{xy}+(1-\delta_{xy})r_{yx}$), and the noise is encoded in the time-inhomogeneous shifted Poisson process $(\mathrm{d}\bm{\varepsilon}(\tau))_{xy}=\mathrm{d}\varepsilon_{xy}(\tau)$ [138, 137, 38, 93].
 This noise has zero mean $\langle\mathrm{d}\varepsilon_{xy}(\tau)\rangle=0$ and the “white noise” property $\langle\mathrm{d}\varepsilon_{xy}(\tau)\mathrm{d}\varepsilon_{ij}(\tau^{\prime})\rangle_{x_{\tau}=k}=\delta_{ix}\delta_{jy}\delta_{ik}\delta(\tau-\tau^{\prime})r_{ij}\mathrm{d}\tau\mathrm{d}\tau^{\prime}$, where the expectation is conditioned on $x_{\tau}=k$.
-We prove the central time-noise correlation lemma with $\mathrm{d}\tau_{i}(\tau)=\mathds{1}_{x_{\tau}=i}\mathrm{d}\tau$,
+The Kronecker $\delta$ show that the noise increments on different edges are independent and only diagonal elements survive, e.g., $\langle\mathrm{d}\varepsilon_{ij}(\tau)\mathrm{d}\varepsilon_{ij}(\tau^{\prime})\rangle_{x_{\tau}=k}=\delta_{ik}\delta(\tau-\tau^{\prime})r_{ij}\mathrm{d}\tau\mathrm{d}\tau^{\prime}$.
+Moreover, since $|\mathrm{d}\varepsilon_{ij}|\sim\sqrt{\mathrm{d}\tau}$, higher moments scale as $|\mathrm{d}\varepsilon_{ij}|^{n}\sim(\mathrm{d}\tau)^{n/2}$, so that all higher moments and cumulants with $n\geq 3$ vanish for small $\mathrm{d}\tau$.
+Note that the Langevin Eq.
+(2) has already been introduced in [93].
+We prove the central time-noise correlation lemma with $\mathrm{d}\tau_{i}(\tau)={\color[rgb]{0,0,0}\delta_{x_{\tau}i}}\mathrm{d}\tau$,
 $\frac{\langle\mathrm{d}\varepsilon_{kl}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle_{x_{\tau}=x}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}$ (3)
-$=\delta_{xk}\mathds{1}_{\tau<\tau^{\prime}}\left[P(i,\tau^{\prime}|l,\tau)-P(i,\tau^{\prime}|k,\tau)\right]r_{kl}+\mathcal{O}(\mathrm{d}\tau),$
+$=\delta_{xk}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\left[P(i,\tau^{\prime}|l,\tau)-P(i,\tau^{\prime}|k,\tau)\right]r_{kl}+\mathcal{O}(\mathrm{d}\tau),$
 which gives access to nontrivial correlations between differentials in Eq. (2)
@@
 $\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}n_{ij}(\tau^{\prime})\rangle_{x_{\tau}=x}^{x_{\tau^{\prime}}=i}$ $=r_{ij}r_{xy}P(i,\tau^{\prime},|y,\tau)p_{x}(\tau)\mathrm{d}\tau\mathrm{d}\tau^{\prime},$
-where $P(x,\tau^{\prime}|i,\tau)$ is the transition probability (i.e., the “propagator”) from state $x$ at $\tau$ to state $i$ at $\tau\leq\tau^{\prime}$ and $p_{i}(\tau)$ is the probability to find the system in state $i$ at time $\tau$.
+where $P(x,\tau^{\prime}|i,\tau)$ is the transition probability (i.e., the “propagator”) from state $i$ at $\tau$ to state $x$ at $\tau^{\prime}\geq\tau$ and $p_{i}(\tau)$ is the probability to find the system in state $i$ at time $\tau$.
 Note that the averages are “pinned”, e.g., $\langle\cdot\rangle^{x_{\tau^{\prime}}=x}_{x_{\tau}=i}$ is the average conditioned on $x_{\tau}=i$} and $x_{\tau^{\prime}}=x$} at times $\tau\leq\tau^{\prime}$.
@@
 These correlations are required for the covariation structure of observables, such as currents and densities. Specifically, time-integrated currents are defined as [38]
-$J_{t}^{\mathrm{c}}$ $=\int_{\tau=0}^{\tau=t}\bm{U}(\bm{x}_{\tau},\tau)^{T}\circ\mathrm{d}\bm{x}_{\tau},$
+$J_{t}^{\mathrm{c}}$ $=\int_{\tau=0}^{\tau=t}\bm{U}(\bm{x}_{\tau},\tau)^{T}\circ\mathrm{d}\bm{x}_{\tau},$ (Stratonovich)
 $J_{t}$ $=\int_{\tau=0}^{\tau=t}\mathrm{Tr}[\bm{\kappa}(\tau)^{T}\mathrm{d}\bm{n}(\tau)],$ (5)
-for continuous-space and MJP, respectively, where the former is a Stratonovich integral over some vector-valued $\bm{U}(\bm{x}_{\tau},\tau)$.
-For MJP the transitions are weighted by a time-dependent antisymmetric transition weight $\bm{\kappa}(\tau)=-\bm{\kappa}(\tau)$.
+for continuous-space and MJP, respectively, where the former is a Stratonovich integral over some vector-valued $\bm{U}(\bm{x}_{\tau},\tau)$ and the $T$ denotes the transpose.
+Recall that the Stratonovich integral is a Riemann–Stieltjes-type integral defined as the limit of of functions evaluated at midpoints, e.g., $\bm{f}(\bm{x}_{\tau})^{T}\circ\mathrm{d}\bm{x}_{\tau}=\bm{f}(\frac{\bm{x}_{\tau+\mathrm{d}\tau}+\bm{x}_{\tau}}{2})^{T}(\bm{x}_{\tau+\mathrm{d}\tau}-\bm{x}_{\tau})$.
+To give an intuition for $\bm{U}$, consider the example $\bm{U}(\bm{x}_{\tau},\tau)]_{i}={\color[rgb]{0,0,0}\mathbb{1}}_{\bm{x}_{\tau}\in\{-h/2,h/2\}^{d}}$ the indicator valued vector being non-zero if $\bm{x}_{\tau}$ is in the hypercube with side length $h$, then the current “counts up” the displacement of the trajectory $\{\bm{x}_{\tau}\}_{0\leq\tau\leq t}$ in that hypercube.
+Note the superscript c throughout this manuscript refers to continuous space quantities and we write superscript (c) as a placeholder when the corresponding relation is valid in both continuous and discrete space systems.
+For MJP the transitions are weighted by a time-dependent antisymmetric transition weight $\bm{\kappa}(\tau)=-\bm{\kappa}(\tau)$ which ensures a sign change in the current when reversing time [144].
 Hence, we identify Stratonovich integrals for MJP as integrals over traces of a two-point function and the jump increments.
 The above currents naturally decompose into a stochastic and “usual” integral [38], $J^{(\mathrm{c})}_{t}=J_{t}^{\mathrm{(c),I}}+J_{t}^{\mathrm{(c),II}}$, where the stochastic part is
-$J_{t}^{\mathrm{c,I}}$ $=\int_{\tau=0}^{\tau=t}\bm{U}(\bm{x}_{\tau},\tau)^{T}\bm{\sigma}\mathrm{d}\bm{W}_{\tau}$
+$J_{t}^{\mathrm{c,I}}$ $=\int_{\tau=0}^{\tau=t}\bm{U}(\bm{x}_{\tau},\tau)^{T}\bm{\sigma}\mathrm{d}\bm{W}_{\tau}$ (Itô)
 $J_{t}^{\mathrm{I}}$ $=\int_{\tau=0}^{\tau=t}\mathrm{Tr}[\bm{\kappa}(\tau)^{T}\mathrm{d}\bm{\varepsilon}(\tau)],$ (6)
-respectively, while the ”usual” part reads $J_{t}^{\mathrm{(c),II}}=\int_{0}^{t}\mathcal{U}_{\tau}^{(c)}\mathrm{d}\tau$ where $\mathcal{U}_{\tau}^{c}\equiv\bm{U}(\bm{x}_{\tau},\tau)^{T}\bm{F}(\bm{x}_{\tau})+\nabla\cdot\left[\bm{D}\bm{U}(\bm{x}_{\tau},\tau)\right]$ and $\mathcal{U}_{\tau}\equiv\mathrm{Tr}[\bm{\kappa}(\tau)^{T}\bm{R}(x_{\tau})]$.
+respectively, while the usual” part reads $J_{t}^{\mathrm{(c),II}}=\int_{0}^{t}\mathcal{U}_{\tau}^{(c)}\mathrm{d}\tau$ where $\mathcal{U}_{\tau}^{c}\equiv\bm{U}(\bm{x}_{\tau},\tau)^{T}\bm{F}(\bm{x}_{\tau})+\nabla\cdot\left[\bm{D}\bm{U}(\bm{x}_{\tau},\tau)\right]$, $\bm{D}=\bm{\sigma}\bm{\sigma}^{T}/2$ is the diffusion matrix, and $\mathcal{U}_{\tau}\equiv\mathrm{Tr}[\bm{\kappa}(\tau)^{T}\bm{R}(x_{\tau})]$.
 Thus, we identify in the second line of Eq.
@@
 $\rho_{t}^{\mathrm{c}}$ $=\int_{0}^{t}\mathrm{d}\tau V(\bm{x}_{\tau},\tau),$
-$\rho_{t}$ $=\int_{\tau=0}^{\tau=t}\bm{V}(\tau)^{T}\mathrm{d}\bm{\tau}(\tau),$ (7)
-for continuous and discrete dynamics, respectively, where $V(\bm{x}_{\tau})$ and $\bm{V}(\tau)=(V_{1}(\tau),V_{2}(\tau),\dots)^{T}$ are scalar- and vector-valued functions, respectively.
+$\rho_{t}$ $=\int_{\tau=0}^{\tau=t}{\color[rgb]{0,0,0}V_{\tau}\mathrm{d}\tau},$ (7)
+for continuous and discrete dynamics, respectively, where $V(\bm{x}_{\tau},\tau)$ and $V_{\tau}\equiv V(x_{\tau},\tau)=\sum_{k}\delta_{x_{\tau}k}V_{k}(\tau)$ are scalar-valued functions.
 (Co)variances of stationary densities and currents can be expressed in terms of integral operators [37, 36]
 $\hat{\mathcal{I}}^{t,\mathrm{c}}_{M,N}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{\tau}^{t}\mathrm{d}\tau^{\prime}\int\mathrm{d}\bm{z}\int\mathrm{d}\bm{z}^{\prime}M(\bm{z})^{T}[\cdot]N(\bm{z}^{\prime}),$
-$\hat{{I}}^{t}_{U^{\alpha},U^{\beta}}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{\tau}^{t}\mathrm{d}\tau^{\prime}\sum_{x,y}U_{xy}^{\alpha}\sum_{i,j}U^{\beta}_{ij}[\cdot],$ (8)
+$\hat{{I}}^{t}_{{\color[rgb]{0,0,0}G}^{\alpha},{\color[rgb]{0,0,0}G}^{\beta}}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{\tau}^{t}\mathrm{d}\tau^{\prime}\sum_{x,y}{\color[rgb]{0,0,0}G}_{xy}^{\alpha}\sum_{i,j}{\color[rgb]{0,0,0}G}^{\beta}_{ij}[\cdot],$ (8)
 for continuous and discrete dynamics respectively.
 The operator in the first line of Eq (8) is already known [36, 35].
-For $X_{t}^{\mathrm{(c)}}\in\{J_{t}^{\mathrm{(c)},k},\rho_{t}^{\mathrm{(c)},k}\}$ and $Y_{t}^{\mathrm{(c)}}\in\{J_{t}^{\mathrm{(c)},l},\rho_{t}^{\mathrm{(c)},l}\}$, $k,l\in\{1,2\}$, we find
+For $X_{t}^{\mathrm{(c)}}\in\{J_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}},\rho_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}}\}$ and $Y_{t}^{\mathrm{(c)}}\in\{J_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}},\rho_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}}\}$, $k,l\in\{1,2\}$, we find
 $\mathrm{cov}(X_{t}^{\mathrm{c}},Y_{t}^{\mathrm{c}})$ $=\hat{\mathcal{I}}^{t,\mathrm{c}}_{X,Y}[{\Xi}_{m}^{\bm{z}\bm{z}^{\prime}}]-\langle X_{t}^{\mathrm{c}}\rangle\langle Y_{t}^{\mathrm{c}}\rangle,$
-$\mathrm{cov}(X_{t},Y_{t})$ $=\hat{{I}}^{t}_{U^{X},U^{Y}}[{\Xi}^{m}_{ijxy}]-\langle X_{t}\rangle\langle Y_{t}\rangle,$ (9)
-where $X_{t}^{\mathrm{(c)}},Y^{\mathrm{(c)}}_{t}$ are the appropriate pathwise observables and $U^{X},U^{Y}$ “window functions” that define them ($\bm{\kappa}$ in the case of MJP).
+$\mathrm{cov}(X_{t},Y_{t})$ $=\hat{{I}}^{t}_{{\color[rgb]{0,0,0}G}^{X},{\color[rgb]{0,0,0}G}^{Y}}[{\Xi}^{m}_{ijxy}]-\langle X_{t}\rangle\langle Y_{t}\rangle,$ (9)
+where $X_{t}^{\mathrm{(c)}},Y^{\mathrm{(c)}}_{t}$ are the appropriate pathwise observables and ${\color[rgb]{0,0,0}G}^{X},{\color[rgb]{0,0,0}G}^{Y}$ functions that define them ( e.g., $\bm{\kappa}$ in the case of currents for MJP).
 Moreover, $m\in\{1,2,3\}$ indicates density-density ($m=1$), current-density ($m=2$), and current-current ($m=3$) covariances.
-For the density covariance ($m=1$), $X_{t}^{\mathrm{(c)}}=\rho_{t}^{\mathrm{(c)},k}$ and $Y_{t}^{\mathrm{(c)}}=\rho_{t}^{\mathrm{(c)},l}$, we get
-${\Xi}_{1}^{\bm{z}\bm{z}^{\prime}}$ $=\langle\mathds{1}\rangle_{\bm{x}_{\tau}=\bm{z}}^{\bm{x}_{\tau^{\prime}}=\bm{z}^{\prime}}+\langle\mathds{1}\rangle_{\bm{x}_{\tau}=\bm{z}^{\prime}}^{\bm{x}_{\tau^{\prime}}=\bm{z}},$ (10)
+For the density covariance ($m=1$), $X_{t}^{\mathrm{(c)}}=\rho_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}}$ and $Y_{t}^{\mathrm{(c)}}=\rho_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}}$, we get
+${\Xi}_{1}^{\bm{z}\bm{z}^{\prime}}$ $=\langle{\color[rgb]{0,0,0}\mathbb{1}}\rangle_{\bm{x}_{\tau}=\bm{z}}^{\bm{x}_{\tau^{\prime}}=\bm{z}^{\prime}}+\langle{\color[rgb]{0,0,0}\mathbb{1}}\rangle_{\bm{x}_{\tau}=\bm{z}^{\prime}}^{\bm{x}_{\tau^{\prime}}=\bm{z}},$ (10)
 ${\Xi}^{1}_{ijxy}$ $=\frac{\langle\mathrm{d}\tau_{x}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle_{x_{\tau}=x}^{x_{\tau^{\prime}}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}+\frac{\langle\mathrm{d}\tau_{i}(\tau)\mathrm{d}\tau_{x}(\tau^{\prime})\rangle^{x_{\tau^{\prime}}=x}_{x_{\tau}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}},$
@@
 For MJP we find, using $\int_{0}^{t}\mathrm{d}\tau\int_{0}^{\tau}\mathrm{d}\tau^{\prime}\to\int_{0}^{t}\mathrm{d}t^{\prime}(t-t^{\prime})$ [36], that
-$\mathrm{cov}({\rho}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P_{\mathrm{s}}(x,\tau;i,\tau^{\prime})+P_{\mathrm{s}}(x,\tau^{\prime};i,\tau)-p_{x}^{\mathrm{s}}p_{i}^{\mathrm{s}}\right]$
-$\mathrm{cov}({J}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\hat{j}_{xy}P(x,t^{\prime}|i)p_{i}^{\mathrm{s}}+\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}p_{i}^{\mathrm{s}}\right],$
-$\mathrm{cov}(J_{t}^{k},{J}_{t}^{l})=$ $t\sum_{x,y}\kappa_{xy}^{k}\kappa_{xy}^{l}r_{xy}p_{x}^{\mathrm{s}}+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\hat{j}_{xy}\hat{j}_{ji}^{\ddagger}P(x,t^{\prime}|j)p_{j}^{\mathrm{s}}+\hat{j}_{ij}\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}J^{\mathrm{s}}_{ij}\right],$ (13)
+$\mathrm{cov}({\rho}_{t,{\color[rgb]{0,0,0}k}},{\rho}_{t,{\color[rgb]{0,0,0}l}})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P_{\mathrm{s}}(x,\tau;i,\tau^{\prime})+P_{\mathrm{s}}(x,\tau^{\prime};i,\tau)-p_{x}^{\mathrm{s}}p_{i}^{\mathrm{s}}\right]$
+$\mathrm{cov}({J}_{t,{\color[rgb]{0,0,0}k}},{\rho}_{t,{\color[rgb]{0,0,0}l}})=$ $\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\hat{j}_{xy}P(x,t^{\prime}|i)p_{i}^{\mathrm{s}}+\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}p_{i}^{\mathrm{s}}\right],$
+$\mathrm{cov}(J_{t,{\color[rgb]{0,0,0}k}},{J}_{t,{\color[rgb]{0,0,0}l}})=$ $t\sum_{x,y}\kappa_{xy}^{k}\kappa_{xy}^{l}r_{xy}p_{x}^{\mathrm{s}}+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\hat{j}_{xy}\hat{j}_{ji}^{\ddagger}P(x,t^{\prime}|j)p_{j}^{\mathrm{s}}+\hat{j}_{ij}\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}J^{\mathrm{s}}_{ij}\right],$ (13)
 where $V^{\alpha/\beta}\delta$ is a shorthand notation for $V^{\alpha/\beta}_{i}\delta_{ij}$, $J^{\mathrm{s}}_{xy}$ is the local steady-state current between $x$ and $y$, and $\hat{j}_{xy}^{\ddagger}$ is the so-called dual-reverse current operator (reversed steady-state current) [37, 36, 35] which for MJP reads $\hat{j}_{xy}=r_{xy}\to\hat{j}_{xy}^{\ddagger}=r_{yx}p_{y}^{\mathrm{s}}/p_{x}^{\mathrm{s}}$ [144].
+In Eq.
+(13), the joint stationary probabilities for state $x$ at $\tau$ and $i$ at $\tau^{\prime}$ are denoted $P_{\mathrm{s}}(x,\tau;i,\tau^{\prime})$, while the marginal stationary distributions to be in a state $x$ is $p^{\mathrm{s}}_{x}$.
+Additionally, we use the fact that the propagator for time-homogeneous systems is time-translation invariant $P(i,\tau^{\prime}|x,\tau)=P(i,\tau^{\prime}-\tau|x)$.
 The results for continuous-space diffusion are given in Refs.
@@
 $\leq\frac{\Delta S_{\mathrm{tot}}^{\mathrm{(c)}}(t)}{2},$ (14)
-where $c(t):\mathds{R}\to\mathds{R}$ is a deterministic weight function for the density and $\Delta S_{\mathrm{tot}}^{\mathrm{(c)}}$ is the discrete (continuous) space entropy production.
-In analogy with [61] we introduced the discrete space modified time-integrated density and current $\tilde{\rho}_{t}=\int_{\tau=0}^{\tau=t}\tau\partial_{\tau}\bm{V}(\tau)^{T}\mathrm{d}\bm{\tau}(\tau)$ and $\tilde{J}_{t}=\int_{\tau=0}^{\tau=t}\tau\mathrm{Tr}\left[(\partial_{\tau}\bm{\kappa}(\tau))^{T}\mathrm{d}\bm{n}(\tau)\right]$ that account for the dependence of $\bm{V}(\tau)$ and $\bm{\kappa}(\tau)$ on $\tau$.
+where $c(t):{\color[rgb]{0,0,0}\mathbb{R}}\to{\color[rgb]{0,0,0}\mathbb{R}}$ is a deterministic weight function for the density and $\Delta S_{\mathrm{tot}}^{\mathrm{(c)}}$ is the discrete (continuous) space entropy production [143]
+$\Delta S_{\mathrm{tot}}^{\mathrm{c}}(t)$ $=\int_{0}^{t}\mathrm{d}\tau\int\mathrm{d}\bm{x}\frac{\bm{j}(\bm{x},\tau)^{T}\bm{D}^{-1}\bm{j}(\bm{x},\tau)}{P(\bm{x},\tau)},$ (15)
+$\Delta S_{\mathrm{tot}}(t)$ $=\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}r_{ij}p_{i}(\tau)\log\frac{r_{ij}p_{i}(\tau)}{r_{ji}p_{j}(\tau)}.$ (16)
+Here, $\bm{j}(\bm{x},\tau)=(\bm{F}(\bm{x})-\bm{D}\nabla)P(\bm{x},\tau)$ is the probability current entering the Fokker Planck equation.
+In analogy with [38] we introduced the discrete space modified time-integrated density and current $\tilde{\rho}_{t}=\int_{\tau=0}^{\tau=t}\tau\partial_{\tau}{\color[rgb]{0,0,0}V_{\tau}\mathrm{d}{\tau}}$ and $\tilde{J}_{t}=\int_{\tau=0}^{\tau=t}\tau\mathrm{Tr}\left[(\partial_{\tau}\bm{\kappa}(\tau))^{T}\mathrm{d}\bm{n}(\tau)\right]$ that account for the dependence of $V_{i}(\tau)$ and $\bm{\kappa}(\tau)$ on $\tau$.
+To compare, the continuous space modified current and density read $\tilde{J}_{t}^{\mathrm{c}}=\int_{\tau=0}^{\tau=t}\tau\partial_{\tau}\bm{U}(\bm{x}_{\tau},\tau)^{T}\circ\mathrm{d}\bm{x}_{\tau}$ and $\tilde{\rho}_{t}^{\mathrm{c}}=\int_{0}^{t}\tau\partial_{\tau}V(\bm{x}_{\tau},\tau)\mathrm{d}\tau$, respectively.
+The transient correlation TUR for MJP in Eq.
+(14) has not yet been proven in this generality since previous results were restricted to steady states [32] or did not consider correlations [89].
 Setting $c(t)=0$ we recover the transient TUR.
-The CTUR (in contrast to the TUR) allows to turn the Cauchy-Schwarz inequality into an equality for all times given a suitable choice of $\bm{V}(\tau)$ and $\bm{\kappa}(\tau)$.
+The CTUR (in contrast to the TUR) allows to turn the Cauchy-Schwarz inequality into an equality for all times given a suitable choice of $V_{\tau}$ and $\bm{\kappa}(\tau)$.
 This choice, however, requires explicit and detailed information about the system and is therefore not applicable in most realistic applications.
 We can nevertheless optimize Eq.
-(14) with respect to $c(t)$ for given $\bm{V}(\tau)$ and $\bm{\kappa}(\tau)$ to find
+(14) with respect to $c(t)$ for given $V_{\tau}$ and $\bm{\kappa}(\tau)$ to find
 $c^{*}(t)$ $=\frac{a(t)\mathrm{cov}({\rho}_{t},{J}_{t})-b(t)\mathrm{var}({J}_{t})}{a(t)\mathrm{var}({\rho}_{t})-b(t)\mathrm{cov}({\rho}_{t},{J}_{t})}\,$, (17)
@@
 Interestingly, in contrast to the steady-state result [32], choosing an arbitrary $c(t)$ may even deteriorate the inference compared to the bare current or density TURs.
-The second thermodynamic inequality we prove is the thermodynamic bound on the transport of any scalar observable $z(t)\equiv z(x_{\tau},\tau)$ [51]
-$\frac{\langle z(t)-z(0)-\int_{0}^{t}\mathrm{d}\tau\partial_{\tau}z(\tau)\rangle^{2}}{t\mathcal{D}^{\mathrm{(c)}}_{t}}\leq\frac{\Delta S_{\mathrm{tot}}^{\mathrm{(c)}}}{2}\,$, (16)
-where $\mathcal{D}^{\mathrm{(c)}}_{t}$ is the so-called fluctuation-scale function accounting for how much the observation stretches microscopic coordinates.
-In continuous space it was shown [113, 51] that
-$\mathcal{D}_{t}^{\mathrm{c}}=\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\left\langle\left[\nabla_{\bm{x}}z_{\tau}\right]^{T}\bm{\gamma}^{-1}(\tau)\nabla_{\bm{x}}z_{\tau}\right\rangle.$ (17)
-For MJP we in turn find
-$\mathcal{D}_{t}\equiv\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\sum_{x,y}(z_{y}(\tau)-z_{x}(\tau))^{2}r_{xy}p_{x}(\tau).$ (18)
+The second thermodynamic inequality we prove is the thermodynamic bound on the transport of any scalar observable $z_{{\color[rgb]{0,0,0}\tau}}\equiv z(x_{\tau},\tau)$ [40]
+$\frac{\langle z_{{\color[rgb]{0,0,0}t}}-z_{{\color[rgb]{0,0,0}0}}-\int_{0}^{t}\mathrm{d}\tau\partial_{\tau}z_{{\color[rgb]{0,0,0}\tau}}\rangle^{2}}{t\mathcal{D}^{\mathrm{(c)}}_{t}}\leq\frac{\Delta S_{\mathrm{tot}}^{\mathrm{(c)}}}{2}\,$, (18)
+where in continuous space it was shown [30, 40] that
+$\mathcal{D}_{t}^{\mathrm{c}}=\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\left\langle\left[\nabla_{\bm{x}}z_{\tau}\right]^{T}\bm{\gamma}^{-1}(\tau)\nabla_{\bm{x}}z_{\tau}\right\rangle,$ (19)
+while for MJP we find
+$\mathcal{D}_{t}\equiv\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\sum_{x,y}[z_{y}(\tau)-z_{x}(\tau)]^{2}r_{xy}p_{x}(\tau).$ (20)
+Here, $\mathcal{D}^{\mathrm{(c)}}_{t}$ is the so-called fluctuation-scale function accounting for how much the observation stretches microscopic coordinates.
+For example, note that stretching $z_{\tau}\mapsto az_{\tau}$ for some constant $a\in\mathbb{R}$ does not alter Eq.
+(18) as the stretch in the nominator is canceled by the change in $\mathcal{D}^{\mathrm{(c)}}_{t}$.
+In Eq.
+(20), $z_{x}(\tau)$ is the specific value of the function $z_{\tau}$ if $x_{\tau}=x$, i.e., $z_{\tau}\equiv z(x_{\tau},\tau)=\sum_{k}\delta_{x_{\tau}k}z_{k}(\tau)$.
 The fluctuation-scale function for both, continuous and discrete space dynamics, is operationally accessible by considering increments $\mathrm{d}z(\tau)=z(x_{\tau+\mathrm{d}\tau},\tau+\mathrm{d}\tau)-z(x_{\tau},\tau)$ along a trajectory and recognizing (as in continuous space [30, 40]) that
@@
 [30].
-It involves correlations of time-independent observables $V_{\tau}\equiv V(x_{\tau})=\sum_{k}\delta_{k,x_{\tau}}V_{k}$ and $z_{\tau}\equiv z(x_{\tau})=\sum_{k}\delta_{k,x_{\tau}}z_{k}$ and their time-average along a trajectory, e.g., $\overline{V}_{t}=t^{-1}\int_{0}^{t}V_{\tau}\mathrm{d}\tau$, resulting in a lower bound of the integrated entropy production rate $\dot{S}_{\mathrm{tot}}(\tau)$ weighted by a “pseudo variance” $\mathrm{pvar}_{\mathrm{ps}}^{F}(z_{\tau})=\sum_{x,y}p_{xy}^{\mathrm{ps}}(\tau)(z_{x}+z_{y}-2F(\tau))$ [114]
+It involves correlations of time-independent observables $V_{\tau}\equiv V(x_{\tau})=\sum_{k}{\color[rgb]{0,0,0}\delta_{x_{\tau}k}}V_{k}$ and $z_{\tau}\equiv z(x_{\tau})=\sum_{k}{\color[rgb]{0,0,0}\delta_{x_{\tau}k}}z_{k}$ and their time-average along a trajectory, e.g., $\overline{V}_{t}=t^{-1}\int_{0}^{t}V_{\tau}\mathrm{d}\tau$, resulting in a lower bound of the integrated entropy production rate $\dot{S}_{\mathrm{tot}}(\tau)$ weighted by a “pseudo variance” $\mathrm{pvar}_{\mathrm{ps}}^{F}(z_{\tau})=\sum_{x,y}p_{xy}^{\mathrm{ps}}(\tau)(z_{x}+z_{y}-2F(\tau))^{{\color[rgb]{0,0,0}2}}$ [42]
 $\frac{\left[2\mathrm{cov}({V}_{t},\overline{{z}}_{t})-\frac{2}{t}\int_{0}^{t}\mathrm{d}\tau\mathrm{cov}(V_{\tau},z_{\tau})\right]^{2}}{t\mathrm{var}(\overline{V}_{t})}-\mathcal{D}_{t}+\frac{2}{t}\left[\mathrm{var}(z_{t})-\mathrm{var}(z_{0})\right]\leq\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\dot{S}_{\mathrm{tot}}(\tau)\mathrm{pvar}_{\mathrm{ps}}^{F}(z_{\tau}),$ (22)
-where $F(\tau)$ may be any time-(in)dependent function.
-The tilted two-point probability $p_{xy}^{\mathrm{ps}}(\tau)$ requires knowledge about transition rates, hence the pseudo variance is not operationally accessible.
+where $F(\tau)$ may be any time-(in)dependent function. The tilted two-point probability,
+$p_{xy}^{\mathrm{ps}}(\tau)=\frac{p_{x}(\tau)r_{xy}Z_{xy}(\tau)^{2}}{\Sigma_{\mathrm{ps}}},$ (23)
+where $Z_{xy}(s)=[r_{xy}p_{x}(\tau)-r_{yx}p_{y}(\tau)]/[r_{xy}p_{x}(\tau)+r_{yx}p_{y}(\tau)]$ and $\Sigma_{\mathrm{ps}}=\sum_{x,y}p_{x}(\tau)r_{xy}Z_{xy}(\tau)^{2}$, requires knowledge about transition rates, hence the pseudo variance is not operationally accessible.
 However, it can be controlled by choosing bounded observables.
+From the expression Eq.
+(23), we can see that $p_{xy}^{\mathrm{ps}}(\tau)$ tilts the probability weight onto dissipating transitions.
+Hence, the pseudo-variance projects the observable onto dissipating transitions.
+Further, we emphasize that the pseudo-variance is a mathematical construct, which needs to be bounded further, hence the physical interpretation is not relevant.
 In NESS, Eq.
@@
 Response to External Perturbations
-The response of a single-time observable $O(t)$ of a continuous-space diffusion to a perturbation in the noise amplitude, which effectively is a temperature perturbation [135], $\bm{\sigma}\to\bm{\sigma}+\delta\tilde{\bm{\sigma}}$ was recently derived using stochastic calculus and reads [136]
+The response of a single-time observable $O(t)$ of a continuous-space diffusion to a perturbation in the noise amplitude, which effectively is a temperature perturbation [91], $\bm{\sigma}\to\bm{\sigma}+\delta\tilde{\bm{\sigma}}$ with perturbation noise amplitude $\tilde{\bm{\sigma}}$ was recently derived using stochastic calculus and reads [86]
 $\partial_{\delta}O_{t}$ $\equiv\lim_{\delta\to 0}\frac{1}{\delta}\left(\langle O(t)\rangle_{\delta}-\langle O(t)\rangle\right)$ (26)
@@
 $r_{xy}^{\delta}$ $=r_{xy}+\delta\tilde{r}_{xy}+\mathcal{O}(\delta^{2}).$ (31)
+Note that similar results have recently been obtained in with a martingale approach [46].
 For equilibrium systems where
@@
 (29) can in some sense be seen as a (mathematical) extension of the FDT to irreversible and transient systems.
-In Appendix A.8 we generalize the response in Eq.
-(26) to perturbations of pathwise observables of MJP, e.g.
-$O(t)=\int_{s=0}^{s=t}\mathrm{Tr}[\bm{b}(s)^{T}\mathrm{d}\bm{\varepsilon}(s)],$
-for some time-dependent hollow matrix $\bm{b}(s)$.
+We further generalize the response in Eq. (29) to perturbations of pathwise observables of MJP (see Appendix A.8). Without loss of generality we set
+$O_{{\color[rgb]{0,0,0}1}}(t)$ $=\int_{s=0}^{s=t}\mathrm{Tr}[\bm{b}(s)^{T}\mathrm{d}\bm{\varepsilon}(s)],$
+${\color[rgb]{0,0,0}O_{2}(t)}$ ${\color[rgb]{0,0,0}=\int_{0}^{t}g_{s}\mathrm{d}s,}$
+for some time-dependent hollow matrix $\bm{b}(s)$ and state function $g_{s}\equiv\sum_{x}\delta_{x_{s}x}g_{x}(s)$.
+For a perturbation strength $h(s)$, i.e., $r_{xy}\to r_{xy}+h(s)\tilde{r}_{xy}$, the result Eq.
+(29) for $O_{\rm p}(t)=O_{1}(t)+O_{2}(t)$ and its path-wise extension in Sec.
+V.3.3 can be recast into the familiar response-function formalism
+$\Delta O_{\rm p}(t)$ $\equiv\langle O_{\rm p}(t)\rangle_{h}-\langle O_{\rm p}(t)\rangle$
+$=\int_{0}^{t}\mathrm{d}sh(s)\chi_{O_{\rm p}}(t,s)+\mathcal{O}(h^{2}),$ (34)
+with response function $\chi_{O_{\rm p}}(t,s)$
+$\chi_{O_{\rm p}}(t,s)$ $=\chi_{O_{1}}(t,s)+\chi_{O_{2}}(t,s)$ (35)
+$={\color[rgb]{0,0,0}\mathbb{1}}_{s<t}\sum_{x,y\neq x}b_{xy}(s)\tilde{r}_{xy}p_{x}(s)+\int_{0}^{t}\mathrm{d}{\color[rgb]{0,0,0}\mathbb{1}}_{z>s}\langle g_{z}R(s)\rangle\,$,
+with $R_{x}(s)=\sum_{y}\tilde{r}_{yx}p_{y}(s)/p_{x}(s)$.
 The results in Eqs.
@@
 when the diagonalization of the transition matrix becomes computationally too demanding.
+Stochastic Calculus and Open Quantum Systems
+Open discrete quantum systems coupled to a thermal bath are described by the Belavkin equation [27]
+$\mathrm{d}\rho_{s}^{\mathrm{u}}=\mathcal{B}(\rho_{s}^{\mathrm{u}})\mathrm{d}s+\sum_{i}\left(\frac{\mathcal{J}_{i}(\rho_{s}^{\mathrm{u}})}{\mathrm{Tr}[\mathcal{J}_{i}(\rho_{s}^{\mathrm{u}})]}-\rho_{s}^{\mathrm{u}}\right)\mathrm{d}\tilde{n}_{is},$ (36)
+where the pure state $\rho_{s}^{\mathrm{u}}$ is a rank-1 matrix and $\mathcal{J}_{i}(\rho_{s}^{\mathrm{u}})=\hat{J}_{i}\rho_{s}^{\mathrm{u}}\hat{J}_{i}^{\dagger}$ with jump operator $\hat{J}_{i}$ for the $i$th quantum jump.
+The deterministic drift in Eq.
+(36) has the form
+$\mathcal{B}(\rho_{s}^{\mathrm{u}})=-i\hat{H}_{\mathrm{eff}}\rho_{s}^{\mathrm{u}}+\rho_{s}^{\mathrm{u}}\hat{H}_{\mathrm{eff}}^{\dagger}-\rho_{s}^{\mathrm{u}}\mathrm{Tr}(-i\hat{H}_{\mathrm{eff}}\rho_{s}^{\mathrm{u}}+\rho_{s}^{\mathrm{u}}\hat{H}_{\mathrm{eff}}^{\dagger}),$ (37)
+where the the effective Hamiltonian $\hat{H}_{\mathrm{eff}}=\hat{H}-\frac{i}{2}\sum_{j}\hat{J}_{j}^{\dagger}\hat{J}_{j}$ includes the Hamiltonian $\hat{H}$ and $\mathrm{d}\tilde{n}_{is}$ is a stochastic differential which takes values 1 if the $i$th transition occurs in $s,s+\mathrm{d}s]$ and 0 otherwise.
+The statistical properties of the noise are $\langle\mathrm{d}\tilde{n}_{is}\rangle=\mathrm{Tr}(\mathcal{J}_{i}(\rho_{s}))$ and $\mathrm{d}\tilde{n}_{is}\mathrm{d}\tilde{n}_{js}=\delta_{ij}\mathrm{d}\tilde{n}_{is}$.
+We show that the classical analogue to Eq. (36) reads
+$\mathrm{d}\delta_{x_{s}x}=$ $\sum_{y\neq x}\left(r_{yx}\delta_{x_{s}y}\mathrm{d}s-\langle\mathrm{d}n_{yx}\rangle-r_{xy}\delta_{x_{s}x}\mathrm{d}s+\langle\mathrm{d}n_{xy}\rangle\right)$
+$+\sum_{y\neq x}(\mathrm{d}\varepsilon_{yx}+\langle\mathrm{d}n_{yx}\rangle-\mathrm{d}\varepsilon_{xy}-\langle\mathrm{d}n_{xy}\rangle).$ (38)
+Taking the average over all noise histories in Eqs.
+(36) and (38) recovers the Lindblad and master equation, respectively.
+Hence, Eqs.
+(36) and (38) are the stochastic evolution of states in the system.
+These, as well as other functionals of stochastic increments, we can study as we know the underlying equations of motion, e.g., Eq.
+(2).
 In the following section we provide detailed definitions and derivations of the above results and present applications. Technical steps and Lemmas are deferred to the Appendix.
@@
 To facilitate the comparison between discrete and continuous state spaces we briefly summarize the basics of overdamped (for simplicity additive-noise) diffusions.
-Let $(\bm{x}_{\tau})_{0\leq\tau\leq t}$ be a trajectory of $\bm{x}_{\tau}\in\mathds{R}^{d}$ evolving according to overdamped Langevin dynamics [61]
+Let $(\bm{x}_{\tau})_{0\leq\tau\leq t}$ be a trajectory of $\bm{x}_{\tau}\in{\color[rgb]{0,0,0}\mathbb{R}}^{d}$ evolving according to overdamped Langevin dynamics [51, 144, 86]
 $\mathrm{d}\bm{x}_{\tau}=\bm{F}(\bm{x}_{\tau})\mathrm{d}\tau+\bm{\sigma}\mathrm{d}\bm{W}_{\tau},$ (39)
@@
 The colors in (d) correspond to the different transitions in (b).
-Consider now a discrete state space $\mathcal{N}\subseteq\mathds{N}$ which, without loss of generality, we can enumerate as $\mathcal{N}=\{1,2,\dots,N\}$.
+Consider now a discrete state space $\mathcal{N}\subseteq{\color[rgb]{0,0,0}\mathbb{N}}$ which, without loss of generality, we can enumerate as $\mathcal{N}=\{1,2,\dots,N\}$.
 Focusing for the time being on time-homogeneous systems (i.e., time-independent transition rates $r_{ij}$ between states $i$ and $j$) the state probability vector $p_{i}(\tau)$ evolves according to the master equation [140, 145]
@@
 where the summation is only over $j$.
-In terms in the incoming and outgoing fluxes $J_{xy}(\tau)=r_{xy}p_{x}(\tau)$ we have as $\frac{\mathrm{d}}{\mathrm{d}\tau}p_{i}(\tau){\color[rgb]{0,0,1}=}\sum_{j\neq i}[J_{ij}(\tau)-J_{ji}(\tau)]$, and in vector notation $\frac{\mathrm{d}}{\mathrm{d}\tau}\bm{p}(\tau)=\bm{L}\bm{p}(\tau)$ with generator $\bm{L}$ having entries $L_{ij}=\delta_{ij}r_{ii}+(1-\delta_{ij})r_{ji}$ and exit rate $r_{ii}=-\sum_{j\neq i}r_{ij}$.
+In terms in the incoming and outgoing fluxes $J_{xy}(\tau)=r_{xy}p_{x}(\tau)$ we have as $\frac{\mathrm{d}}{\mathrm{d}\tau}p_{i}(\tau){\color[rgb]{0,0,0}=}\sum_{j\neq i}[J_{ij}(\tau)-J_{ji}(\tau)]$, and in vector notation $\frac{\mathrm{d}}{\mathrm{d}\tau}\bm{p}(\tau)=\bm{L}\bm{p}(\tau)$ with generator $\bm{L}$ having entries $L_{ij}=\delta_{ij}r_{ii}+(1-\delta_{ij})r_{ji}$ and exit rate $r_{ii}=-\sum_{j\neq i}r_{ij}$.
 The transition probability (or “propagator”) is therefore given by $P(i,\tau^{\prime}-\tau|j)=P(i,\tau^{\prime}|j,\tau)=\left(\exp{(\tau^{\prime}-\tau)\bm{L}}\right)_{ij}$.
 A path is fully determined by the sequence of states $(x_{i})$ and transition times $t_{i}$, $(x_{\tau})_{0\leq\tau\leq t}=(x_{0},t_{0}=0;x_{1},t_{1};\dots)$.
-The probability of such a trajectory is then [6, 36, 95, 138, 139]
-$\mathds{P}\left[(x_{\tau})_{0\leq\tau\leq t}\right]=p_{x_{0}}(0)\mathrm{e}^{\sum_{k}\tau_{k}(t)r_{kk}}\prod_{i,j}(r_{ij})^{n_{ij}(t)},$ (33)
+The normalized probability of such a trajectory is then [144, 109, 95, 69, 147]
+${\color[rgb]{0,0,0}\mathbb{P}}\left[(x_{\tau})_{0\leq\tau\leq t}\right]=p_{x_{0}}(0)\mathrm{e}^{\sum_{k}\tau_{k}(t)r_{kk}}\prod_{i,j}(r_{ij})^{n_{ij}(t)},$ (41)
 where $p_{x_{0}}(0)$ is the probability of the initial state $x_{0}$ and $n_{ij}(t)$ and $\tau_{k}(t)$ are the number of transitions $i\to j$ and the total time spent in state $k$ along the trajectory of length $t$, respectively, and can be formulated as [38]
@@
 Rearranging Eq. (45) yields the consistent stochastic equation of motion for $x\neq y$ for MJP
-$\underbrace{\mathrm{d}n_{xy}(\tau)}_{\text{``displacement"}}=\underbrace{r_{xy}\mathds{1}_{x_{\tau}=x}}_{\text{``drift"}}\mathrm{d}\tau+\underbrace{\mathrm{d}\varepsilon_{xy}(\tau)}_{\text{``noise"}}.$ (38)
-Defining the matrix $(\bm{R}(x_{\tau}))_{ij}\equiv(1-\delta_{ij})r_{ij}\mathds{1}_{x_{\tau}=i}$, we can write Eq. (38) as a stochastic matrix equation
+$\underbrace{\mathrm{d}n_{xy}(\tau)}_{\text{``displacement"}}=\underbrace{r_{xy}{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}}_{\text{``drift"}}\mathrm{d}\tau+\underbrace{\mathrm{d}\varepsilon_{xy}(\tau)}_{\text{``noise"}}.$ (46)
+Defining the matrix $(\bm{R}(x_{\tau}))_{ij}\equiv(1-\delta_{ij})r_{ij}{\color[rgb]{0,0,0}\delta_{x_{\tau}i}}$, we can write Eq. (46) as a stochastic matrix equation
 $\mathrm{d}\bm{n}(\tau)=\bm{R}(x_{\tau})\mathrm{d}\tau+\mathrm{d}\bm{\varepsilon}(\tau).$ (47)
@@
 (47) are to be neglected, as they do not have any physically relevant meaning.
+Table 1: Overview of quantities entering the stochastic equations of motion in continuous and discrete space.
+Meaning Continuous Discrete
+Displacement $\mathrm{d}\bm{x}_{\tau}$ $\mathrm{d}\bm{n}(\tau)$
+Deterministic Drift $\bm{F}(\bm{x}_{\tau})\mathrm{d}\tau$ $\bm{R}(x_{\tau})\mathrm{d}\tau$
+Noise $\bm{\sigma}\mathrm{d}\bm{W}_{\tau}$ $\mathrm{d}\bm{\varepsilon}(\tau)$
 Note that $\mathrm{d}\bm{\varepsilon}(\tau)$ is not Gaussian noise, but rather a shifted (due to the $\bm{R}(x_{\tau})\mathrm{d}\tau$ term) time-inhomogeneous Poisson process.
 Explicitly, because for $x_{\tau}=x$ and $y\neq x$ the differential $\mathrm{d}n_{xy}(\tau)\sim\mathrm{Poi}(r_{xy}\mathrm{d}\tau)$ is Poisson, the values $\mathrm{d}\varepsilon_{xy}(\tau)$ can take for any $x_{\tau}$ are $\mathrm{d}\varepsilon_{xy}(\tau)\in\{0,-r_{xy}\mathrm{d}\tau,1-r_{xy}\mathrm{d}\tau,2-r_{xy}\mathrm{d}\tau,3-r_{xy}\mathrm{d}\tau,\dots\}$, where $\mathrm{d}\varepsilon_{xy}(\tau)=0$ happens if $x_{\tau}\neq x$.
-In case $x_{\tau}=x$, one can identify $P(\mathrm{d}\varepsilon_{xy}(\tau)=k-r_{xy}\mathrm{d}\tau|x_{\tau}=x)=\mathrm{poi}_{r_{xy}\mathrm{d}\tau}(k)$ for any $k\in\mathds{N}_{\geq 0}$.
-In the limit of small $\mathrm{d}\tau$ to linear order in $\mathrm{d}\tau$ the noise therefore takes values $\mathrm{d}\varepsilon_{xy}(\tau)\in\{0,-r_{xy}\mathrm{d}\tau,1-r_{xy}\mathrm{d}\tau\}$ with probabilities $1-p_{x}(\tau),\left(1-r_{xy}\mathrm{d}\tau\right)p_{x}(\tau),r_{xy}\mathrm{d}\tau p_{x}(\tau)$, respectively.
+In case $x_{\tau}=x$, one can identify $P(\mathrm{d}\varepsilon_{xy}(\tau)=k-r_{xy}\mathrm{d}\tau|x_{\tau}=x)=\mathrm{poi}_{r_{xy}\mathrm{d}\tau}(k)$ for any $k\in{\color[rgb]{0,0,0}\mathbb{N}}_{\geq 0}$, where $\mathrm{poi}_{\lambda}(k)=\lambda^{k}\mathrm{e^{-\lambda}/k!}$ is the Poisson distribution function.
+In the limit of small $\mathrm{d}\tau$, to linear order in $\mathrm{d}\tau$, the noise therefore takes values $\mathrm{d}\varepsilon_{xy}(\tau)\in\{0,-r_{xy}\mathrm{d}\tau,1-r_{xy}\mathrm{d}\tau\}$ with probabilities $1-p_{x}(\tau),\left(1-r_{xy}\mathrm{d}\tau\right)p_{x}(\tau),r_{xy}\mathrm{d}\tau p_{x}(\tau)$, respectively.
 The probability for any $k\geq 2$ are $\mathcal{O}(\mathrm{d}\tau^{k})$ and can therefore be neglected.
@@
 $\frac{\langle\mathrm{d}\varepsilon_{kl}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle_{x_{\tau}=x}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}$ (50)
-$=\delta_{xk}\mathds{1}_{\tau<\tau^{\prime}}\left[P(i,\tau^{\prime}|l,\tau)-P(i,\tau^{\prime}|k,\tau)\right]r_{kl}+\mathcal{O}(\mathrm{d}\tau).$
+$=\delta_{xk}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\left[P(i,\tau^{\prime}|l,\tau)-P(i,\tau^{\prime}|k,\tau)\right]r_{kl}+\mathcal{O}(\mathrm{d}\tau).$
 The above lemma has been shown before [93]; in App.
@@
 Note that, due to the Markovianty of the system, we do not need to restrict on $\tau\leq\tau^{\prime}$ in evaluating the expectation, as $\tau>\tau^{\prime}$ vanishes.
-The above correlators allow, for example, to evaluate the covariance between the number of jumps ate different space-time points along the trajectory, i.e., for $\tau\leq\tau^{\prime}$, [60]
+The above correlators allow, for example, to evaluate the covariance between the number of jumps at different space-time points along the trajectory, i.e., for $\tau\leq\tau^{\prime}$, [93]
 $\langle$ $\mathrm{d}n_{kl}(\tau)\mathrm{d}n_{ij}(\tau^{\prime})\rangle_{x_{\tau}=x}=\mathrm{d}\tau\mathrm{d}\tau^{\prime}\delta_{kx}r_{kl}$ (51)
-$\times\left[\delta(\tau-\tau^{\prime})\delta_{ik}\delta_{jl}+\mathds{1}_{\tau=\tau^{\prime}}\delta_{ik}r_{ij}+\mathds{1}_{\tau<\tau^{\prime}}r_{ij}P(i,\tau^{\prime}|l,\tau)\right].$
+$\times\left[\delta(\tau-\tau^{\prime})\delta_{ik}\delta_{jl}+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau=\tau^{\prime}}\delta_{ik}r_{ij}+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}r_{ij}P(i,\tau^{\prime}|l,\tau)\right].$
 The middle term can be neglected when integrating over $\tau$ and $\tau^{\prime}$ as it has measure zero.
@@
 We may rewrite the path measure Eq. (41) with the stochastic differentials discussed above as
-$\mathds{P}[(x_{\tau})_{0\leq\tau\leq t}]=p_{x_{0}}(0)\exp\left[\int_{s=0}^{s=t}\sum_{x}r_{xx}\mathrm{d}\tau_{x}(s)+\int_{s=0}^{s=t}\sum_{x,y\neq x}\log r_{xy}\mathrm{d}n_{xy}(s)\right],$ (44)
+${\color[rgb]{0,0,0}\mathbb{P}}[(x_{\tau})_{0\leq\tau\leq t}]=p_{x_{0}}(0)\exp\left[\int_{s=0}^{s=t}\sum_{x}r_{xx}\mathrm{d}\tau_{x}(s)+\int_{s=0}^{s=t}\sum_{x,y\neq x}\log r_{xy}\mathrm{d}n_{xy}(s)\right],$ (52)
 where the second sum is taken over both $x$ and $y\neq x$. The advantage of writing the path measure as in Eq. (52) instead of Eq. (41) is that it allows for possible time dependencies of the rates.
-Meaning Continuous Discrete
-Displacement $\mathrm{d}\bm{x}_{\tau}$ $\mathrm{d}\bm{n}(\tau)$
-Deterministic Drift $\bm{F}(\bm{x}_{\tau})\mathrm{d}\tau$ $\bm{R}(x_{\tau})\mathrm{d}\tau$
-Noise $\bm{\sigma}\mathrm{d}\bm{W}_{\tau}$ $\mathrm{d}\bm{\varepsilon}(\tau)$
-Table 1: Overview of quantities entering the stochastic equations of motion in continuous and discrete space.
 III.2 Pathwise Observables: Currents and Densities
@@
 Similarly, a density assigns each state $i$ a weight $V_{i}(\tau)$ and then integrates these along the trajectory.
-These weights can be combined into the state function vector $\bm{V}(\tau)=(V_{1}(\tau),V_{2}(\tau),\dots)^{T}$.
-With $\bm{V}(\tau)$ we define the density [61]
-$\rho_{t}=\int_{\tau=0}^{\tau=t}\bm{V}(\tau)^{T}\mathrm{d}\bm{\tau}(\tau),$ (49)
-where $\mathrm{d}\bm{\tau}(\tau)=(\mathrm{d}\tau_{1}(\tau),\mathrm{d}\tau_{2}(\tau),\dots)^{T}$.
+These weights can be combined into the state function $V_{\tau}\equiv V(x_{\tau},\tau)=\sum_{k}\delta_{x_{\tau}k}V_{k}(\tau)$.
+With ${V}_{\tau}$ we define the density [38]
+[…]
 An exemplary trajectory with corresponding time-integrated density and current is shown in Fig. 3 to visualize Eqs. (55) and (57).
-Figure 3: Example trajectory and examples of an arising current and density.
-The part of the trajectory in (a) which contributes to the density $\rho_{\tau}$ with state function, $\bm{V}^{T}=(0,1,0)$, is marked in blue.
-Similarly, the transitions contributing to the current $J_{\tau}$ with transition weights, $\kappa_{ij}=\delta_{i2}\delta_{j3}-\delta_{i3}\delta_{j2}$, are marked by orange arrows with numbers representing the weight of the transition.
-The resulting density and current can be seen in (b) and (c), respectively.
+[…]
 III.3 From Stratonovich Increments to Observable Correlations
@@
 $\hat{\mathcal{I}}^{t,\mathrm{c}}_{M,N}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{\tau}^{t}\mathrm{d}\tau^{\prime}\int\mathrm{d}\bm{z}\int\mathrm{d}\bm{z}^{\prime}M(\bm{z})^{T}[\cdot]N(\bm{z}^{\prime}),$ (58a)
-$\hat{{I}}^{t}_{U^{\alpha},U^{\beta}}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{\tau}^{t}\mathrm{d}\tau^{\prime}\sum_{x,y}U_{xy}^{\alpha}\sum_{i,j}U^{\beta}_{ij}[\cdot],$ (50b)
+$\hat{{I}}^{t}_{{\color[rgb]{0,0,0}G}^{\alpha},{\color[rgb]{0,0,0}G}^{\beta}}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{\tau}^{t}\mathrm{d}\tau^{\prime}\sum_{x,y}{\color[rgb]{0,0,0}G}_{xy}^{\alpha}\sum_{i,j}{\color[rgb]{0,0,0}G}^{\beta}_{ij}[\cdot],$ (58b)
 where $\cdot$] in the first line denotes functions which depend on $\bm{z}$, $\bm{z}^{\prime}$, $\tau$, and $\tau^{\prime}$.
-Similarly, the functions entering $\hat{{I}}^{t}_{U^{\alpha},U^{\beta}}[\cdot]$ depend on $x$, $y$, $i$, $j$, $\tau$, and $\tau^{\prime}$.
-The functions $M(\bm{z})$ and $N(\bm{z})$ may be scalar- or vector-valued functions, while $U^{\alpha}$ and $U^{\beta}$ are two-point quantities, e.g., $\bm{\kappa}$.
+Similarly, the functions entering $\hat{{I}}^{t}_{{\color[rgb]{0,0,0}G}^{\alpha},{\color[rgb]{0,0,0}G}^{\beta}}[\cdot]$ depend on $x$, $y$, $i$, $j$, $\tau$, and $\tau^{\prime}$.
+The functions $M(\bm{z})$ and $N(\bm{z})$ may be scalar- or vector-valued functions, while ${\color[rgb]{0,0,0}G}^{\alpha}$ and ${\color[rgb]{0,0,0}G}^{\beta}$ are two-point quantities, e.g., $\bm{\kappa}$.
 Equation (58b) corresponds to Eq.
@@
 While one may expect the integrals to be substituted by a single sum, the double sum is indeed required to include non-local quantities, e.g., currents.
-We will now consider two observables $X_{t}^{\mathrm{(c)}}\in\{J_{t}^{\mathrm{(c),k}},\rho_{t}^{\mathrm{(c),k}}\}$ and $Y_{t}^{\mathrm{(c)}}\in\{J_{t}^{\mathrm{(c),l}},\rho_{t}^{\mathrm{(c),l}}\}$, where the superscript is used to distinguish discrete from the continuous setting and $k,l\in\{1,2\}$ is the index for currents $J_{t}^{\mathrm{(c),1}},J_{t}^{\mathrm{(c),2}}$ and densities $\rho_{t}^{\mathrm{(c),1}},\rho_{t}^{\mathrm{(c),2}}$.
-These currents and densities may have different current-/density-defining “window” functions, which we take into consideration here.
+We will now consider two observables $X_{t}^{\mathrm{(c)}}\in\{J_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}},\rho_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}}\}$ and $Y_{t}^{\mathrm{(c)}}\in\{J_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}},\rho_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}}\}$, where the superscript (c) is used to distinguish discrete (no superscript) from the continuous (superscript $\mathrm{c}$) setting and $k,l\in\{1,2\}$ is the index for currents $J_{t,{\color[rgb]{0,0,0}1}}^{\mathrm{(c)}},J_{t,{\color[rgb]{0,0,0}2}}^{\mathrm{(c)}}$ and densities $\rho_{t,{\color[rgb]{0,0,0}1}}^{\mathrm{(c)}},\rho_{t,{\color[rgb]{0,0,0}2}}^{\mathrm{(c)}}$.
+These currents and densities may have different current-/density-defining functions, which we take into consideration here.
 As shown in Refs.
@@
 $\mathrm{cov}(X_{t},Y_{t})$ $=\hat{{I}}^{t}_{U^{X},U^{Y}}[{\Xi}^{m}_{ijxy}]-\langle X_{t}\rangle\langle Y_{t}\rangle,$ (59)
-where $m=1$ for $X_{t}^{\mathrm{(c)}}=\rho_{t}^{\mathrm{(c)},k}$ and $Y_{t}^{\mathrm{(c)}}=\rho_{t}^{\mathrm{(c)},l}$, w.l.o.g.
-$m=2$ for $X_{t}^{\mathrm{(c)}}=J_{t}^{\mathrm{(c)},k}$ and $Y_{t}^{\mathrm{(c)}}=\rho_{t}^{\mathrm{(c)},l}$, and $m=3$ for $X_{t}^{\mathrm{(c)}}=J_{t}^{\mathrm{(c)},k}$ and $Y_{t}^{\mathrm{(c)}}=J_{t}^{\mathrm{(c)},l}$.
+where the $\Xi$ matrices are defined through increment correlations, see Eqs.
+(60), (62), and (67).
+The index $m$ we use to distinguish between different observable correlations: $m=1$ for $X_{t}^{\mathrm{(c)}}=\rho_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}}$ and $Y_{t}^{\mathrm{(c)}}=\rho_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}}$, w.l.o.g.
+$m=2$ for $X_{t}^{\mathrm{(c)}}=J_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}}$ and $Y_{t}^{\mathrm{(c)}}=\rho_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}}$, and $m=3$ for $X_{t}^{\mathrm{(c)}}=J_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}}$ and $Y_{t}^{\mathrm{(c)}}=J_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}}$.
 For density-density covariance ($m=1$), the increment correlations read [36]
-${\Xi}_{1}^{\bm{z}\bm{z}^{\prime}}$ $=\langle\mathds{1}\rangle_{\bm{x}_{\tau}=\bm{z}}^{\bm{x}_{\tau^{\prime}}=\bm{z}^{\prime}}+\langle\mathds{1}\rangle_{\bm{x}_{\tau}=\bm{z}^{\prime}}^{\bm{x}_{\tau^{\prime}}=\bm{z}},$ (52)
+${\Xi}_{1}^{\bm{z}\bm{z}^{\prime}}$ $=\langle{\color[rgb]{0,0,0}\mathbb{1}}\rangle_{\bm{x}_{\tau}=\bm{z}}^{\bm{x}_{\tau^{\prime}}=\bm{z}^{\prime}}+\langle{\color[rgb]{0,0,0}\mathbb{1}}\rangle_{\bm{x}_{\tau}=\bm{z}^{\prime}}^{\bm{x}_{\tau^{\prime}}=\bm{z}},$ (60)
 ${\Xi}^{1}_{ijxy}$ $=\frac{\langle\mathrm{d}\tau_{x}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle_{x_{\tau}=x}^{x_{\tau^{\prime}}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}+\frac{\langle\mathrm{d}\tau_{x}(\tau^{\prime})\mathrm{d}\tau_{i}(\tau)\rangle^{x_{\tau^{\prime}}=x}_{x_{\tau}=i}}{\mathrm{d}\tau\mathrm{d}\tau^{\prime}}.$
@@
 Evaluating these correlations yields the stationary covariances [36, 35]
-$\mathrm{cov}({\rho}_{t}^{\mathrm{c},k},{\rho}_{t}^{\mathrm{c},l})=$ $\hat{\mathcal{I}}^{t,\mathrm{c}}_{V^{k},V^{l}}\left[P_{\mathrm{s}}(\bm{z},\tau;\bm{z}^{\prime},\tau^{\prime})+P_{\mathrm{s}}(\bm{z}^{\prime},\tau;\bm{z},\tau^{\prime})-p_{\mathrm{s}}(\bm{z})p_{\mathrm{s}}(\bm{z}^{\prime})\right],$
-$\mathrm{cov}({\rho}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P_{\mathrm{s}}(x,\tau;i,\tau^{\prime})+P_{\mathrm{s}}(x,\tau^{\prime};i,\tau)-p_{x}^{\mathrm{s}}p_{i}^{\mathrm{s}}\right].$ (53)
+$\mathrm{cov}({\rho}_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{c}},{\rho}_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{c}})=$ $\hat{\mathcal{I}}^{t,\mathrm{c}}_{V^{k},V^{l}}\left[P_{\mathrm{s}}(\bm{z},\tau;\bm{z}^{\prime},\tau^{\prime})+P_{\mathrm{s}}(\bm{z}^{\prime},\tau;\bm{z},\tau^{\prime})-p_{\mathrm{s}}(\bm{z})p_{\mathrm{s}}(\bm{z}^{\prime})\right],$
+$\mathrm{cov}({\rho}_{t,{\color[rgb]{0,0,0}k}},{\rho}_{t,{\color[rgb]{0,0,0}l}})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P_{\mathrm{s}}(x,\tau;i,\tau^{\prime})+P_{\mathrm{s}}(x,\tau^{\prime};i,\tau)-p_{x}^{\mathrm{s}}p_{i}^{\mathrm{s}}\right].$ (61)
 The two-point function $V^{k}\delta$ stands for the (effectively one-point) quantity $V^{k}_{i}\delta_{ij}$  [^1] Note that the $\delta$ is introduced for convenience to allow for compact notation using the integration operator $\hat{{I}}^{t}_{U^{\alpha},U^{\beta}}[\cdot]$.
@@
 In both cases, the dual system (denoted by ) has the same steady state probability whereas the irreversible drift $\bm{j}_{\mathrm{s}}(\bm{x})$ has the opposite sign $\bm{j}^{\ddagger}_{\mathrm{s}}(\bm{x})=-\bm{j}_{\mathrm{s}}(\bm{x})$.
-It is useful to define current and dual current operators; in the continuous-space setting, these are [133, 96, 125] $\hat{\bm{j}}_{\bm{x}}=p_{s}(\bm{x})^{-1}\bm{j}_{\mathrm{s}}(\bm{x})-p_{\mathrm{s}}(\bm{x})\bm{D}\nabla_{\bm{x}}p_{\mathrm{s}}(\bm{x})^{-1}$ and $\hat{\bm{j}}_{\bm{x}}^{\ddagger}=p_{s}(\bm{x})^{-1}\bm{j}_{\mathrm{s}}(\bm{x})+p_{\mathrm{s}}(\bm{x})\bm{D}\nabla_{\bm{x}}p_{\mathrm{s}}(\bm{x})^{-1}$.
+It is useful to define current and dual current operators; in the continuous-space setting, these are [37, 36, 35] $\hat{\bm{j}}_{\bm{x}}=p_{s}(\bm{x})^{-1}\bm{j}_{\mathrm{s}}(\bm{x})-p_{\mathrm{s}}(\bm{x})\bm{D}\nabla_{\bm{x}}{\color[rgb]{0,0,0}\left[{\color[rgb]{0,0,0}p_{\mathrm{s}}(\bm{x})^{-1}}\right]}$ and $\hat{\bm{j}}_{\bm{x}}^{\ddagger}=p_{s}(\bm{x})^{-1}\bm{j}_{\mathrm{s}}(\bm{x})+p_{\mathrm{s}}(\bm{x})\bm{D}\nabla_{\bm{x}}{\color[rgb]{0,0,0}\left[{\color[rgb]{0,0,0}p_{\mathrm{s}}(\bm{x})^{-1}}\right]}$.
 The corresponding operators for discrete-state systems are much simpler and read $\hat{j}_{xy}=r_{xy}$ and $\hat{j}_{xy}^{\ddagger}=r_{yx}p_{y}^{\mathrm{s}}/p_{x}^{\mathrm{s}}$, respectively.
@@
 Moreover, we can rewrite the current-density covariances compactly and in a unified form as [36, 35]
-$\mathrm{cov}({J}_{t}^{\mathrm{c},k},{\rho}_{t}^{\mathrm{c},l})$ $=\hat{\mathcal{I}}^{t,\mathrm{c}}_{\bm{U}^{k},V^{l}}\left[\hat{\bm{j}}_{\bm{z}^{\prime}}P(\bm{z}^{\prime},t^{\prime}|\bm{z})p_{\mathrm{s}}(\bm{z})+\hat{\bm{j}}^{\ddagger}_{\bm{z}}P(\bm{z}^{\prime},t^{\prime}|\bm{z})p_{\mathrm{s}}(\bm{z})-\bm{j}_{\mathrm{s}}(\bm{z})p_{\mathrm{s}}(\bm{z}^{\prime})\right],$
-$\mathrm{cov}({J}_{t}^{k},{\rho}_{t}^{l})$ $=\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\hat{j}_{xy}P(x,t^{\prime}|i)p_{i}^{\mathrm{s}}+\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}p_{i}^{\mathrm{s}}\right].$ (58)
+$\mathrm{cov}({J}_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{c}},{\rho}_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{c}})$ $=\hat{\mathcal{I}}^{t,\mathrm{c}}_{\bm{U}^{k},V^{l}}\left[\hat{\bm{j}}_{\bm{z}^{\prime}}P(\bm{z}^{\prime},t^{\prime}|\bm{z})p_{\mathrm{s}}(\bm{z})+\hat{\bm{j}}^{\ddagger}_{\bm{z}}P(\bm{z}^{\prime},t^{\prime}|\bm{z})p_{\mathrm{s}}(\bm{z})-\bm{j}_{\mathrm{s}}(\bm{z})p_{\mathrm{s}}(\bm{z}^{\prime})\right],$
+$\mathrm{cov}({J}_{t,{\color[rgb]{0,0,0}k}},{\rho}_{t,{\color[rgb]{0,0,0}l}})$ $=\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\hat{j}_{xy}P(x,t^{\prime}|i)p_{i}^{\mathrm{s}}+\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}p_{i}^{\mathrm{s}}\right].$ (66)
 The only time dependence in the integrands in Eq.
@@
 Using again the generalized time-reversal symmetry, we can evaluate these correlations to obtain
-$\mathrm{cov}({J}_{t}^{\mathrm{c},k},J_{t}^{\mathrm{c},l})=$ $2t\int\mathrm{d}\bm{z}(\bm{U}^{k}(\bm{z}))^{T}\bm{D}\bm{U}^{l}(\bm{z})p_{\mathrm{s}}(\bm{z})$
+$\mathrm{cov}({J}_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{c}},J_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{c}})=$ $2t\int\mathrm{d}\bm{z}(\bm{U}^{k}(\bm{z}))^{T}\bm{D}\bm{U}^{l}(\bm{z})p_{\mathrm{s}}(\bm{z})$
 $+\hat{\mathcal{I}}^{t,\mathrm{c}}_{\bm{U}^{k},\bm{U}^{l}}\left[\hat{\bm{j}}_{\bm{z}}\left(\hat{\bm{j}}^{\ddagger}_{\bm{z}^{\prime}}\right)^{T}P(\bm{z},t^{\prime}|\bm{z}^{\prime})p_{\mathrm{s}}(\bm{z}^{\prime})+\hat{\bm{j}}^{\ddagger}_{\bm{z}}\hat{\bm{j}}_{\bm{z}^{\prime}}^{T}P(\bm{z}^{\prime},t^{\prime}|\bm{z})p_{\mathrm{s}}(\bm{z})-\bm{j}_{\mathrm{s}}(\bm{z})\bm{j}_{\mathrm{s}}(\bm{z}^{\prime})^{T}\right]$
-$\mathrm{cov}(J_{t}^{k},{J}_{t}^{l})=$ $t\sum_{x,y}\kappa_{xy}^{k}\kappa_{xy}^{l}r_{xy}p_{x}^{\mathrm{s}}+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\hat{j}_{xy}\hat{j}_{ji}^{\ddagger}P(x,t^{\prime}|j)p_{j}^{\mathrm{s}}+\hat{j}_{ij}\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}J^{\mathrm{s}}_{ij}\right],$ (60)
+$\mathrm{cov}(J_{t,{\color[rgb]{0,0,0}k}},{J}_{t,{\color[rgb]{0,0,0}l}})=$ $t\sum_{x,y}\kappa_{xy}^{k}\kappa_{xy}^{l}r_{xy}p_{x}^{\mathrm{s}}+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\hat{j}_{xy}\hat{j}_{ji}^{\ddagger}P(x,t^{\prime}|j)p_{j}^{\mathrm{s}}+\hat{j}_{ij}\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}J^{\mathrm{s}}_{ij}\right],$ (68)
 where the analogy between continuous and discrete state spaces is evident.
@@
 Here we highlight the similarities and differences between discrete space and continuous space.
-The mathematical structure of the results is (up to the analogy) the same, which should not come as a surprise, at least (but not only) in the sense that the MJP process can (under appropriate conditions on the drift $\bm{F}$ in Eq.
-(1), i.e., that the reversible contribution stems from a generalized potential with deep minima separated by high barriers [1, 2, 3, 11, 16]) be seen as the long-time coarse-grained limit of the continuous system.
-The continuous space correlations in Eqs.
-(52), (54), and (59) are valid for all $t\geq 0$ and similarly for systems with nominally discrete degrees of freedom (such as spin systems).
-Conversely, if the MJP is meant as an approximate, long-time effective description of a system with continuous degrees of freedom, Eqs.
-(52), (54), and (59) described covariances of pathwise observables at sufficiently long times.
-A non-zero $\mathrm{cov}(J_{t}^{\mathrm{c}},\rho_{t}^{\mathrm{c}})$ necessarily indicates that detailed balance is broken [96].
-The same observation can be made in the discrete case, as the equilibrium dual operator $\hat{j}_{xy}^{{\ddagger},\mathrm{eq}}=\hat{j}_{xy}^{\mathrm{eq}}$ together with $\bm{\kappa}=\bm{\kappa}(\tau)=-\bm{\kappa}^{T}$ yields $\mathrm{cov}_{\mathrm{eq}}(J_{t},\rho_{t})=0$.
-To provide preliminary intuition, density (co)variances (scaled by $t^{-1}$) are shown in Fig.
-4 for the secondary active transport (SAT) model, which will be discussed in more detail in Sec.
-IV.
-In addition, in Fig.
-5c, we show the mean current $J_{t}$ arising in the stationary SAT model with the magnitude of the typical fluctuations using the standard deviation, $\sqrt{\mathrm{var}(J_{t})}$, and how these can be bounded from below using the stationary thermodynamic uncertainty relation (TUR).
 Figure 4: Two densities, ${\rho}_{t}^{\alpha}$ and ${\rho}_{t}^{\beta}$, are used to visualize density (co)variances in the SAT model, see Sec.
@@
 The initial condition is $p_{i}=(\delta_{i1}+\delta_{i3}+\delta_{i5})/3$ and the values of the parameters can be found in Tab.
-3.
+4.
 In (b), $e^{\mathrm{out}}_{B}=40$ is used.
+The mathematical structure of the results is (up to the analogy) the same, which should not come as a surprise, at least (but not only) in the sense that the MJP process can (under appropriate conditions on the drift $\bm{F}$ in Eq.
+(1), i.e., that the reversible contribution stems from a generalized potential with deep minima separated by high barriers [113, 52, 54, 48, 67]) be seen as the long-time coarse-grained limit of the continuous system.
+The continuous space correlations in Eqs.
+(60), (62), and (67) are valid for all $t\geq 0$ and similarly for systems with nominally discrete degrees of freedom (such as spin systems).
+Conversely, if the MJP is meant as an approximate, long-time effective description of a system with continuous degrees of freedom, Eqs.
+(60), (62), and (67) described covariances of pathwise observables at sufficiently long times.
 Figure 5: Statistics of transitions and waiting times in the stationary SAT model.
@@
 We use $t=10$ in every panel.
+A non-zero $\mathrm{cov}(J_{t}^{\mathrm{c}},\rho_{t}^{\mathrm{c}})$ necessarily indicates that detailed balance is broken [36].
+The same observation can be made in the discrete case, as the equilibrium dual operator $\hat{j}_{xy}^{{\ddagger},\mathrm{eq}}=\hat{j}_{xy}^{\mathrm{eq}}$ together with $\bm{\kappa}=\bm{\kappa}(\tau)=-\bm{\kappa}^{T}$ yields $\mathrm{cov}_{\mathrm{eq}}(J_{t},\rho_{t})=0$.
+To provide preliminary intuition, density (co)variances (scaled by $t^{-1}$) are shown in Fig.
+4 for the secondary active transport (SAT) model, which will be discussed in more detail in Sec.
+IV.
+In addition, in Fig.
+5c, we show the mean current $J_{t}$ arising in the stationary SAT model with the magnitude of the typical fluctuations using the standard deviation, $\sqrt{\mathrm{var}(J_{t})}$, and how these can be bounded from below using the stationary thermodynamic uncertainty relation (TUR).
 III.4 Transient Covariances
@@
 (58b)
-$\hat{{I}}^{t}_{U^{\alpha},U^{\beta}}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{0}^{t}\mathrm{d}\tau^{\prime}\sum_{x,y}U_{xy}^{\alpha}(\tau)\sum_{i,j}U^{\beta}_{ij}(\tau^{\prime})[\cdot],$ (61)
+$\hat{{I}}^{t}_{{\color[rgb]{0,0,0}G}^{\alpha},{\color[rgb]{0,0,0}G}^{\beta}}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{0}^{t}\mathrm{d}\tau^{\prime}\sum_{x,y}{\color[rgb]{0,0,0}G}_{xy}^{\alpha}(\tau)\sum_{i,j}{\color[rgb]{0,0,0}G}^{\beta}_{ij}(\tau^{\prime})[\cdot],$ (69)
 where $U^{\alpha}$ and $U^{\beta}$ are explicitly time-dependent.
@@
 (61), (66), and (68) generalize to
-$\mathrm{cov}({\rho}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P(x,\tau;i,\tau^{\prime})-p_{x}(\tau)p_{i}(\tau^{\prime})\right],$
-$\mathrm{cov}({J}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\mathds{1}_{\tau>\tau^{\prime}}r_{xy}P(x,\tau|i,\tau^{\prime})p_{i}(\tau^{\prime})+\mathds{1}_{\tau<\tau^{\prime}}r_{yx}P(i,\tau^{\prime}|y,\tau)p_{y}(\tau)-r_{xy}p_{x}(\tau)p_{i}(\tau^{\prime})\right],$
-$\mathrm{cov}(J_{t}^{k},{J}_{t}^{l})=$ $\int_{0}^{t}\mathrm{d}\tau\sum_{x,y}\kappa_{xy}^{k}(\tau)\kappa_{xy}^{l}(\tau)r_{xy}p_{x}(\tau)$
-$+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\mathds{1}_{\tau>\tau^{\prime}}r_{xy}r_{ij}P(x,\tau|j,\tau^{\prime})p_{i}(\tau^{\prime})+\mathds{1}_{\tau<\tau^{\prime}}r_{ij}r_{xy}P(i,\tau^{\prime}|y,\tau)p_{y}(\tau)-r_{xy}p_{x}(\tau)r_{ij}p_{i}(\tau^{\prime})\right],$ (62)
+$\mathrm{cov}({\rho}_{t,{\color[rgb]{0,0,0}k}},{\rho}_{t,{\color[rgb]{0,0,0}l}})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P(x,\tau;i,\tau^{\prime})-p_{x}(\tau)p_{i}(\tau^{\prime})\right],$
+$\mathrm{cov}({J}_{t,{\color[rgb]{0,0,0}k}},{\rho}_{t,{\color[rgb]{0,0,0}l}})=$ $\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[{\color[rgb]{0,0,0}\mathbb{1}}_{\tau>\tau^{\prime}}r_{xy}P(x,\tau|i,\tau^{\prime})p_{i}(\tau^{\prime})+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}r_{yx}P(i,\tau^{\prime}|y,\tau)p_{y}(\tau)-r_{xy}p_{x}(\tau)p_{i}(\tau^{\prime})\right],$
+$\mathrm{cov}(J_{t,{\color[rgb]{0,0,0}k}},{J}_{t,{\color[rgb]{0,0,0}l}})=$ $\int_{0}^{t}\mathrm{d}\tau\sum_{x,y}\kappa_{xy}^{k}(\tau)\kappa_{xy}^{l}(\tau)r_{xy}p_{x}(\tau)$
+$+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[{\color[rgb]{0,0,0}\mathbb{1}}_{\tau>\tau^{\prime}}r_{xy}r_{ij}P(x,\tau|j,\tau^{\prime})p_{i}(\tau^{\prime})+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}r_{ij}r_{xy}P(i,\tau^{\prime}|y,\tau)p_{y}(\tau)-r_{xy}p_{x}(\tau)r_{ij}p_{i}(\tau^{\prime})\right],$ (70)
 For a generalization to time-dependent rates see Sec. VII, and for time-inhomogeneous diffusion see Ref. [37].
@@
 IV.1 Model systems
-While discussing various applications of the developed stochastic calculus for MJP we will throughout elude to biophysically relevant model systems shown in Fig.
+While discussing various applications of the developed stochastic calculus for MJP we will throughout allude to biophysically relevant model systems shown in Fig.
 6.
 Thereby we aim mostly at illustrating how to apply the theoretical results rather than to discuss the respective biological implications.
-Figure 6: Overview of models used to visualize the results: (a) Markov model for secondary active transport (SAT); the states $(1,3,5)$ correspond to the protein funnel being open toward the cell exterior while states $(2,4,6)$ describe the funnel being open toward the cell interior.
+The three models we introduce in the following sections are the secondary active transport across cell membranes (and an extension of this model), the calmodulin folding dynamics, and a four-state toy model in various settings.
+Figure 6: Overview of models used to visualize the results: (a) Markov model for secondary active transport (SAT); the states $(1,3,5)$ correspond to the protein funnel being open toward the cell exterior while states $(2,4,6)$ describe the funnel being open toward the cell interior (see Tab.
+2).
 The transition rate for the conversion of the opening from the interior to the exterior (and vice versa) is $\gamma$, which in the presence of ligands AA or $B$ is scaled by a dimensionless factor $x$.
-AA and $B$ leave the funnel with rate $l_{A/B}$ and enter it with rate ${\rm e}^{\mathrm{out}/\mathrm{in}}_{A/B}$ outside/inside the cell.
-Of particular interest in the (extended) SAT model is the impact of ${\rm e}^{\mathrm{out}}_{B}$, i.e., the concentration of the helper molecule.
+AA and $B$ leave the funnel with rate $l_{A/B}$ and enter it with rate ${\color[rgb]{0,0,0}e}^{\mathrm{out}/\mathrm{in}}_{A/B}$ outside/inside the cell.
+Of particular interest in the (extended) SAT model is the impact of ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$, i.e., the concentration of the helper molecule.
 (b) Schematic of how the SAT model in (a) may be extended to a “train” of SAT models to capture the number of AA molecules, $n$, transported through the cell membrane.
@@
 (71)], and upon constant driving, respectively.
-We are interested in the response to temperature perturbations and investigate the effect on a single-time observable $A_{t}=\delta_{x_{t}1}+\delta_{x_{t}2}$ and a current with transition weights $\kappa_{ij}=\delta_{i4}\delta_{j1}-\delta_{i1}\delta_{j4}+\delta_{i3}\delta_{j2}-\delta_{i2}\delta_{j3}$.
-(e) 4-state ring model with constant driving having rates 1 in the clockwise and $\epsilon$ in counterclockwise direction; the stationary distribution is independent of $\epsilon$.
+We are interested in the response to temperature perturbations and investigate the effect on a single-time observable ${\color[rgb]{0,0,0}A}_{t}=\delta_{x_{t}1}+\delta_{x_{t}2}$ and a current with transition weights $\kappa_{ij}=\delta_{i4}\delta_{j1}-\delta_{i1}\delta_{j4}+\delta_{i3}\delta_{j2}-\delta_{i2}\delta_{j3}$.
+(e) 4-state ring model with constant driving having rates 1 in the clockwise and $\epsilon$ in countercockwise direction; the stationary distribution is independent of $\epsilon$.
 The model is used for a comparison of the quality of thermodynamic bounds in situations where only the mesostates A and B, corresponding to states $(1,2)$ and $(3,4)$, respectively, are observed without the ability to distinguish the respective transition pathways between them.
 IV.1.1 Secondary Active Transport
+Table 2: Overview of states in SAT model.
+State Meaning
+1 Funnel outside with $A$ bound
+2 Funnel inside with $A$ bound
+3 Funnel outside with nothing bound
+4 Funnel inside with nothing bound
+5 Funnel outside with $B$ bound
+6 Funnel inside with $B$ bound
 We first consider a Markov model of secondary active transport (SAT) of molecule $A$ through a protein funnel facilitated by molecule $B$ [14, 16, 15].
 The model, shown in Fig.
-6a, consists of six states; three [i.e., $(1,3,5)$] corresponding to the funnel being open toward the exterior of the cell and other three [i.e., $(2,4,6)$] open toward the interior of cell.
+6a and listed in Tab.
+2, consists of six states; three [i.e., $(1,3,5)$] corresponding to the funnel being open toward the exterior of the cell and other three [i.e., $(2,4,6)$] open toward the interior of cell.
+Hence, the state space is $\mathcal{N}=\{1,2,3,4,5,6\}$.
 In the absence of $A$ or $B$ bound, the funnel changes conformation with rate $\gamma$.
@@
 AA and $B$ leave or enter the funnel inside and outside the cell with rates $l_{A/B}$ and ${\rm e}^{\mathrm{out/in}}_{A/B}$, respectively.
-We further consider a periodic extension of coupled SAT models to describe the transport of molecules $A$ through the membrane (Fig.
+The adopted parametrization is listed in Tab.
+4 in Appendix A.11.
+We further consider a periodic extension of coupled SAT models to describe the total transport of molecules $A$ through the membrane (Fig.
 6b).
-We leave the rates in the system unchanged, whereas the state space is extended to $\mathcal{N}\times\mathds{Z}$, with the additional degree of freedom accounting for the total number of $A$ molecules outside the cell.
+We leave the rates in the system unchanged, whereas the state space is extended to $\mathcal{N}\times{\color[rgb]{0,0,0}\mathbb{Z}}$, with the additional degree of freedom accounting for the total number of $A$ molecules outside the cell.
+The transport can be split into two categories: symporters and antiporters.
+These refer to to molecule $A$ being transported parallel or anti-parallel to molecule $B$, respectively.
+We are interested in the fraction of the total steady-state entropy production that we can estimate via the thermodynamic bounds for various values of transition rates ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$.
+These transition rates are assumed proportional to the concentration of $B$ outside the cell [14], allowing us to “control” the value of ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$, leaving it as free parameter we can tune.
 IV.1.2 Calmodulin Folding Dynamics
@@
 [12] summarized in Tab.
-2 while the topology of the folding network is shown in Fig.
+3 in Appendix A.11 while the topology of the folding network is shown in Fig.
 6c.
@@
 $r_{xy}=D\mathrm{e}^{(\lambda E_{x}-(1-\lambda)E_{y})/T}.$ (71)
-We use the free energies listed in Tab. 4.
+We use the free energies listed in Tab. 5 in Appendix A.11.
 In the constant driving setting, we set the clockwise rates to unity whereas the counterclockwise rates are controlled by the parameter $\epsilon$.
@@
 In the direct proofs in Refs. [38, 40], the stochastic integral
-$A_{t}^{\mathrm{c}}\equiv\int_{\tau=0}^{\tau=t}\frac{\bm{j}(\bm{x}_{\tau},\tau)}{P(\bm{x}_{\tau},\tau)}\left[2\bm{D}(\bm{x}_{\tau})\right]^{-1}\bm{\sigma}(\bm{x}_{\tau})\mathrm{d}\bm{W}_{\tau},$ (65)
+$A_{t}^{\mathrm{c}}\equiv\int_{\tau=0}^{\tau=t}\frac{\bm{j}(\bm{x}_{\tau},\tau){\color[rgb]{0,0,0}{}^{T}}}{P(\bm{x}_{\tau},\tau)}\left[2\bm{D}(\bm{x}_{\tau})\right]^{-1}\bm{\sigma}(\bm{x}_{\tau})\mathrm{d}\bm{W}_{\tau},$ (73)
 is a central object, chosen to immediately yield $\langle(A_{t}^{\mathrm{c}})^{2}\rangle=\Delta S_{\mathrm{tot}}^{\mathrm{c}}(t)/2$ which allows for crucial simplifications in the proofs.
@@
 As a key quantity in the various proofs of thermodynamic bounds we introduce the auxiliary integral
-$A_{t}$ $=\int_{\tau=0}^{\tau=t}\mathrm{Tr}[z_{\tau}^{T}\mathrm{d}\bm{\varepsilon}(\tau)],$ (67)
+$A_{t}$ $=\int_{\tau=0}^{\tau=t}\mathrm{Tr}[{\color[rgb]{0,0,0}\bm{Z}(\tau)}^{T}\mathrm{d}\bm{\varepsilon}(\tau)],$ (75)
 where
-$(z_{\tau})_{ij}=\frac{r_{ij}p_{i}(\tau)-r_{ji}p_{j}(\tau)}{r_{ij}p_{i}(\tau)+r_{ji}p_{j}(\tau)}.$ (68)
+$({\color[rgb]{0,0,0}\bm{Z}(\tau)})_{ij}=\frac{r_{ij}p_{i}(\tau)-r_{ji}p_{j}(\tau)}{r_{ij}p_{i}(\tau)+r_{ji}p_{j}(\tau)}.$ (76)
 Using Eq. (49) the mean is $\langle A_{t}\rangle=0$ and the variance can be identified as a “pseudo entropy production” [150, 148]
@@
 $\langle$ $A_{t}J_{t}^{\mathrm{II}}\rangle$
-$=$ $\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}\mathds{1}_{\tau<\tau^{\prime}}$ (75)
+$=$ $\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}$ (83)
 $\times\sum_{x,y}\left[P(i,\tau^{\prime}|y,\tau)-P(i,\tau^{\prime}|x,\tau)\right]p_{x}(\tau)r_{xy}Z_{xy}(\tau),$
@@
 $\langle A_{t}\rho_{t}\rangle$ $=(t\partial_{t}-1)\langle\rho_{t}\rangle-\langle\tilde{\rho}_{t}\rangle,$
-$\tilde{\rho}_{t}$ $\equiv\int_{\tau=0}^{\tau=t}\tau\partial_{\tau}\bm{V}(\tau)^{T}\mathrm{d}\bm{\tau}(\tau).$ (79)
-By shifting the current $\Delta J_{t}\to\Delta J_{t}-c(t)\Delta\rho_{t}$ with some $c:\mathds{R}\to\mathds{R}$ in Eq. (74) already yields the transient correlation TUR (CTUR)
+[…]
+By shifting the current $\Delta J_{t}\to\Delta J_{t}-c(t)\Delta\rho_{t}$ with some $c:{\color[rgb]{0,0,0}\mathbb{R}}\to{\color[rgb]{0,0,0}\mathbb{R}}$ in Eq.
+(82) already yields the transient correlation TUR (CTUR)
 $\frac{\left[t\partial_{t}\langle J_{t}\rangle-\langle\tilde{J}_{t}\rangle-c(t)\left((t\partial_{t}-1)\langle\rho_{t}\rangle-\langle\tilde{\rho}_{t}\rangle\right)\right]^{2}}{\mathrm{var}(J_{t}-c(t)\rho_{t})}\leq\frac{\Delta S_{\mathrm{tot}}(t)}{2}.$ (88)
@@
 The former requires $\Delta J_{t}-c(t)\Delta\rho_{t}\propto A_{t}$.
-Since $c(t)\neq 0$, we see that the saturation is achieved when $V_{i}(\tau)=\hat{\bm{e}}_{i}^{T}\bm{\kappa}(\tau)\bm{L}\hat{\bm{e}}_{i}/c(t)=\sum_{j}\kappa_{ij}(\tau)r_{ij}/c(t)$ and $\bm{\kappa}=c^{\prime}\bm{Z}$ for any $c^{\prime}\in\mathds{R}$.
+Since $c(t)\neq 0$, we see that the saturation is achieved when $V_{i}(\tau)=\hat{\bm{e}}_{i}^{T}\bm{\kappa}(\tau)\bm{L}\hat{\bm{e}}_{i}/c(t)=\sum_{j}\kappa_{ij}(\tau)r_{ij}/c(t)$ and $\bm{\kappa}=c^{\prime}\bm{Z}$ for any $c^{\prime}\in{\color[rgb]{0,0,0}\mathbb{R}}$.
 Here, $\hat{\bm{e}}_{i}$ is the $i$th unit vector.
@@
 Hence, it seems to be more practical to optimize the bound w.r.t. $c(t)$ for a given $J_{t}$ and $\rho_{t}$. The l.h.s. of Eq. (88), which we denote $h(t)$, has two extrema in $c(t)$
-$c^{\dagger}(t)$ $=\frac{a(t)}{b(t)}\,$,
-$c^{*}(t)$ $=\frac{a(t)\mathrm{cov}({\rho}_{t},{J}_{t})-b(t)\mathrm{var}({J}_{t})}{a(t)\mathrm{var}({\rho}_{t})-b(t)\mathrm{cov}({\rho}_{t},{J}_{t})}\,$, (83)
+$c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$ $=\frac{a(t)}{b(t)}\,$,
+$c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$ $=\frac{a(t)\mathrm{cov}({\rho}_{t},{J}_{t})-b(t)\mathrm{var}({J}_{t})}{a(t)\mathrm{var}({\rho}_{t})-b(t)\mathrm{cov}({\rho}_{t},{J}_{t})}\,$, (91)
 where $a(t)\equiv t\partial_{t}\langle{J}_{t}\rangle-\langle{\tilde{J}}_{t}\rangle$ and $b(t)\equiv(t\partial_{t}-1)\langle{\rho}_{t}\rangle-\langle{\tilde{\rho}}_{t}\rangle$.
@@
 (91) are accessible from $J_{t}$ and $\rho_{t}$, making the optimization practically feasible.
-Note that $c^{\dagger}(t)$ and $c^{*}(t)$ correspond to the minimum and maximum of $h(t)$, respectively, as
-$\frac{\partial^{2}}{\partial c^{2}}h(c,t)|_{c(t)=c^{\dagger}(t)}$ $=\frac{4b(t)^{2}}{\mathrm{var}\left({J}_{t}-\frac{a(t)}{b(t)}\rho_{t}\right)}\geq 0\,$, (84)
-$\frac{\partial^{2}}{\partial c^{2}}h(c,t)|_{c(t)=c^{*}(t)}$ $=\frac{-4b(t)^{2}\mathrm{var}\left({J}_{t}-\frac{a(t)}{b(t)}\rho_{t}\right)}{\left(\mathrm{var}\left({J}_{t}-c^{*}(t)\rho_{t}\right)\right)^{2}}\leq 0\,.$
-Moreover, one can immediately see that using $c^{\dagger}(t)$ one recovers the second law of thermodynamics, as $t\partial_{t}\langle J_{t}\rangle-\langle\tilde{J}_{t}\rangle-c^{\dagger}(t)\left[(t\partial_{t}-1)\langle\rho_{t}\rangle-\langle\tilde{\rho}_{t}\rangle\right]=0$.
+Note that $c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$ and $c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$ correspond to the minimum and maximum of $h(t)$, respectively, as
+$\frac{\partial^{2}}{\partial c^{2}}h(c,t)|_{c(t)=c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)}$ $=\frac{4b(t)^{2}}{\mathrm{var}\left({J}_{t}-\frac{a(t)}{b(t)}\rho_{t}\right)}\geq 0\,$, (92)
+$\frac{\partial^{2}}{\partial c^{2}}h(c,t)|_{c(t)=c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)}$ $=\frac{-4b(t)^{2}\mathrm{var}\left({J}_{t}-\frac{a(t)}{b(t)}\rho_{t}\right)}{\left(\mathrm{var}\left({J}_{t}-c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)\rho_{t}\right)\right)^{2}}\leq 0\,.$
+Moreover, one can immediately see that using $c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$ one recovers the second law of thermodynamics, as $t\partial_{t}\langle J_{t}\rangle-\langle\tilde{J}_{t}\rangle-c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)\left[(t\partial_{t}-1)\langle\rho_{t}\rangle-\langle\tilde{\rho}_{t}\rangle\right]=0$.
 The optimal quality factor therefore reads
@@
 Only the saturated quality factor remains finite for $t\to\infty$.
-As expected, the optimal $\mathcal{Q}^{\mathrm{CTUR}}_{\mathrm{opt}}$ gives the best inferred entropy production when $\bm{\kappa}(\tau)$ and $\bm{V}(\tau)$ do not saturate the TUR, whereas $\mathcal{Q}^{\mathrm{CTUR}}_{c=10}\leq\mathcal{Q}^{J\mathrm{TUR}}_{(23)}$ for $t\leq 3$.
+As expected, the optimal $\mathcal{Q}^{\mathrm{CTUR}}_{\mathrm{opt}}$ gives the best inferred entropy production when $\bm{\kappa}(\tau)$ and ${V}_{\tau}$ do not saturate the TUR, whereas $\mathcal{Q}^{\mathrm{CTUR}}_{c=10}\leq\mathcal{Q}^{J\mathrm{TUR}}_{(23)}$ for $t\leq 3$.
 Hence, simply using an arbitrary $c(t)$ in the CTUR may in fact lead to a worse estimate for the entropy production compared to the current TUR.
@@
 8.
-The expressions of $c^{*}(t)$ and $c^{\dagger}(t)$ Eqs.
+The expressions of $c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$ and $c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$ Eqs.
 (91) can be both positive and negative.
-Furthermore, they are in no particular relation to another, i.e., one cannot easily say if and when $c^{*}(t)<c^{\dagger}(t)$ and vice versa.
+Furthermore, they are in no particular relation to another, i.e., one cannot easily say if and when $c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)<c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$ and vice versa.
 A more detailed discussion is given in App.
@@
 6c).
-We present the various CTUR quality factors $\mathcal{Q}$ arising from various combinations of density, $\rho$, and current, $J_{t}$, as a function of time.
-The black line denotes the saturation of the TUR, i.e., when $\bm{\kappa}(\tau)=z_{\tau}$ and $c(t)\rho_{t}=J_{t}^{\mathrm{II}}$, corresponding to the estimated entropy production being the pseudo-entropy production $\Sigma_{\mathrm{est}}=2\langle A_{t}^{2}\rangle$.
+We present the various CTUR quality factors $\mathcal{Q}$ arising from various combinations of density, $\rho_{{\color[rgb]{0,0,0}\tau}}$, and current, $J_{t}$, as a function of time.
+The black line denotes the saturation of the TUR, i.e., when $\bm{\kappa}(\tau)={\color[rgb]{0,0,0}\bm{Z}(\tau)}$ and $c(t)\rho_{t}=J_{t}^{\mathrm{II}}$, corresponding to the estimated entropy production being the pseudo-entropy production $\Sigma_{\mathrm{est}}=2\langle A_{t}^{2}\rangle$.
 Above the black line is the region $\Sigma_{\mathrm{est}}>2\langle A_{t}^{2}\rangle$ which is not accessible using the (C)TUR.
 The remaining quality factors shown use the current ($\mathcal{Q}_{(23)}^{J\mathrm{TUR}}$), density ($\mathcal{Q}_{12}^{\rho\mathrm{TUR}}$), or a combination of these ($\mathcal{Q}_{c=10}^{\mathrm{CTUR}}$ and $\mathcal{Q}_{\mathrm{opt}}^{\mathrm{CTUR}}$) in orange, green, purple, and red, respectively.
-For the CTURs, the proportionality $c(t)=10$ (purple) and $c(t)=c^{*}(t)$ (red) are chosen, the latter optimizing the quality factor for given $J_{t}$ and $\rho_{t}$.
+For the CTURs, the proportionality $c(t)=10$ (purple) and $c(t)=c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$ (red) are chosen, the latter optimizing the quality factor for given $J_{t}$ and $\rho_{t}$.
 The rates can be found in Tab.
-2.
+3.
 Figure 8: Quality factor as a function of density weighting function $c(t)$ in the calmodulin system at a fixed $t=0.05$.
@@
 The state function entering the density is $V_{i}(\tau)=\delta_{i3}$ and the initial condition is $p_{i}(0)=\delta_{i6}$.
-The circles are the quality factors for most optimal $c(t)=c^{*}(t)$, while the squares denote the least optimal $c(t)=c^{\dagger}(t)$.
+The circles are the quality factors for most optimal $c(t)=c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$, while the squares denote the least optimal $c(t)=c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$.
 IV.3 Thermodynamic Transport Bound
@@
 $B_{t}=\int_{\tau=0}^{\tau=t}\mathrm{Tr}[\bm{H}(\tau)\mathrm{d}\bm{\varepsilon}(\tau)],$ (97)
-where $\bm{H}(\tau)]_{ij}=z_{i}(\tau)-z_{j}(\tau)$ is the difference of some observable state function $z_{\tau}\equiv z(x_{\tau},\tau)=\sum_{k}\delta_{k,x_{\tau}}z_{k}(\tau)$.
+where $\bm{H}(\tau)]_{ij}=z_{i}(\tau)-z_{j}(\tau)$ is the difference of some observable state function $z_{\tau}\equiv z(x_{\tau},\tau)=\sum_{k}{\color[rgb]{0,0,0}\delta_{x_{\tau}k}}z_{k}(\tau)$.
 The mean of $B_{t}$ vanishes while the second moment reads
@@
 Namely, consider the infinitesimal difference $\mathrm{d}z(\tau)=z_{\tau+\mathrm{d}\tau}-z_{\tau}$ along the trajectory.
-Then for any $n\in\mathds{N}_{>0}$ we then get $\langle[\mathrm{d}z(\tau)]^{n}\rangle=\sum_{x,y}(z_{y}-z_{x})^{n}r_{xy}p_{x}(\tau)\mathrm{d}\tau+\mathcal{O}(\mathrm{d}\tau^{2})$.
+Then for any $n\in{\color[rgb]{0,0,0}\mathbb{N}}_{>0}$ we then get $\langle[\mathrm{d}z(\tau)]^{n}\rangle=\sum_{x,y}(z_{y}-z_{x})^{n}r_{xy}p_{x}(\tau)\mathrm{d}\tau+\mathcal{O}(\mathrm{d}\tau^{2})$.
 Thus, $\mathrm{var}[\mathrm{d}z(\tau)]=\sum_{x,y}(z_{y}-z_{x})^{2}r_{xy}p_{x}(\tau)\mathrm{d}\tau+\mathcal{O}(\mathrm{d}\tau^{2})$ and
@@
 $=\langle z_{t}-z_{0}-\int_{0}^{t}\mathrm{d}\tau\partial_{\tau}z_{\tau}\rangle.$ (100)
-The average of $z_{t}$ (and of any vector) is given by $\langle z_{\tau}\rangle=\sum_{k}\sum_{l}\delta_{k,l}z_{k}(\tau)p_{l}(t)$.
+The average of $z_{t}$ is given by $\langle z_{\tau}\rangle=\sum_{k}\sum_{l}\delta_{kl}z_{k}(\tau)p_{l}(t)$.
 By combining the various averages entering the Cauchy-Schwarz inequality $\langle A_{t}B_{t}\rangle\leq\langle A_{t}^{2}\rangle\langle B_{t}^{2}\rangle$ we obtain the transport bound for MJP
@@
 Saturation of Eq.
-(93) entails choosing $z_{\tau}$ such that $\bm{H}(\tau)^{T}\propto z_{\tau}$ to saturate the Cauchy-Schwarz inequality.
+(101) entails choosing $z_{\tau}$ such that $\bm{H}(\tau)^{T}\propto{\color[rgb]{0,0,0}\bm{Z}(\tau)}$ to saturate the Cauchy-Schwarz inequality.
 Explicitly, for all $\tau$
@@
 A similar connection can be made in continuous space, where a recent TUR for underdamped Langevin dynamics [104] can be seen as a generalization of the continuous TB [40].
-It should be noted that, while the TB in general cannot be saturated, the gTB can be saturated with the choice of weight $\bm{\kappa}(\tau)\propto z_{\tau}$.
+It should be noted that, while the TB in general cannot be saturated, the gTB can be saturated with the choice of weight $\bm{\kappa}(\tau)\propto{\color[rgb]{0,0,0}\bm{Z}(\tau)}$.
 Moreover, in some cases the transient TB of one system may be identical to the stationary gTB of another system.
@@
 The parameters are listed in Tab.
-3.
-The SAT describes the transport of a large molecule (denoted molecule $A$) through a cell membrane against its chemical gradient using the transport of helper molecule (denoted molecule $B$) [145, 146, 147], see Fig.
-6a-b.
-The reduction to a six-state Markov model, $\mathcal{N}=\{1,2,3,4,5,6\}$, is commonly used to describe this secondary active transport [145, 146, 147].
-The efficiency of the process is related to the number of molecules $A$ transported given a concentration of $B$.
-The transport can be split into two categories: symporters and antiporters.
-These refer to molecule $A$ being transported parallel or anti-parallel to molecule $B$, respectively.
-To effectively describe the total number of transported molecules, we extend the six-state model to a chain of six-state models, i.e., $\tilde{\mathcal{N}}=\mathcal{N}\times\mathds{Z}$, where the additional degree of freedom corresponds to the number of $A$ molecules transported through the membrane.
-We are interested in the fraction of the total steady-state entropy production that we can estimate via the thermodynamic bounds for various values of transition rates ${\rm e}^{\mathrm{out}}_{B}$.
-These transition rates are assumed proportional to the concentration of $B$ outside the cell [145].
-To make reasonable comparisons, we either have to consider symporter or antiporter systems.
+4.
+To make reasonable comparisons in the SAT model, we either have to consider symporter or antiporter systems.
 We are going to focus on the latter.
-We first vary ${\rm e}^{\mathrm{out}}_{B}$ to assess which values of transition rates give rise to antiporter behaviour before adressing thermodynamic inferrence.
+We first vary ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$ to assess which values of transition rates give rise to antiporter behaviour before adressing thermodynamic inferrence.
 We start with the cycle affinities of the model.
@@
 If $\langle J_{t}^{\mathrm{SAT}}\rangle_{\mathrm{s}}>0$, the average total number of molecules $A$ transported towards the outside of the cell is positive, corresponding to an average transport of $A$ from the inside to the outside of the cell.
-Thus, the condition for an antiporter behavior is $l_{A}/e^{\mathrm{out}}_{A}>p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}$, which depends on ${\rm e}^{\mathrm{out}}_{B}$ via the steady-state distribution.
+Thus, the condition for an antiporter behavior is $l_{A}/e^{\mathrm{out}}_{A}>p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}$, which depends on ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$ via the steady-state distribution.
 With the parameters we use, see Tab.
-3, we need $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}<2/3$.
+4, we need $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}<2/3$.
 The value $2/3$ is marked as a black line in the inset of Fig.
-9a, where the green line corresponds to the value of $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}$ as a function of ${\rm e}^{\mathrm{out}}_{B}$.
-The dashed red line marks the value of ${\rm e}^{\mathrm{out}}_{B}$ s.t.
+9a, where the green line corresponds to the value of $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}$ as a function of ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$.
+The dashed red line marks the value of ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$ s.t.
 $\mathcal{A}_{\mathcal{C}_{3}}=0$, while the dashed orange line is where $p_{3}^{\mathrm{s}}/p_{1}^{\mathrm{s}}=2/3$.
 The same can be readily seen in Fig.
-9b, where the mean current is presented for various values of ${\rm e}^{\mathrm{out}}_{B}$.
+9b, where the mean current is presented for various values of ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$.
 Clearly, the condition $\mathcal{A}_{\mathcal{C}_{3}}=0$ is not sufficient for a positive current.
-Additionally, we can consider the transient extended model and apply the TB.
-For the extended model, consider $p_{x,n}(\tau)$ the probability of being in state $x$ in system $n$ at time $\tau$, where $n\in\mathds{Z}$ corresponds to the number of $A$ molecules transported through the membrane.
-When marginalizing over $\mathds{Z}$ we recover the steady-state probability $p_{x}^{\mathrm{s}}=\sum_{n\in\mathds{Z}}p_{x,n}(\tau)$.
-As a state function we choose the number of $A$ molecules $z_{\tau}=n$, such that
-$\langle z_{\tau}\rangle=\sum_{n\in\mathds{Z}}n\sum_{x=1}^{6}p_{x,n}(\tau)=\sum_{n\in\mathds{Z}}n\tilde{p}_{n}(\tau),$ (103)
+[…]
+$\langle z_{\tau}\rangle=\sum_{n\in{\color[rgb]{0,0,0}\mathbb{Z}}}n\sum_{x=1}^{6}p_{x,n}(\tau)=\sum_{n\in{\color[rgb]{0,0,0}\mathbb{Z}}}n\tilde{p}_{n}(\tau),$ (111)
 where $\tilde{p}_{n}(\tau)$ is the probability distribution marginalized over $\mathcal{N}$.
@@
 The fluctuation-scale function is recognized to be
-$\mathcal{D}_{t}^{\mathrm{ext}}=\frac{1}{t}\sum_{n\in\mathds{Z}}\sum_{x,y\in\mathcal{N}}\int_{0}^{t}\mathrm{d}\tau$ $\left[(z_{y,n}-z_{x,n})^{2}\delta_{x\to y,n}+(z_{y,n+1}-z_{x,n})^{2}\delta_{x\to y,n\to n+1}\right$.
+$\mathcal{D}_{t}^{\mathrm{ext}}=\frac{1}{t}\sum_{n\in{\color[rgb]{0,0,0}\mathbb{Z}}}\sum_{x,y\in\mathcal{N}}\int_{0}^{t}\mathrm{d}\tau$ $\left[(z_{y,n}-z_{x,n})^{2}\delta_{x\to y,n}+(z_{y,n+1}-z_{x,n})^{2}\delta_{x\to y,n\to n+1}\right$.
 $+\left.(z_{y,n-1}-z_{x,n})^{2}\delta_{x\to y,n\to n-1}\right]r_{xy}p_{x,n}(\tau)\,.$ (112)
 Note that $\delta_{x\to y,n\to n+1}$ is one if the transition $x\to y$ is observed while the system goes from $n$ to $n+1$ and zero otherwise. We can simplify Eq. (112) to find
-$\mathcal{D}_{t}^{\mathrm{ext}}=\frac{1}{t}\sum_{n\in\mathds{Z}}\int_{0}^{t}\mathrm{d}\tau\left[r_{13}p_{1,n}(\tau)+r_{31}p_{3,n}(\tau)\right],$ (105)
+$\mathcal{D}_{t}^{\mathrm{ext}}=\frac{1}{t}\sum_{n\in{\color[rgb]{0,0,0}\mathbb{Z}}}\int_{0}^{t}\mathrm{d}\tau\left[r_{13}p_{1,n}(\tau)+r_{31}p_{3,n}(\tau)\right],$ (113)
 because $z_{x,n}-z_{y,n}=0$ and $z_{x,n}-z_{y,n\pm 1}=\pm 1$. Marginalizing over $n$ finally gives
@@
 The quality factors of the TUR $\mathcal{Q}^{\mathrm{TUR}}$ and TB $\mathcal{Q}^{\mathrm{TB}}$ are shown in Fig.
-9c (solid and dashed line, respectively) for ${\rm e}^{\mathrm{out}}_{B}$ in the antiporter regime.
+9c (solid and dashed line, respectively) for ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$ in the antiporter regime.
 Because $t\mathcal{D}_{t}^{\mathrm{ext}}>\mathrm{var}_{\mathrm{s}}(J_{t}^{\mathrm{SAT}})$ for all $t>0$, the TUR infers a higher fraction of the entropy production, i.e., $\mathcal{Q}^{\mathrm{TUR}}\geq\mathcal{Q}^{\mathrm{TB}}$.
-Additionally, one may show that taking ${\rm e}^{\mathrm{out}}_{B}\to\infty$ leads to the quality factors converging to a finite value [12, 152].
+Additionally, one may show that taking ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}\to\infty$ leads to the quality factors converging to a finite value [45, 157].
 Hence, there is no improvement in the inference of dissipation at higher concentrations of $B$ outside the cell beyond a certain point.
@@
 ${B}_{t}$ $=\frac{1}{\sqrt{t}}\int_{\tau=0}^{\tau=t}\mathrm{Tr}\left[\left(\bm{H}(\tau)-\tilde{\bm{Z}}(\tau)^{T}\right)\mathrm{d}\bm{\varepsilon}(\tau)\right],$ (115)
-where $\tilde{Z}_{xy}(\tau)=Z_{xy}(\tau)\left[z_{x}(\tau)+z_{y}(\tau)-2F(\tau)\right]$ for some time-dependent function $F(\tau)$.
+where $\tilde{Z}_{xy}(\tau)=Z_{xy}(\tau)\left[z_{x}(\tau)+z_{y}(\tau)-2F(\tau)\right]$ for some time-dependent function $F(\tau)$, $Z_{xy}$ is given by Eq.
+(76), and $\bm{H}$ is the same matrix also entering Eq.
+(97).
 The second key ingredient is the shifted and rescaled density observable
-${C}_{t}$ $=\frac{1}{\sqrt{t}}\int_{\tau=0}^{\tau=t}\left[\bm{V}-\langle{V}_{\tau}\rangle\bm{1}\right]^{T}\mathrm{d}\bm{\tau}(\tau),$ (108)
-where $\bm{1}$ is the one-vector [i.e.
-$(\bm{1})_{i}=1$ for all $i\in\mathcal{N}$] and $\langle V_{\tau}\rangle=\sum_{i}V_{i}p_{i}(\tau)$.
-While the TUR and TB allow the state functions $z_{\tau}$ and $\bm{V}(\tau)$ to be explicitly time-dependent, the correlation bound requires these to be strictly time-independent.
+${C}_{t}$ $=\frac{1}{\sqrt{t}}\int_{\tau=0}^{\tau=t}{\color[rgb]{0,0,0}\left[V_{\tau}-\langle{V}_{\tau}\rangle\right]\mathrm{d}{\tau}},$ (116)
+where $\langle V_{\tau}\rangle=\sum_{i}V_{i}p_{i}(\tau)$.
+While the TUR and TB allow the state functions $z_{\tau}$ and ${V}_{\tau}$ to be explicitly time-dependent, the correlation bound requires these to be strictly time-independent.
 Furthermore, we write, e.g., $\langle z_{\tau}\rangle=\sum_{i}z_{i}p_{i}(\tau)$ to explicitly highlight that the average depends on time through $p_{i}(\tau)$.
@@
 $t\mathrm{var}(\overline{V}_{t})=\langle C_{t}^{2}\rangle.$ (120)
-With the shorthand notations $\bm{V}^{\Delta}(\tau)=\bm{V}-\langle{V}(\tau)\rangle\bm{1}$ and $z^{F}_{\tau}=z_{\tau}-\langle F(\tau)\rangle\bm{1}$, the cross term is computed to be
-$t\langle B_{t}C_{t}\rangle$ $=\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{x,y,i}\mathds{1}_{\tau<\tau^{\prime}}V^{\Delta}_{i}(\tau^{\prime})\left(H_{yx}-Z_{xy}(\tau)\left[z_{x}^{F}(\tau)+z_{y}^{F}(\tau)\right]r_{xy}p_{x}(\tau)\right)\left[P(i,\tau^{\prime}|y,\tau)-P(i,\tau^{\prime}|x,\tau)\right],$ (113)
+With the shorthand notations $V_{x}^{\Delta}(\tau)=V_{x}-\langle V_{\tau}\rangle$ and $z^{F}_{\tau}=z_{\tau}-\langle F(\tau)\rangle$, the cross term is computed to be
+$t\langle B_{t}C_{t}\rangle$ $=\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{x,y,i}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}V^{\Delta}_{i}(\tau^{\prime})\left(H_{yx}-Z_{xy}(\tau)\left[z_{x}^{F}(\tau)+z_{y}^{F}(\tau)\right]r_{xy}p_{x}(\tau)\right)\left[P(i,\tau^{\prime}|y,\tau)-P(i,\tau^{\prime}|x,\tau)\right],$ (121)
 where the sum over $x,y$ can be simplified because of symmetries (we suppress the time indices here) to get
-$\sum_{x,y}H_{yx}r_{xy}p_{x}[P(i|y)-P(i|x)]$
-$=-\sum_{x,y}z_{x}^{F}(r_{xy}p_{x}+r_{yx}p_{y})[P(i|y)-P(i|x)],$ (114)
+$\sum_{x,y}H_{yx}r_{xy}p_{x}[P(i|y)-P(i|x)]=-\sum_{x,y}z_{x}^{F}(r_{xy}p_{x}+r_{yx}p_{y})[P(i|y)-P(i|x)],$ (122)
 and
-$\sum_{x,y}Z_{xy}(z_{x}^{F}+z_{y}^{F})r_{xy}p_{x}\left[P(i|y)-P(i|x)\right]$
-$=\sum_{x,y}z_{x}^{F}(r_{xy}p_{x}-r_{yx}p_{y})\left[P(i|y)-P(i|x)\right].$ (115)
+$\sum_{x,y}Z_{xy}(z_{x}^{F}+z_{y}^{F})r_{xy}p_{x}\left[P(i|y)-P(i|x)\right]=\sum_{x,y}z_{x}^{F}(r_{xy}p_{x}-r_{yx}p_{y})\left[P(i|y)-P(i|x)\right].$ (123)
 Using time-homogeneity of the semi-group $\bm{P}(\tau^{\prime}|\tau)=\bm{P}(\tau^{\prime}-\tau)\equiv{\rm e}^{\bm{L}(\tau^{\prime}-\tau)}$ which obeys the master equation Eq.
@@
 $\partial_{\tau}\bm{P}(\tau)=\bm{L}\bm{P}(\tau)=\bm{P}(\tau)\bm{L}$, and $\partial_{\tau}\bm{P}(\tau^{\prime}-\tau)=-\partial_{\tau^{\prime}}\bm{P}(\tau^{\prime}-\tau)$, the integrand in Eq.
-(113) simplifies to $2\sum_{x,i}\mathds{1}_{\tau<\tau^{\prime}}V^{\Delta}_{i}(\tau^{\prime})z^{F}_{x}(\tau)p_{x}(\tau)\partial_{\tau^{\prime}}P(i,\tau^{\prime}|x,\tau)$.
+(121) simplifies to $2\sum_{x,i}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}V^{\Delta}_{i}(\tau^{\prime})z^{F}_{x}(\tau)p_{x}(\tau)\partial_{\tau^{\prime}}P(i,\tau^{\prime}|x,\tau)$.
 Since
@@
 the particular choice $F(\tau)=\langle z_{\tau}\rangle$ yields
-$t\langle B_{t}C_{t}\rangle$ $=2\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\mathds{1}_{\tau<\tau^{\prime}}\partial_{\tau^{\prime}}\mathrm{cov}({V}_{\tau^{\prime}},z_{\tau}).$ (117)
-An integration by parts in $\tau^{\prime}$ with $\partial_{\tau^{\prime}}\mathds{1}_{\tau<\tau^{\prime}}=\delta(\tau^{\prime}-\tau)$ further yields
+$t\langle B_{t}C_{t}\rangle$ $=2\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\partial_{\tau^{\prime}}\mathrm{cov}({V}_{\tau^{\prime}},z_{\tau}).$ (125)
+An integration by parts in $\tau^{\prime}$ with $\partial_{\tau^{\prime}}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}=\delta(\tau^{\prime}-\tau)$ further yields
 $\langle B_{t}C_{t}\rangle$ $=2\mathrm{cov}\left[{V}_{t},\overline{{z}}_{t}\right]-\frac{2}{t}\int_{0}^{t}\mathrm{d}\tau\mathrm{cov}\left[{V}_{\tau},z_{\tau}\right].$ (126)
@@
 In NESS, Eq. (130) simplifies to [we suppress time indices for stationary single-time (pseudo/co-)variances]
-$\frac{4\left[\mathrm{cov}({V}_{t},\overline{{z}}_{t})-\mathrm{cov}_{\mathrm{s}}({V},{z})\right]^{2}}{t\mathrm{var}(\overline{V}_{t})}-\mathcal{D}_{t}$
-$\leq\frac{\langle A_{t}^{2}\rangle_{\mathrm{s}}}{t}\mathrm{pvar}_{\mathrm{ps}}^{F}({z})\leq\frac{\Delta S_{\mathrm{tot}}(t)}{2t}\mathrm{pvar}_{\mathrm{ps}}^{F}({z}).$ (123)
+$\frac{4\left[\mathrm{cov}({V}_{t},\overline{{z}}_{t})-\mathrm{cov}_{\mathrm{s}}({V},{z})\right]^{2}}{t\mathrm{var}(\overline{V}_{t})}-\mathcal{D}_{t}\leq\frac{\langle A_{t}^{2}\rangle_{\mathrm{s}}}{t}\mathrm{pvar}_{\mathrm{ps}}^{F}({z})\leq\frac{\Delta S_{\mathrm{tot}}(t)}{2t}\mathrm{pvar}_{\mathrm{ps}}^{F}({z}).$ (131)
 The two point probability density $p_{xy}(\tau)$ is not accessible, since it requires knowing $\bm{L}$, hence the pseudo variance in Eqs.
@@
 The steady-state distribution $p_{i}^{\mathrm{s}}=p=1/4$ is independent of $\epsilon$.
-Suppose that we can only distinguish between a pair of mesostates A and B it can only be distinguished whether the states are in either states 1 and 2 or states 3 and 4.
+Suppose that we can only distinguish between a pair of mesostates A and B, i.e., it can only be distinguished whether the system is in either states 1 and 2 (mesostate A) or states 3 and 4 (mesostate B).
 These we refer to mesostates A (containing $\{1,2\}$) and B (containing $\{3,4\}$), see Fig.
@@
 (141)]
-$\langle O(t)\rangle_{(\delta)}=\sum_{(x_{\tau})_{0\leq\tau\leq t}}O[(x_{\tau})_{0\leq\tau\leq t}]\mathds{P}^{(\delta)}[(x_{\tau})_{0\leq\tau\leq t}].$ (134)
+$\langle O(t)\rangle_{(\delta)}=\sum_{(x_{\tau})_{0\leq\tau\leq t}}O[(x_{\tau})_{0\leq\tau\leq t}]{\color[rgb]{0,0,0}\mathbb{P}}^{(\delta)}[(x_{\tau})_{0\leq\tau\leq t}].$ (142)
 The path measures Eq. (41) of processes in Eqs. (47) and (141) can be conveniently written in terms of the stochastic differentials as
-$\mathds{P}^{\delta}[(x_{\tau})_{0\leq\tau\leq t}]=p_{x_{0}}(0)\exp\left[\int_{s=0}^{s=t}\sum_{x}r^{\delta}_{xx}\mathrm{d}\tau_{x}(s)+\int_{s=0}^{s=t}\sum_{x,y\neq x}\log r_{xy}^{\delta}\mathrm{d}n_{xy}(s)\right],$ (135)
-where $p_{x_{0}}(0)$ is the distribution of the initial state $x_{\tau=0}=x_{0}$.
-Note that a given path realization defines a unique sequence of transitions and hence $\mathrm{d}\tau_{x}(s)$ and $\mathrm{d}n_{xy}(s)$, which in turn enter equally $\mathds{P}$ in Eq.
-(33) and $\mathds{P}^{\delta}$ in Eq.
-(135).
-However, because of the different rates ($r^{\delta}_{xy}$ versus $r_{xy}$), the measure of the same path is generally different under $\mathds{P}$ and $\mathds{P}^{\delta}$.
-Moreover, writing the path measure in terms of stochastic differentials as in Eq.
-(135) further has the advantage of allowing for time-dependent rates.
+${\color[rgb]{0,0,0}\mathbb{P}}^{\delta}[(x_{\tau})_{0\leq\tau\leq t}]=p_{x_{0}}(0)\exp\left[\int_{s=0}^{s=t}\sum_{x}r^{\delta}_{xx}\mathrm{d}\tau_{x}(s)+\int_{s=0}^{s=t}\sum_{x,y\neq x}\log r_{xy}^{\delta}\mathrm{d}n_{xy}(s)\right],$ (143)
+[…]
 Consider the perturbed rates
 $r_{xy}^{\delta}=r_{xy}(\chi+\delta)=r_{xy}(\chi)+\delta\tilde{r}_{xy}(\chi)+\mathcal{O}(\delta^{2}).$ (144)
-Note that we assume that the rates $r_{xy}=r_{xy}(\chi)$ and $r_{xy}^{\delta}$ are functions of $\chi$ and $\chi+\delta$, respectively, ($\chi$ is a placeholder for the perturbed quantity, e.g., temperature $T$) and that the function is analytic at $\delta=0$, i.e., that it has a well-defined Taylor expansion at that point.
-Therefore, $\mathds{P}^{\delta}$ is absolutely continuous w.r.t.
-$\mathds{P}$ for sufficiently small $\delta$  [^3] For MJP, the perturbed measure $\mathds{P}^{\delta}\ll\mathds{P}$ is absolutely continuous w.r.t.
-the unperturbed measure $\mathds{P}$ as long as for all $x\neq y$ $r_{xy}>0\implies r_{xy}^{\delta}>0$, i.e., as long as the underlying topology of the MJP remains unchanged.
-and the perturbed average $\langle O(t)\rangle_{\delta}$ can be written in terms of the unperturbed process in Eq.
-(39) using the Cameron-Martin-Girsanov theorem with the Radon-Nikodym derivative [163]
-$\langle O(t)\rangle_{\delta}=\left\langle O(t)\frac{\mathrm{d}\mathds{P}^{\delta}}{\mathrm{d}\mathds{P}}\right\rangle$ (137)
+[…]
+$\langle O(t)\rangle_{\delta}=\left\langle O(t)\frac{\mathrm{d}{\color[rgb]{0,0,0}\mathbb{P}}^{\delta}}{\mathrm{d}{\color[rgb]{0,0,0}\mathbb{P}}}\right\rangle$ (145)
 where we assumed that the processes start from the same initial distribution. The Radon-Nikodym derivative reads
-$\frac{\mathrm{d}\mathds{P}^{\delta}}{\mathrm{d}\mathds{P}}\left[(x_{\tau})_{0\leq\tau\leq t}\right]$ $=\exp\left[\int_{s=0}^{s=t}\sum_{x}(r_{xx}^{\delta}-r_{xx})\mathrm{d}\tau_{x}(s)+\int_{s=0}^{s=t}\sum_{x,y\neq x}\log\frac{r_{xy}^{\delta}}{r_{xy}}\mathrm{d}n_{xy}(s)\right]$
+$\frac{\mathrm{d}{\color[rgb]{0,0,0}\mathbb{P}}^{\delta}}{\mathrm{d}{\color[rgb]{0,0,0}\mathbb{P}}}\left[(x_{\tau})_{0\leq\tau\leq t}\right]$ $=\exp\left[\int_{s=0}^{s=t}\sum_{x}(r_{xx}^{\delta}-r_{xx})\mathrm{d}\tau_{x}(s)+\int_{s=0}^{s=t}\sum_{x,y\neq x}\log\frac{r_{xy}^{\delta}}{r_{xy}}\mathrm{d}n_{xy}(s)\right]$
 $=\exp\left[-\delta\int_{s=0}^{s=t}\sum_{x,y\neq x}\tilde{r}_{xy}\mathrm{d}\tau_{x}(s)+\delta\int_{s=0}^{s=t}\sum_{x,y\neq x}\frac{\tilde{r}_{xy}}{r_{xy}}\mathrm{d}n_{xy}(s)+\mathcal{O}(\delta^{2})\right]$
@@
 for any $\lambda\in[0,1]$ (commonly used are $\lambda=1$ and $\lambda=1/2$) and $D$ may scale with $T$ ($D=\tilde{D}T$ with $\tilde{D}=\mathrm{const.}$). The rates (148) satisfy detailed balance.
-The aim here is to relate the response in Eq. (139) of a single-time observable $A(t)$ to the equilibrium correlation function,
+The aim here is to relate the response in Eq. (147) of a single-time observable $O(t)$ to the equilibrium correlation function,
 $C_{OE}(t)=\left(\langle OE\rangle_{\mathrm{eq}}-\langle O(t)E(0)\rangle_{\mathrm{eq}}\right).$ (149)
@@
 Equation (154) is the linear response of an equilibrium MJP to a temperature perturbation.
-V.3 Examples
+V.3 Connection to Response Function Formalism
+In this section, we connect the results derived in the previous sections to the well established response function formalism, which is often used to describe the fluctuation dissipation theorem.
+While the results of response functions for MJP has been previously derived in, e.g., [64, 110, 153, 173, 142] , we show how these may be extracted from our results to bridge the “gap” to the literature.
+V.3.1 Equilibrium Response
+For a first order perturbation theory of a single-time observable $B(t)$ in an equilibrium system with perturbation $h(s)$, $0\leq s\leq t$, conjugate to some observable $R(s)$, the response reads [64, 110]
+$\Delta B(t)\equiv\langle B(t)\rangle_{h}-\langle B\rangle_{\mathrm{eq}}=\int_{0}^{t}\mathrm{d}sh(s)\chi(t-s)\,.$ (155)
+Here, $\chi(s)$ is the response function (recall that $k_{\mathrm{B}}=1$ throughout the manuscript) [64]
+$\chi(s)=-\frac{1}{T}\frac{\mathrm{d}}{\mathrm{d}s}\langle R(0)B(s)\rangle_{\mathrm{eq}}\,.$ (156)
+To connect this to the result we derived in the previous section, we first consider the case a constant perturbation to an equilibrium system, i.e., $h(s)=h$.
+Comparing Eq.
+(154) to Eq.
+(155) (with $\delta\to h$), we see that to first order in $h$
+$\Delta O(t)$ $=\frac{h}{T^{2}}\left(\langle OE\rangle_{\mathrm{eq}}-\langle O(t)E(0)\rangle_{\mathrm{eq}}\right)$
+$=\frac{1}{T^{2}}\int_{0}^{t}\mathrm{d}s\frac{\mathrm{d}}{\mathrm{d}s}\langle O(t)E(s)\rangle_{\mathrm{eq}}h\,.$ (157)
+Since the equilibrium correlation is time-translation invariant, we can rewrite the above response Eq. (157) as
+$\Delta O(t)=\int_{0}^{t}\mathrm{d}s\chi(t-s)h+\mathcal{O}(h^{2})\,$, (158)
+with $\chi(t-s)\equiv\frac{1}{T^{2}}\frac{\mathrm{d}}{\mathrm{d}s}\langle{O}(t)E(s)\rangle_{\mathrm{eq}}=-\frac{1}{T^{2}}\frac{\mathrm{d}}{\mathrm{d}t}\langle O(t-s)E(0)\rangle_{\mathrm{eq}}$ consistent with Eq.
+(156) with $R=T^{-1}E$.
+For time-dependent $h(s)$ we get to first order
+$\Delta O(t)=\left\langle O(t)\int_{s=0}^{s=t}h(s)\sum_{x\neq y}\frac{\tilde{r}_{xy}}{r_{xy}}\mathrm{d}\varepsilon_{xy}(s)\right\rangle\,.$ (159)
+The simplifications performed in Eqs. (151), (152), and (153) can still be made, as these are independent of the perturbation $h(s)$
+$\Delta O(t)$ $=\frac{1}{T^{2}}\int_{0}^{t}h(s)\sum_{i}\sum_{x}O_{i}E_{x}\partial_{s}P(i,t|x,s)p_{x}^{\mathrm{eq}}\mathrm{d}s$
+$=\frac{1}{T^{2}}\int_{0}^{t}h(s)\partial_{s}\langle O(t)E(s)\rangle_{\mathrm{eq}}\mathrm{d}s$
+$=\int_{0}^{t}h(s)\chi(t-s)\,$, (160)
+again with
+$\chi(t-s)$ $=\frac{1}{T^{2}}\frac{\mathrm{d}}{\mathrm{d}s}\langle O(t)E(s)\rangle_{\mathrm{eq}}$
+$=-\frac{1}{T^{2}}\frac{\mathrm{d}}{\mathrm{d}t}\langle O(t-s)E(0)\rangle_{\mathrm{eq}}\,.$ (161)
+The extension of these results, e.g., in case of non-stationary systems, has already been discussed in Ref. [64].
+V.3.2 Response of a Stationary System
+To further connect our result to responses of (non-equilibrium) steady states, we realize that the perturbation rates $\tilde{r}_{xy}$ are the elements of the perturbation generator $\bm{L}^{1}$, i.e., $\bm{L}=\bm{L}_{0}+h\bm{L}^{1}$ with $L_{xy}=r_{yx}$ and $L^{1}_{xy}=\tilde{r}_{yx}$.
+Using the correlation lemma Eq.
+(50), the response of a single-time observable in steady-state reads to first order in $h$
+$\Delta_{\mathrm{s}}O(t)\equiv$ $\langle O(t)\rangle_{h}-\langle O(t)\rangle_{\mathrm{s}}$
+$=$ $\int_{0}^{t}\mathrm{d}sh(s)\sum_{i}O_{i}\sum_{x,y\neq x}\frac{\tilde{r}_{xy}}{r_{xy}}$
+$\times\left[P(i,t|y,s)-P(i,t|x,s)\right]r_{xy}p_{x}^{\mathrm{s}}\,.$ (162)
+The last double-sum simplifies as follows
+$\sum_{y,x\neq y}\tilde{r}_{xy}[P(i,t|y,s)-P(i,t|x,s)]p_{x}^{\mathrm{s}}$
+$=$ $\sum_{y,x\neq y}\tilde{r}_{yx}P(i,t|x,s)p_{y}^{\mathrm{s}}-\underbrace{\sum_{y,x\neq y}\tilde{r}_{xy}}_{\tilde{r}_{xx}}P(i,t|x,s)]p_{x}^{\mathrm{s}}$
+$=$ $\sum_{y,x}\tilde{r}_{yx}P(i,t|x,s)p_{y}^{\mathrm{s}}\,$, (163)
+where we split the sum and relabel $x\leftrightarrow y$ in the first line, and then use the definition of the diagonal elements to simplify the expression.
+As a consequence, we can now write the conjugate observable $R_{x}^{\mathrm{s}}\equiv\sum_{y}L_{xy}^{1}p_{y}^{\mathrm{s}}/p_{x}^{\mathrm{s}}$, and identify that inserting Eq.
+(163) into Eq.
+(162) yields to first order
+$\Delta_{\mathrm{s}}O(t)=$ $\int_{0}^{t}\mathrm{d}sh(s)\sum_{i}O_{i}\sum_{y,x}\underbrace{\tilde{r}_{yx}}_{L_{xy}^{1}}P(i,t|x,s)p_{y}^{\mathrm{s}}$
+$=$ $\int_{0}^{t}\mathrm{d}sh(s)\sum_{i}O_{i}\sum_{x}P(i,t|x,s)\underbrace{\sum_{y}L_{xy}^{1}\frac{p_{y}^{\mathrm{s}}}{p_{x}^{\mathrm{s}}}}_{R^{\mathrm{s}}_{x}}p_{x}^{\mathrm{s}}$
+$=$ $\int_{0}^{t}\mathrm{d}sh(s)\sum_{i,x}O_{i}P(i,t|x,s)R^{\mathrm{s}}_{x}p_{x}^{\mathrm{s}}\,.$ (164)
+Here, we can identify the stationary response function
+$\chi_{\mathrm{s}}(t-s)\equiv{\color[rgb]{0,0,0}\mathbb{1}}_{s<t}\langle O(t)R^{\mathrm{s}}(s)\rangle_{\mathrm{s}}={\color[rgb]{0,0,0}\mathbb{1}}_{s<t}\langle O(t-s)R^{\mathrm{s}}(0)\rangle_{\mathrm{s}}\,$, (165)
+so that Eq. (164) can be written as
+$\Delta_{\mathrm{s}}O(t)=\int_{0}^{t}\mathrm{d}sh(s)\chi_{\mathrm{s}}(t-s)+\mathcal{O}(h^{2})\,.$ (166)
+This is consistent with the results from Ref. [142].
+V.3.3 Response of Path-Wise Observables
+Lastly, we consider the response of, possibly transient, path-wise observables. Specifically, consider the observables
+$O_{1}(t)$ $=\int_{0}^{t}\sum_{x,y\neq x}b_{xy}(s)\mathrm{d}\varepsilon_{xy}(s)\,$, (167)
+$O_{2}(t)$ $=\int_{0}^{t}g_{s}\mathrm{d}s\,.$ (168)
+Using Eq. (48) the response of Eq. (167) can be stated directly to first order
+$\Delta O_{1}(t)\equiv$ $\langle O_{1}(t)\rangle_{h}-\langle O_{1}(t)\rangle$
+$=$ $\int_{0}^{t}\mathrm{d}z\sum_{x,y\neq x}b_{xy}(z)\int_{0}^{t}\mathrm{d}sh(s)\sum_{i,j\neq i}\frac{\tilde{r}_{ij}}{r_{ij}}$
+$\times\delta_{ix}\delta_{jy}\delta(s-z)r_{ij}p_{i}(s)$
+$=$ $\int_{0}^{t}\mathrm{d}z\sum_{x,y\neq x}b_{xy}(z)h(z)\tilde{r}_{xy}p_{x}(z)\,.$ (169)
+Similarly, using Eq. (50), the response of $O_{2}(t)$ becomes to first order (note that this result is compared to the approach used in Ref. [64] in App. A.8)
+$\Delta O_{2}(t)\equiv$ $\langle O_{2}(t)\rangle_{h}-\langle O_{2}(t)\rangle$
+$=$ $\int_{0}^{t}\mathrm{d}z\int_{0}^{t}\mathrm{d}s{\color[rgb]{0,0,0}\mathbb{1}}_{s<z}h(s)\sum_{i,x,y\neq x}$ (170)
+$\times g_{i}(z)\tilde{r}_{xy}p_{x}(s)\left[P(i,z|y,s)-P(i,z|x,s)\right].$
+The sum over $x,y$ can again be simplified using Eq. (163), so that
+$\Delta O_{2}(t)=$ $\int_{0}^{t}\mathrm{d}z\int_{0}^{t}\mathrm{d}s{\color[rgb]{0,0,0}\mathbb{1}}_{s<z}h(s)\sum_{i,x,y}g_{i}(z)\tilde{r}_{yx}p_{y}(s)P(i,z|x,s)$
+$+\mathcal{O}(h^{2})\,.$ (171)
+Defining $R_{x}(s)=\sum_{y}\tilde{r}_{yx}p_{y}(s)/p_{x}(s)=\sum_{y}L_{xy}^{1}p_{y}(s)/p_{x}(s)$}, the response functions for $O_{1}(t)$ and $O_{2}(t)$ become
+$\chi_{O_{1}}(t,s)$ $={\color[rgb]{0,0,0}\mathbb{1}}_{t>s}\sum_{x,y\neq x}b_{xy}(s)\tilde{r}_{xy}p_{x}(s),$ (172)
+$\chi_{O_{2}}(t,s)$ $=\int_{0}^{t}\mathrm{d}z{\color[rgb]{0,0,0}\mathbb{1}}_{z>s}\langle g_{z}R(s)\rangle,$ (173)
+so that the response to any additive functional $O_{\mathrm{p}}(t)=O_{1}(t)+O_{2}(t)$ becomes
+$\Delta O_{\mathrm{p}}(t)\equiv$ $\langle O_{\mathrm{p}}\rangle_{h}-\langle O_{\mathrm{p}}\rangle$
+$=$ $\int_{0}^{t}\mathrm{d}sh(s)\chi_{O_{\mathrm{p}}}(t,s)+\mathcal{O}(h^{2})\,$,
+$\chi_{O_{\mathrm{p}}}(t,s)=$ $\chi_{O_{1}}(t,s)+\chi_{O_{2}}(t,s)\,.$ (174)
+Note that observables of integrals over $\mathrm{d}\bm{n}(s)$ is a special case where $g_{x}(s)$ and $b_{xy}(s)$ have been chosen appropriately to satisfy Eq. (46).
+V.4 Examples
 To visualize the results in Eqs.
@@
 (71) with the free energies listed in Tab.
-4.
+5.
 In Fig.
@@
 11d.
-The diamonds on the x-axis denote the largest relaxation time of the respective generators.
+The diamonds on the horizontal axis denote the largest relaxation time of the respective generators.
 As a visual aid, the value of the equilibrium covariance between $O$ and $E$, $\mathrm{cov}_{\mathrm{eq}}(O,E)=\langle OE\rangle_{\mathrm{eq}}-\langle O\rangle_{\mathrm{eq}}\langle E\rangle_{\mathrm{eq}}$, is included in Fig.
@@
 The free energies, $E_{i}$, used in (a) and (c) are listed in Tab.
-4.
+5.
 The simulations are performed with the celebrated Gillespie algorithm.
@@
 $\mathrm{cov}({\rho}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P(x,\tau;i,\tau^{\prime};v)-p_{x}(\tau;v)p_{i}(\tau^{\prime};v)\right],$
-$\mathrm{cov}({J}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\mathds{1}_{\tau>\tau^{\prime}}r_{xy}P(x,\tau|i,\tau^{\prime};v)p_{i}(\tau^{\prime};v)+\mathds{1}_{\tau<\tau^{\prime}}r_{yx}(v\tau)P(i,\tau^{\prime}|y,\tau;v)p_{y}(\tau;v)-r_{xy}(v\tau)p_{x}(\tau;v)p_{i}(\tau^{\prime};v)\right],$
+$\mathrm{cov}({J}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[{\color[rgb]{0,0,0}\mathbb{1}}_{\tau>\tau^{\prime}}r_{xy}P(x,\tau|i,\tau^{\prime};v)p_{i}(\tau^{\prime};v)+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}r_{yx}(v\tau)P(i,\tau^{\prime}|y,\tau;v)p_{y}(\tau;v)-r_{xy}(v\tau)p_{x}(\tau;v)p_{i}(\tau^{\prime};v)\right],$
 $\mathrm{cov}(J_{t}^{k},{J}_{t}^{l})=$ $\int_{0}^{t}\mathrm{d}\tau\sum_{x,y}\kappa_{xy}^{k}(v\tau)\kappa_{xy}^{l}(v\tau)r_{xy}(v\tau)p_{x}(\tau;v)$
-$+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\mathds{1}_{\tau>\tau^{\prime}}r_{xy}(v\tau)r_{ji}(v\tau^{\prime})P(x,\tau|j,\tau^{\prime};v)p_{i}(\tau^{\prime};v)+\mathds{1}_{\tau<\tau^{\prime}}r_{ij}(v\tau^{\prime})r_{yx}(v\tau)P(i,\tau^{\prime}|y,\tau;v)p_{y}(\tau;v)\right$.
+$+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[{\color[rgb]{0,0,0}\mathbb{1}}_{\tau>\tau^{\prime}}r_{xy}(v\tau)r_{ji}(v\tau^{\prime})P(x,\tau|j,\tau^{\prime};v)p_{i}(\tau^{\prime};v)+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}r_{ij}(v\tau^{\prime})r_{yx}(v\tau)P(i,\tau^{\prime}|y,\tau;v)p_{y}(\tau;v)\right$.
 $-\left.r_{xy}(v\tau)p_{x}(\tau;v)r_{ij}(v\tau^{\prime})p_{i}(\tau^{\prime};v)\right],$ (182)
@@
 With this identity, we can extend the time-dependent TUR in Eq. (183) to also include densities,
-$\rho_{t}$ $=\int_{\tau=0}^{\tau=t}\bm{V}(v\tau)^{T}\mathrm{d}\bm{\tau}(\tau),$ (158)
+[…]
 as well. To do so, we require $\langle A_{t}\rho_{t}\rangle$, which after some simplification using Eq. (50) reads
-$\langle A_{t}\rho_{t}\rangle=$ $\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i}V_{i}(v\tau^{\prime})\mathds{1}_{\tau<\tau^{\prime}}\sum_{x,y}\left[P(i,\tau^{\prime}|y,\tau;v)-P(i,\tau^{\prime}|x,\tau;v)\right]p_{x}(\tau;v)r_{xy}(v\tau)Z_{xy}(v\tau).$ (159)
+$\langle A_{t}\rho_{t}\rangle=$ $\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i}V_{i}(v\tau^{\prime}){\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\sum_{x,y}\left[P(i,\tau^{\prime}|y,\tau;v)-P(i,\tau^{\prime}|x,\tau;v)\right]p_{x}(\tau;v)r_{xy}(v\tau)Z_{xy}(v\tau).$ (187)
 Rearranging the inner sum ($x,y$), we find that Eq. (187) simplifies to
-$\langle$ $A_{t}\rho_{t}\rangle$ (160)
-$=$ $\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{\tau^{\prime}}\mathrm{d}\tau\sum_{i,x}V_{i}(v\tau^{\prime})P(i,\tau^{\prime}|x,\tau;v)\partial_{\tau}p_{x}(\tau;v)$
-$=$ $-\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{\tau^{\prime}}\mathrm{d}\tau\sum_{i,x}V_{i}(v\tau^{\prime})\left[\partial_{\tau}P(i,\tau^{\prime}|x,\tau;v)\right]p_{x}(\tau;v),$
+$\langle$ $A_{t}\rho_{t}\rangle\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{\tau^{\prime}}\mathrm{d}\tau\sum_{i,x}V_{i}(v\tau^{\prime})P(i,\tau^{\prime}|x,\tau;v)\partial_{\tau}p_{x}(\tau;v)=-\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{\tau^{\prime}}\mathrm{d}\tau\sum_{i,x}V_{i}(v\tau^{\prime})\left[\partial_{\tau}P(i,\tau^{\prime}|x,\tau;v)\right]p_{x}(\tau;v),$
 where we performed a partial integration over $\tau$ to obtain the third line.
@@
 (121)] does not reduce to an accessible quantity.
-VIII Outlook
+VIII Connection to Quantum Unraveling
+Stochastic calculus has also found traction in the theory of open quantum systems [13, 50, 25, 27, 9], i.e., in the context of quantum unraveling.
+In the following, we provide some insight into how the classical stochastic-calculus approach presented here is connected to quantum unraveling.
+Before doing so, we introduce some basic notation and theory.
+Subsequently, we highlight the relation to our work.
+Let $\rho_{s}^{\mathrm{u}}$ be a rank-1 matrix denoting a pure state in a discrete-state open quantum system at time $s$. It evolves according to the Belavkin equation [27]
+$\mathrm{d}\rho_{s}^{\mathrm{u}}=\mathcal{B}(\rho_{s}^{\mathrm{u}})\mathrm{d}s+\sum_{i}\left(\frac{\mathcal{J}_{i}(\rho_{s}^{\mathrm{u}})}{\mathrm{Tr}[\mathcal{J}_{i}(\rho_{s}^{\mathrm{u}})]}-\rho_{s}^{\mathrm{u}}\right)\mathrm{d}\tilde{n}_{is},$ (191)
+where $\mathcal{J}_{i}(\rho_{s}^{\mathrm{u}})=\hat{J}_{i}\rho_{s}^{\mathrm{u}}\hat{J}_{i}^{\dagger}$ with jump operator $\hat{J}_{i}$ for the $i$th quantum jump.
+The first term in Eq.
+(191) has the form
+$\mathcal{B}(\rho_{s}^{\mathrm{u}})=-i\hat{H}_{\mathrm{eff}}\rho_{s}^{\mathrm{u}}+\rho_{s}^{\mathrm{u}}\hat{H}_{\mathrm{eff}}^{\dagger}-\rho_{s}^{\mathrm{u}}\mathrm{Tr}(-i\hat{H}_{\mathrm{eff}}\rho_{s}^{\mathrm{u}}+\rho_{s}^{\mathrm{u}}\hat{H}_{\mathrm{eff}}^{\dagger}),$ (192)
+where the Hamiltonian $\hat{H}$ of the system enters through the effective Hamiltonian $\hat{H}_{\mathrm{eff}}=\hat{H}-\frac{i}{2}\sum_{j}\hat{J}_{j}^{\dagger}\hat{J}_{j}$.
+Lastly, $\mathrm{d}\tilde{n}_{is}$ is a stochastic differential which takes values 1 if the $i$th transition occurs in $s,s+\mathrm{d}s]$ and 0 otherwise.
+It has the properties $\langle\mathrm{d}\tilde{n}_{is}\rangle=\mathrm{Tr}(\mathcal{J}_{i}(\rho_{s}))$ and $\mathrm{d}\tilde{n}_{is}\mathrm{d}\tilde{n}_{js}=\delta_{ij}\mathrm{d}\tilde{n}_{is}$.
+Taking the average over noise histories in Eq.
+(191), the resulting density matrix $\rho_{s}=\langle\rho_{s}^{\mathrm{u}}\rangle$ evolves according to the Lindblad equation, often referred to as “quantum master equation”.
+With this short insight into quantum SDE’s, we now turn to the question of how these concepts are related to the work we present in this article.
+Since classical stochastic dynamics should correspond to a special case of quantum open-system evolutions [22], one may at first think that Eq.
+(46) describes the classical counterpart to Eq.
+(191); they both aim to describe a trajectory (the former a classical sequence of discrete states and the latter the evolution of pure states in a Hilbert space) and both include Poissonian noise $\mathrm{d}\bm{n}$ and $\mathrm{d}\tilde{n}$, respectively.
+However, Eq.
+(46) is not the classical counterpart to Eq.
+(191), which can be seen from two observations: (i) Eq.
+(46) is an SDE describing the evolution of transitions and (ii) Eq.
+(191) is a functional of the transitions describing the evolution of the states.
+Moreover, Eq.
+(46) allows us to consider the (stochastic) evolution of transitions occuring in a discrete systems and further enables us to form and study additive functionals of the dynamics.
+Hence, using (i) enables us to study functionals such as (ii).
+To be precise on this last point, we can construct a classical functional of the transitions which is the classical counterpart to Eq.
+(191).
+It turns out that this functional simply is $\mathrm{d}{\color[rgb]{0,0,0}\delta_{x_{s}x}}\equiv\frac{\mathrm{d}}{\mathrm{d}\tau}{\color[rgb]{0,0,0}\delta_{x_{s}x}}\mathrm{d}\tau$, see Eq.
+(212).
+Explicitly, using Eq.
+(46) in Eq.
+(212) yields
+$\mathrm{d}{\color[rgb]{0,0,0}\delta_{x_{s}x}}$ $=\sum_{y\neq x}(\mathrm{d}n_{yx}-\mathrm{d}n_{xy})$ (193)
+$=\sum_{y\neq x}(r_{yx}\delta_{x_{s}y}\mathrm{d}s-r_{xy}\delta_{x_{s}x}\mathrm{d}s)+\sum_{y\neq x}(\mathrm{d}\varepsilon_{yx}-\mathrm{d}\varepsilon_{xy}).$
+The first sum is the deterministic evolution of the state, i.e., the master operator acting on the “population vector” $\delta_{x_{s}x}$ and the second sum is the zero-mean stochastic noise in the system.
+Shifting the mean of the noise increments by adding and subtracting $\sum_{y\neq x}(\langle\mathrm{d}n_{yx}\rangle-\langle\mathrm{d}n_{xy}\rangle)$ [this corresponds to the trace terms in Eq.
+(191)], we recover the classical analogue of Eq.
+(191)
+$\mathrm{d}\delta_{x_{s}x}=$ $\sum_{y\neq x}\left(r_{yx}\delta_{x_{s}y}\mathrm{d}s-\langle\mathrm{d}n_{yx}\rangle-r_{xy}\delta_{x_{s}x}\mathrm{d}s+\langle\mathrm{d}n_{xy}\rangle\right)$
+$+\sum_{y\neq x}(\mathrm{d}\varepsilon_{yx}+\langle\mathrm{d}n_{yx}\rangle-\mathrm{d}\varepsilon_{xy}-\langle\mathrm{d}n_{xy}\rangle).$ (194)
+Indeed, taking the average of Eq.
+(193) w.r.t.
+the noise history recovers the master equation, i.e., the classical analogue of the Lindblad equation, further showing the correspondence between Eqs.
+(191) and (193).
+For completeness it should be mentioned that the quantum analogue of Eq. (46) has already been identified in Ref. [93] as
+$\mathrm{d}\tilde{n}_{it}=\mathrm{Tr}(\mathcal{J}_{i}(\rho_{t}))\mathrm{d}t+\mathrm{d}\tilde{\varepsilon}_{it},$ (195)
+where the noise $\mathrm{d}\tilde{\varepsilon}_{it}$ corresponds to Eq.
+(45) and the statistical properties can be evaluated in a similar fashion; $\langle\mathrm{d}\tilde{\varepsilon}_{it}\rangle=0$ and $\langle\mathrm{d}\tilde{\varepsilon}_{it}\mathrm{d}\tilde{\varepsilon}_{jt^{\prime}}\rangle=\delta_{ij}\delta(t-t^{\prime})\mathrm{Tr}(\mathcal{J}_{i}(\rho_{t}))\mathrm{d}t\mathrm{d}t^{\prime}$.
+While the connection to the unraveling of the Lindblad equation and, thus, the Belavkin equation is therefore established, the question of how this in turn relates to the classical counterpart remains elusive, further motivating our discussion here.
+IX Outlook
 We developed a stochastic-calculus for path-wise observables (i.e., functionals) of Markov-jump processes that unifies the different approaches to Markov-jump dynamics, as well as the descriptions of diffusion and jump dynamics.
@@
 We further derived the response of any (generally path-wise) observable to a general (incl.
-thermal) perturbation.
-Finally, we established the continuum limit to achieve the complete unification of diffusion and jump dynamics.
+thermal) perturbation and constructed a corresponding response-function formalism for single-time as well as path-wise observables.
+Subsequently, we established the continuum limit to achieve the complete unification of diffusion and jump dynamics.
+Lastly, we derived the classical analogue of the Belavkin equation and commented on the analogy with open quantum systems.
 Our results ultimately place functionals of diffusion and jump processes on an equal footing on the level of individual stochastic realizations, and hence achieve a “contraction” of two until now disjoint frameworks in time-average statistical mechanics.
+Our newly developed framework has recently already been applied in [174].
 The results enable, and are expected to inspire, new directions of research.
@@
 Uffbassa! Mir hen da wos g’scheid’s gmochdd!
-IX Data availability
+X Data availability
 The data that support the findings of this article are openly available at [158].
 Acknowledgments
+We thank Jorge ”Jefe” Tabanera-Bravo for discussions on quantum systems.
 The financial support from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program (Grant Agreement No.
@@
 Specifically, we can write
-$\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle=$ $\mathds{1}_{\tau\geq\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$
-$+\mathds{1}_{\tau<\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle,$ (164)
+$\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle=$ ${\color[rgb]{0,0,0}\mathbb{1}}_{\tau\geq\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$
+$+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle,$ (196)
 where the first term can be written as
-$\mathds{1}_{\tau\geq\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$
+${\color[rgb]{0,0,0}\mathbb{1}}_{\tau\geq\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$
 $=\underbrace{\sum_{\mathrm{d}\varepsilon_{xy}}\mathrm{d}\varepsilon_{xy}p(\mathrm{d}\varepsilon_{xy})}_{=0}P(x,\tau|i,\tau^{\prime})p_{i}(\tau^{\prime}).$ (197)
 Since the system is Markovian, (ii) and (iii) do not contribute. Hence, without loss of generality we can write
-$\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle=\mathds{1}_{\tau<\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle.$ (166)
+$\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle={\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle.$ (198)
 For small $\mathrm{d}\tau\to 0$, the noise increment can only take two values
-$\mathrm{d}\varepsilon_{xy}(\tau)=\mathds{1}_{x}(\tau)\begin{cases}1-r_{xy}\mathrm{d}\tau&\text{transition},\\ -r_{xy}\mathrm{d}\tau&\text{no transition}.\end{cases}$ (167)
+$\mathrm{d}\varepsilon_{xy}(\tau)={\color[rgb]{0,0,0}\delta_{x_{\tau}x}}(\tau)\begin{cases}1-r_{xy}\mathrm{d}\tau&\text{transition},\\ -r_{xy}\mathrm{d}\tau&\text{no transition}.\end{cases}$ (199)
 These values occur with probability $r_{xy}\mathrm{d}\tau$ and $1-r_{xy}\mathrm{d}\tau$, respectively. Defining the state function
@@
 allows us to explicitly evaluate Eq. (198) by introducing an intermediate point (see Fig. 12), such that to leading order
-$\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$ $=\mathds{1}_{\tau<\tau^{\prime}}\mathrm{d}\tau^{\prime}\sum_{\mathrm{d}{\varepsilon}_{xy}(\tau)}p(\mathrm{d}{\varepsilon}_{xy}(\tau))\mathrm{d}{\varepsilon}_{xy}(\tau)P(i,\tau^{\prime}|\gamma(\mathrm{d}{\varepsilon}_{xy}(\tau)),\tau+\mathrm{d}\tau)p_{x}(\tau)$
-$=\mathds{1}_{\tau<\tau^{\prime}}\mathrm{d}\tau^{\prime}p_{x}(\tau)r_{xy}\mathrm{d}\tau\left(1-r_{xy}\mathrm{d}\tau\right)\left[P(i,\tau^{\prime}|y,\tau+\mathrm{d}\tau)-P(i,\tau^{\prime}|x,\tau+\mathrm{d}\tau)\right]$
-$=\mathds{1}_{\tau<\tau^{\prime}}\mathrm{d}\tau^{\prime}\left[P(i,\tau^{\prime}|y,\tau)-P(i,\tau^{\prime}|x,\tau)\right]p_{x}(\tau)r_{xy}\mathrm{d}\tau\,.$ (169)
+$\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$ $={\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\mathrm{d}\tau^{\prime}\sum_{\mathrm{d}{\varepsilon}_{xy}(\tau)}p(\mathrm{d}{\varepsilon}_{xy}(\tau))\mathrm{d}{\varepsilon}_{xy}(\tau)P(i,\tau^{\prime}|\gamma(\mathrm{d}{\varepsilon}_{xy}(\tau)),\tau+\mathrm{d}\tau)p_{x}(\tau)$
+$={\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\mathrm{d}\tau^{\prime}p_{x}(\tau)r_{xy}\mathrm{d}\tau\left(1-r_{xy}\mathrm{d}\tau\right)\left[P(i,\tau^{\prime}|y,\tau+\mathrm{d}\tau)-P(i,\tau^{\prime}|x,\tau+\mathrm{d}\tau)\right]$
+$={\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\mathrm{d}\tau^{\prime}\left[P(i,\tau^{\prime}|y,\tau)-P(i,\tau^{\prime}|x,\tau)\right]p_{x}(\tau)r_{xy}\mathrm{d}\tau\,.$ (201)
 Figure 12: Visualization of the intermediate point considered in the proof of Noise-Time correlation lemma.
@@
 Using the last two equations in Eqs. (49), we get
-$\frac{\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle}{r_{xy}\mathrm{d}\tau\mathrm{d}\tau^{\prime}}=$ $\mathds{1}_{\tau<\tau^{\prime}}P(i,\tau^{\prime}|y,\tau)p_{x}(\tau)$
-$+\mathds{1}_{\tau\geq\tau^{\prime}}P(x,\tau|i,\tau^{\prime})p_{i}(\tau^{\prime}).$ (171)
+$\frac{\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle}{r_{xy}\mathrm{d}\tau\mathrm{d}\tau^{\prime}}=$ ${\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}P(i,\tau^{\prime}|y,\tau)p_{x}(\tau)$
+$+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau\geq\tau^{\prime}}P(x,\tau|i,\tau^{\prime})p_{i}(\tau^{\prime}).$ (203)
 A.2.2 Jump-Jump Correlations
@@
 Using Eqs. (49) therefore yields
-$\frac{\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}n_{ij}(\tau^{\prime})\rangle}{r_{xy}r_{ij}\mathrm{d}\tau\mathrm{d}\tau^{\prime}}=$ $\delta_{ix}\delta_{jy}\delta(\tau-\tau^{\prime})p_{x}(\tau)$
-$+\mathds{1}_{\tau<\tau^{\prime}}P(i,\tau^{\prime}|y,\tau)p_{x}(\tau)$
-$+\mathds{1}_{\tau\geq\tau^{\prime}}P(x,\tau|j,\tau^{\prime})p_{i}(\tau^{\prime}).$ (173)
+$\frac{\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}n_{ij}(\tau^{\prime})\rangle}{r_{xy}r_{ij}\mathrm{d}\tau\mathrm{d}\tau^{\prime}}=$ ${\color[rgb]{0,0,0}r_{ij}}\delta_{ix}\delta_{jy}\delta(\tau-\tau^{\prime})p_{x}(\tau)$
+$+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}P(i,\tau^{\prime}|y,\tau)p_{x}(\tau)$
+$+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau\geq\tau^{\prime}}P(x,\tau|j,\tau^{\prime})p_{i}(\tau^{\prime}).$ (205)
 A.3 From Equations of Motion to Ensemble Description
@@
 (39).
-The calculation is well ḱnown (see e.g., Ref.
+The calculation is well known (see e.g., Ref.
 [51]).
-Let $K(x)$ be an arbitrary test function that obeys the boundary conditions.
+Let $K(x)$ be an arbitrary test function that is sufficiently smooth and obeys the boundary conditions.
 Considering the expectation $\langle K(\bm{x}(t))\rangle=\int\mathrm{d}\bm{x}K(\bm{x})P(\bm{x},t)$ for the dummy function at some time $t$, the temporal derivative can be evaluated using Itô’s lemma [51]
 $\frac{\mathrm{d}}{\mathrm{d}t}\langle K(\bm{x}(t))\rangle=\langle\bm{F}(\bm{x})\cdot\nabla K(\bm{x})+\nabla\cdot\bm{D}\nabla K(\bm{x})\rangle$ (206)
-$=\int\mathrm{d}\bm{x}\left[\bm{F}(\bm{x})\cdot\nabla K(\bm{x})+\nabla\cdot\bm{D}\nabla K(\bm{x})\right]P(\bm{x},t).$
+$=\int\mathrm{d}\bm{x}{\color[rgb]{0,0,0}P(\bm{x},t)}\left[\bm{F}(\bm{x})\cdot\nabla K(\bm{x})+\nabla\cdot\bm{D}\nabla K(\bm{x})\right].$
 Equivalently, the derivative can be written as
@@
 $\int$ $\mathrm{d}\bm{x}K(\bm{x})\frac{\partial}{\partial t}P(\bm{x},t)$
-$=$ $\int\mathrm{d}\bm{x}\left[\bm{F}(\bm{x})\cdot\nabla K(\bm{x})+\nabla\cdot\bm{D}\nabla K(\bm{x})\right]P(\bm{x},t)$ (176)
+$=$ $\int\mathrm{d}\bm{x}{\color[rgb]{0,0,0}P(\bm{x},t)}\left[\bm{F}(\bm{x})\cdot\nabla K(\bm{x})+\nabla\cdot\bm{D}\nabla K(\bm{x})\right]$ (208)
 $=$ $\int\mathrm{d}\bm{x}K(\bm{x})\left[-\nabla\left\{\bm{F}(\bm{x})P(\bm{x},t)\right\}\right.+\left.\nabla\cdot\bm{D}\nabla P(\bm{x},t)\right]\,$,
@@
 To show how the master equation emerges from the stochastic equations of motion in Eq.
-(38), we first need to consider the indicator function $\mathds{1}_{x}(\tau)$ of being in state $x$ at time $\tau$.
-Consider $\mathds{1}_{x}(\tau+\mathrm{d}\tau)=\mathds{1}_{x}(\tau)+\frac{\mathrm{d}}{\mathrm{d}\tau}\mathds{1}_{x}(\tau)\mathrm{d}\tau+\mathcal{O}(\mathrm{d}\tau^{2})$, where the derivative should be understood as
-$\frac{\mathrm{d}}{\mathrm{d}\tau}\mathds{1}_{x}(\tau)=\lim_{\mathrm{d}\tau\to 0^{+}}\frac{\mathds{1}_{x}(\tau+\mathrm{d}\tau)-\mathds{1}_{x}(\tau)}{\mathrm{d}\tau}.$ (178)
+(46), we first need to consider the indicator function $\color[rgb]{0,0,0}\delta_{x_{\tau}x}$} of being in state $x$ at time $\tau$.
+Consider ${\color[rgb]{0,0,0}\delta_{x_{\tau+\mathrm{d}\tau}x}}={\color[rgb]{0,0,0}\delta_{x_{\tau}x}}+\frac{\mathrm{d}}{\mathrm{d}\tau}{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}\mathrm{d}\tau+\mathcal{O}(\mathrm{d}\tau^{2})$, where the derivative should be understood as
+$\frac{\mathrm{d}}{\mathrm{d}\tau}{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}=\lim_{\mathrm{d}\tau\to 0^{+}}\frac{{\color[rgb]{0,0,0}\delta_{x_{\tau+\mathrm{d}\tau}x}}-{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}}{\mathrm{d}\tau}.$ (210)
 By the non-explosive property of MJP [21] a finite number of jumps occur in any finite time interval with probability 1.
 Therefore, we do not need to consider any higher order correction, as the probability of a second transition occurring in an infinitesimal increment vanishes.
-The term $\frac{\mathrm{d}}{\mathrm{d}\tau}\mathds{1}_{x}(\tau)\mathrm{d}\tau$ can therefore take on values
+The term $\frac{\mathrm{d}}{\mathrm{d}\tau}{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}\mathrm{d}\tau$ can therefore take on values
 […]
 As a consequence, we can write
-$\frac{\mathrm{d}}{\mathrm{d}\tau}\mathds{1}_{x}(\tau)\mathrm{d}\tau=\sum_{y\neq x}\left(\mathrm{d}n_{yx}(\tau)-\mathrm{d}n_{xy}(\tau)\right),$ (180)
-because at most one jump differential will be non-zero as $\mathrm{d}\tau\to 0$. The probability of being in a state $x$ at time $\tau$ is $p_{x}(\tau)=\langle\mathds{1}_{x}(\tau)\rangle$. Therefore,
-$p_{x}(\tau+\mathrm{d}\tau)-p_{x}(\tau)$ $=\left\langle\mathds{1}_{x}(\tau+\mathrm{d}\tau)-\mathds{1}_{x}(\tau)\right\rangle$
-$=\left\langle\frac{\mathrm{d}}{\mathrm{d}\tau}\mathds{1}_{x}(\tau)\mathrm{d}\tau\right\rangle$ (181)
-$=\left\langle\sum_{y\neq x}(\mathrm{d}n_{yx}(\tau)-\mathrm{d}n_{xy}(\tau))\right\rangle$
-$=\mathrm{d}\tau\sum_{y\neq x}\left(r_{yx}p_{y}(\tau)-r_{xy}p_{x}(\tau)\right)$
+$\frac{\mathrm{d}}{\mathrm{d}\tau}{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}\mathrm{d}\tau=\sum_{y\neq x}{\color[rgb]{0,0,0}[}\mathrm{d}n_{yx}(\tau)-\mathrm{d}n_{xy}(\tau){\color[rgb]{0,0,0}]},$ (212)
+because at most one jump differential will be non-zero as $\mathrm{d}\tau\to 0$.
+The probability of being in a state $x$ at time $\tau$ is $p_{x}(\tau)=\langle{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}\rangle$.
+Therefore,
+$p_{x}(\tau+\mathrm{d}\tau)-p_{x}(\tau)$ $=\left\langle{\color[rgb]{0,0,0}\delta_{x_{\tau+\mathrm{d}\tau}x}}-{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}\right\rangle$
+$=\left\langle\frac{\mathrm{d}}{\mathrm{d}\tau}{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}\mathrm{d}\tau\right\rangle$ (213)
+$=\left\langle\sum_{y\neq x}{\color[rgb]{0,0,0}[}\mathrm{d}n_{yx}(\tau)-\mathrm{d}n_{xy}(\tau){\color[rgb]{0,0,0}]}\right\rangle$
+$=\mathrm{d}\tau\sum_{y\neq x}{\color[rgb]{0,0,0}[}r_{yx}p_{y}(\tau)-r_{xy}p_{x}(\tau){\color[rgb]{0,0,0}]}$
 Rearranging Eq. (213) and taking the limit $\mathrm{d}\tau\to 0$ yields the master equation (40) in a way equivalent to the Fokker-Planck equation (209).
@@
 $\langle$ $A_{t}J_{t}^{\mathrm{II}}\rangle$
-$=$ $\frac{1}{2}\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}\mathds{1}_{\tau<\tau^{\prime}}$
+$=$ $\frac{1}{2}\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}$
 $\times\sum_{x,y}\left[P(i,\tau^{\prime}|y,\tau)-P(i,\tau^{\prime}|x,\tau)\right]\left[p_{x}(\tau)r_{xy}-p_{y}(\tau)r_{yx}\right]$
-$=$ $-\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}\mathds{1}_{\tau<\tau^{\prime}}$ (182)
+$=$ $-\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}$ (214)
 $\times\sum_{x}P(i,\tau^{\prime}|x,\tau)\sum_{y,y\neq x}\left[p_{x}(\tau)r_{xy}-p_{y}(\tau)r_{yx}\right].$
@@
 $=$ $\int_{0}^{t}\mathrm{d}\tau^{\prime}\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}\left[-\sum_{x}P(i,\tau^{\prime}|x,0)p_{x}(0)\right$.
-$-\left.\int_{0}^{t}\mathrm{d}\tau\sum_{x}p_{x}(\tau)\partial_{\tau}\left\{P(i,\tau^{\prime}|x,\tau)\mathds{1}_{\tau<\tau^{\prime}}\right\}\right]$
+$-\left.\int_{0}^{t}\mathrm{d}\tau\sum_{x}p_{x}(\tau)\partial_{\tau}\left\{P(i,\tau^{\prime}|x,\tau){\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\right\}\right]$
 $=$ $-\langle J_{t}\rangle+\int_{0}^{t}\mathrm{d}\tau^{\prime}\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}\partial_{\tau^{\prime}}\left[\tau^{\prime}p_{i}(\tau^{\prime})\right],$ (215)
-where we use that $\partial_{\tau}P(i,\tau^{\prime}|x,\tau)=-\partial_{\tau^{\prime}}P(i,\tau^{\prime}|x,\tau)$, $\partial_{\tau}\mathds{1}_{\tau<\tau^{\prime}}=-\partial_{\tau^{\prime}}\mathds{1}_{\tau<\tau^{\prime}}$, and $\int_{0}^{\tau^{\prime}}\mathrm{d}\tau\mathds{1}_{\tau<\tau^{\prime}}=\tau^{\prime}$.
+where we use that $\partial_{\tau}P(i,\tau^{\prime}|x,\tau)=-\partial_{\tau^{\prime}}P(i,\tau^{\prime}|x,\tau)$, $\partial_{\tau}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}=-\partial_{\tau^{\prime}}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}$, and $\int_{0}^{\tau^{\prime}}\mathrm{d}\tau{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}=\tau^{\prime}$.
 With another integration by parts, the correlator finally becomes
@@
 IV.1 with the assumption that $\epsilon>1$.
-We start by realizing that $\Delta\bm{V}\equiv\bm{V}^{\mathrm{A}}-\langle{V}^{\mathrm{A}}\rangle\bm{1}=\frac{1}{2}(1,1,-1,-1)^{T}$.
+We start by realizing that $V^{\Delta}_{\tau}=V_{\tau}-\langle V_{\tau}\rangle=\bm{\delta}(\tau)^{T}\Delta\bm{V}$ with $\bm{\delta}(\tau)^{T}=(\delta_{x_{\tau}1},\delta_{x_{\tau}2},\delta_{x_{\tau}3},\delta_{x_{\tau}4})$ and $\Delta\bm{V}\equiv\bm{V}^{\mathrm{A}}-\langle{V}^{\mathrm{A}}\rangle\bm{1}=\frac{1}{2}(1,1,-1,-1)^{T}$.
 Hence, we can write
@@
 (147) to obtain
-$\partial_{\delta}O_{t}=$ $\lim_{\delta\to 0}\frac{1}{\delta}\left(\left\langle O(t)\frac{\mathrm{d}\mathds{P^{\delta}}}{\mathrm{d}\mathds{P}}\right\rangle-\langle O(t)\rangle\right)$
+$\partial_{\delta}O_{t}=$ $\lim_{\delta\to 0}\frac{1}{\delta}\left(\left\langle O(t)\frac{\mathrm{d}\mathbb{P^{\delta}}}{\mathrm{d}{\color[rgb]{0,0,0}\mathbb{P}}}\right\rangle-\langle O(t)\rangle\right)$
 $=$ $\frac{1}{T^{2}}\left\langle O(t)\int_{s=0}^{s=t}\sum_{x,y\neq x}\left(V_{xy}+T\right)\mathrm{d}\varepsilon_{xy}(s)\right\rangle$ (235)
@@
 Now let the observable $O(t)$ be a dwell-time integral
-$O(t)=\int_{s=0}^{s=t}\bm{g}^{T}\mathrm{d}\bm{\tau}(s),$ (208)
-with some state function vector $\bm{g}$. The response is
+$O(t)=\int_{s=0}^{s=t}{\color[rgb]{0,0,0}g_{s}\mathrm{d}s},$ (240)
+with some state function $g_{\tau}\equiv\sum_{k}\delta_{x_{\tau}k}g_{k}$. The response is
 $\partial_{\delta}O_{t}$ $=\lim_{\delta\to 0}\frac{1}{\delta}\int_{0}^{t}\mathrm{d}s\sum_{x}g_{x}\left[p^{\delta}_{x}(s)-p_{x}(s)\right]$ (241)
@@
 Conversely, employing Eqs. (147) and (50) yields
-$\partial_{\delta}O_{t}=\sum_{i,x,y\neq x}\int_{0}^{t}\mathrm{d}\tau$ $\int_{0}^{t}\mathrm{d}s\mathds{1}_{s<\tau}g_{i}\tilde{r}_{xy}p_{x}(s)$ (214)
+$\partial_{\delta}O_{t}=\sum_{i,x,y\neq x}\int_{0}^{t}\mathrm{d}\tau$ $\int_{0}^{t}\mathrm{d}s{\color[rgb]{0,0,0}\mathbb{1}}_{s<\tau}g_{i}\tilde{r}_{xy}p_{x}(s)$ (246)
 $\times\left[P(i,\tau|y,s)-P(i,\tau|x,s)\right],$
@@
 (147) applies for jump-integrated observables, e.g., currents, as well.
-Lastly, although not explicitly written, the matrix $\bm{b}$ and vector $\bm{g}$ may also depend on time. Equations (205) and (213) are written to readily accommodate for this.
+Lastly, although not explicitly written, the matrix $\bm{b}$ and functions $g_{k}$ may also depend on time. Equations (237) and (245) are written to readily accommodate for this.
 A.9 Time-Dependent Rates
@@
 Figure 14: Shown are quality factor of the CTUR in (a-c) using values of $c(t)$ shown in (d-f) in the calmodulin model.
-The values of $c(t)=c^{*}(t)+\frac{\gamma}{5}(c^{*}(t)-c^{\dagger}(t))$ are shown for values $10\leq\gamma\leq 10$, where $\gamma=-10$ is black and $\gamma=10$ is yellow.
+The values of $c(t)=c^{{\color[rgb]{0,0,0}\text{opt}}}(t)+\frac{\gamma}{5}(c^{{\color[rgb]{0,0,0}\text{opt}}}(t)-c^{{\color[rgb]{0,0,0}\min}}(t))$ are shown for values $10\leq\gamma\leq 10$, where $\gamma=-10$ is black and $\gamma=10$ is yellow.
 The initial condition is $p_{i}(0)=\delta_{i6}$ and the density uses $V_{i}(\tau)=\delta_{i3}$ in each panel.
 The transition weights are $\bm{\kappa}^{(12)}$ in (a) and (d), $\bm{\kappa}^{(56)}$ in (b) and (e), and $\bm{\kappa}^{(26)}$ in (c) and (f).
-The most optimal $c(t)=c^{*}(t)$ is marked by diamonds and the least optimal $c(t)=c^{\dagger}(t)$ is marked by squares.
+The most optimal $c(t)=c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$ is marked by diamonds and the least optimal $c(t)=c^{{\color[rgb]{0,0,0}\min}}(t)$ is marked by squares.
 In (a-c) the $\mathcal{Q}$ for $c=0$ is included as a blue line.
@@
 14c.
-Various values of $c(t)$ are used, including $c(t)=0$, $c^{\dagger}(t)$, and $c^{*}(t)$.
+Various values of $c(t)$ are used, including $c(t)=0$, $c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$, and $c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$.
 The absolute values of the considered $c(t)$ are shown in Fig.
@@
 14f.
-As can be easily seen, the behavior of $c^{\dagger}(t)$ and $c^{*}(t)$ is highly nontrivial and changes greatly in magnitude.
+As can be easily seen, the behavior of $c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$ and $c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$ is highly nontrivial and changes greatly in magnitude.
 A.11 Simulation Parameters
+Table 3: Transition rates used in the calmodulin example from Ref. [12] which are adapted from Ref. [156].
 Transition $i\to j$ Rate $r_{ij}$
@@
 $6\to 2$ 2.922
-Table 2: Transition rates used in the calmodulin example from Ref. [149] which are adapted from Ref. [88].
+Table 4: Values of fixed parameters in SAT simulations which are adapted from Ref. [16].
 Symbol Value
@@
 $e^{\mathrm{in}}_{B}$ 2
-Table 3: Values of fixed parameters in SAT simulations which are adapted from Ref. [146].
+Table 5: Overview of free energies for the four state ring with equilibrium rates.
 State $E_{i}$
@@
 44 $0.2$
-Table 4: Overview of free energies for the four state ring with equilibrium rates.